A Complete Bibliography of Publications in Advances in Computational Mathematics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

05 February 2018 Version 1.00

Title word cross-reference

(3, 1) [DOZ94]. (m, l, s) [BL99]. −1 [Cum96]. 1 [FF05, GT14, Han05, JRS15, NV10, NSW16, TZ07]. 1 < p < ∞ [LWY06]. 2 [Boc04, DL98, GM14, HKR04, Lee03, Möß10, Pot04]. 2.1933 [Pla00]. 3 [CK11, GL13, Hem95, Hem96, HS13, MU14, VW02, vdHS00]. 4 [Han98, Jia11]. 4n [SC08]. 6 [DLUS17]. [−1, 1] [CT05, CKK12]. Ax = b [Cul96]. c [Rip99, Sch11b]. C^{0} [SS14]. C^{1} [KJ09, LL04, LPSSP00, NZ04, NSZ04, Sab04, SS04, DL98]. C^{2} [AF96, Rem12]. C^{3} [CGM01, CM03]. C^{k} [MSS02, PR99]. C^{r} [LLS96, LL98]. C^{ρ} [LLS94]. d [HX02, Rob98]. E [MP96, Pré94]. ℓ^{1} [GN08, GS13, MZ13]. ℓ_{1} [SHTS14, ZYY16]. ℓ_{1}, 2 [Fli18]. F (z + 1) = b^{F (z)} [PC17]. F_{A} [BW15]. F_{D} [BW15, TZ05]. GC_{n} [BH17]. H [Dai16, MS00, Peñ11, GH06, Xu04a, ZQ13, ZG16]. H^{−1} [She09]. H_{F}^{p} [HLT11]. H_{1} [MSS02]. hp [MW01]. L^{2} [CLZ02, LL10, LL12b, MS16]. L^{2}(R) [CHL17]. L^{2}(R^{d}) [CR08]. L^{∞} [Noa14]. L^{p} [War13, RS99]. L^{p}(R) [LWY06]. L_{0} [TGA96]. L_{1} [LW04, DSS09]. l_{2} [DSS09]. L_{2}(R) [Atr12]. L_{p} [Jia95]. L_{p}(R^{d}) [DL04]. L_{r} [CDP09]. M [CTZ02, Jia00, Tem98]. R^{d} [War13]. R^{n} [Sor18]. S^{3} [Grä12]. H_{∈} [BGH15]. n [BH17, DKLT93, SC08]. O(N) [GM14]. P [Dai16, Peñ11, CS03, GH06, Jan98, SL12]. P_{0} [ZT15]. P_{1} [ZT15]. π(x) [Kot08]. q [Rob98]. QR [CM99, LMV11]. R^{2} [KL00]. R^{3} [KL00]. s [Hes09, KL04, KL07]. S^{2} [HL08, Hes09]. σ [SX07]. SO(3) [Grä12]. τ [gTpM02]. V [ZXC12]. ε [GH16]. φ [WZtX16]. W [CGH^{+}96, IB17].

-adaptive [ZQ13, ZG16]. -algorithm [MP96, Pré94]. -analysis [ZYY16]. -bases [MS00]. -channel [Jia00]. -convergence [CDP09]. -d [NV10, TZ07, DL98, FF05, VW02, JRS15]. -energy [Hes09]. -error [BGH15]. -estimate [CLZ02]. -finite [LLS96]. -fold [Jia11]. -Frames [CS03]. -function [IB17]. -functions [WZtX16]. -fundamental [LL98]. -geometrically [ZXC12]. -harmonic [Xu04a]. -isometry [SL12]. -matrices [Dai16, Peñ11]. -methods [gTpM02]. -minimization [Fli18]. -monotone [KL04, KL07]. -orthogonal [SX07]. -penalty [GS13]. -piecewise [PR99]. -point [DLUS17, SC08]. -poised [BH17]. -refinable [CTZ02]. -regularization [MZ13, SHTS14]. -sphere [HKR04, Pot04]. -splines [BL99, MSS02, NZ04]. -stable [TGA96]. -subgradient [GH16]. -subspaces [HLT11]. -superconvergence [ZSB01a]. -synthesis [ZYY16]. -term [Tem98]. -version [Jan98]. -wave [HX02]. -weighted [MS16]. -widths [DKLT93].

1 [BF94]. 19.07.1998 [Sil99].

29 [Fus08a].

4 [Hem96]. 409b [CN96].

96c [Hem96].

a-posteriori [MRH15]. absence [GM14]. Acceleration [Pré94, XY15, DMMS11, LZY98, MP96, WMB13]. accelerations [Noa14]. Accuracy [FF05]. accurate [Del17, HBMY14]. acoustic [ARPR01, BBR02, HR02, ZP06]. action [Las16]. active [KPY09]. active-set [KPY09]. Adams [Wil98]. adapted [AL13, Che06, CZS16, Str13]. Adaptive [Ano01b, BDD06, BDMR10, BLS02, DFR07, FYL08, Hol01, KHM16, MSZ06, PP97, QW11, RWT13, XZ03, ASU17, APV14, BKN11, BMK15, BCE^{+}09, BF94, CCE^{+}12, DMMS11, Dũn11, Dũn12, KCW17, KC16, RSZ11, SW96, WLPV15, ZL16, ZQ13, ZG16]. adaptively [GJS14]. adaptivity [AK01]. adding [RY13]. additional [CaL06]. additive [Bre95, PTC12]. Adjoint [LSLS14, FGO14, FMPS13]. Adv. [Fus08a, Hem96]. Advances [Ano02d, BOP^{+}15, Ano03c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano16]. advection [CLZW13]. aerial [ZB99]. aeroacoustic [CEL15]. Affine [CSS98, Jia03, CHSS03, Mai06, PW94]. affine-invariant [Mai06]. age [Kim06]. age-structured [Kim06]. algebra [BKN11, BP11, CLM02, FH05]. Algebraic [JF02, RSA14, CGR13, DMMS11, Hau97, LMV11, Pan95, SXYY06]. algorithm [BMB14, BPK96, CP99, Che02, CX04, zCpST09, CZS16, DLK12, DGK15, FKS05, GM14, Han05, HZHL12, HFH15, JZ10, KPY09, KC16, LD07, Maz07, MU14, MP96, Mur08, Pré94, QW11, Rip99, Rob98, RHZ17, Ven94, nCV13, WW12, WLW16, XWL13, ZSHZ15, ZXC12, Zwi94, dH94]. Algorithmic [Cai02, GH99]. Algorithms [Ano02d, FMR00, GN01, BGM93, BEPS96, CHM18, DT96, DT12, FJ99, FRT09, Gon93, GN08, GH16, LLS09, LT06, Li04, LSXZ15, LT12, MSZ06, MSXZ13, Sab04, Tem98, Tem00, Tem01, Tem02, VM17, VK96, WLPV15, XZ01, Yin07]. all-electron [SL15]. Allen [RWT13]. almost [GP95, yGmW98]. alternate [MP96]. Alternative [LPY10, Sch11b]. American [lZmCRX03]. Ampère [Awa15]. analog [LPY10]. analyses [KKL03, MNC16]. Analysis [ADL11, JS05, KS17, KvL95, KL16, Lee03, VW02, dFGAJN18, Ang00, AK15, BMSR^{+}16, BEPS96, CDP03, CCSS09, CCG10, Che06, CGH^{+}13, CKW10, DG13, DLZ16, Flo96a, FM04, yGW07, GL17, JK08, KK05, KA97, Lee16, Lev99, LMV11, LY08, Lut04, MWW13, Nie98, OS09, OBS15, Onc14, PM17, PS98, RAB^{+}09, RHZ17, SXZ06, SLN14, TZ03, TS05, VW05, WYW11, WZ10, WOBL17, XB05, Xu14, YDF97, Yao16, Yin07, ZSB01b, ZGGW12, ZYY16, Zwi94]. Analytic [CSW96, MXO13, Kot08, Pet00, RAB^{+}09, Sto96]. analyzing [KL15]. angled [MB96]. angled-derivative [MB96]. angular [SF14]. anisotropic [CGRS15, GO95, HS03, Kun01, Pei05, ZC14]. Announcement [Ano01c]. anomalous [ZD17]. antenna [SFDE15]. antennas [MKS02, Ryn00]. Antisymmetric [Jia03]. any [KXZ04, Mon09]. Appell [BCZ05]. Appending [Kun95]. Application [FE15, LL12a, Nee96, lZmCRX03, BEG17, BDD06, CLL08, CZS16, CGS93, FRT09, yGSlX03, Li10a, LD07, MU14, PZ16, Pré94, SL12, Wal95, ZL16]. applications [ABI15, ByLl01, DST04, DS13, yGS08, HKR04, HAS05, HX17, JGW15, JW95, Low05, MRS18, RT14, RZ10, RSZ11, SW06, Sto96, WZtX16, vdHS00]. applied [AC05, BP11, CEL15, Smo07, ZB12]. Approach [Zoz15b, BPJ02, CCSS09, CHRX06, DGMES13, FY13, Grä12, HT12, JKK^{+}12, KvD01, LSW17, Leo02, Pan95, Wal95, ZD14]. approaches [ZD17, Zoz15a]. approximants [BCZ05, CDTV99, VK96]. Approximate [CLMR10, FPR^{+}12, FGO14, GN01, FOW14, IMS99, JLZ08, KK05, Kim06, NBL11, Pan95]. approximated [GLS15]. approximately [DM15]. approximating [CGR11]. Approximation [AB11, BDKY02, BGN17, BR15, CP04, KP04, Le 05, Li09, Lut04, Mha93, MNW99, NSW07, Pot04, SS16, SS08, Xu04a, ZJ06, AK01, ASS10, BX94, BC99, BMK15, BS05, BA15, BW15, CLLS06, Cas07, CL07, CLM96b, CP07b, CG17, DPS93, Dar03, DY01, EGL13, FJ99, Flo96a, FHN94, FW15, GHdN12, Grä12, GS10, yGmW98, Han12, HW00, KORS17, KL04, KL07, KvD01, LS98, LW04, LW14, LT06, Li94, Li03, Li10a, Li10b, LT98, LB93, LL07b, LP14, MM00, Mai05, MN15, MRH15, MNW96, MWW13, MS04, MS16, NW04, NS04, Nee96, NG99, NZ04, PS98, Spe17, Tem98, TYY11, TDG16, War13, WH07, YZ08, Yoo01]. approximations [ByLl01, CCG10, Che06, CGH^{+}13, DG13, DG16, FF05, yGSlX03, yGS08, HL05, HR01, HX02, IB17, IMS99, JGW15, Kot08, MRS03, YLBL12, Ye00]. arbitrarily [Han10]. arbitrary [CMP07a, CHR00, HW00, SL05, Ven94]. arc [Leo02]. architecture [Jan98]. Area [HR17]. arising [AM14, BNP14, JM00, LN14, LL99, LSXZ15, MV01, TBD94]. Arnoldi [WZtX16]. array [AH96]. arrays [Cal05, SFDE15]. art [FRT09]. ARTICLE [Zoz15b]. artificial [Mha93]. aspects [Le 00, Sau95, Yoo01, dB00]. asset [PGB15]. Associated [GH03, ADL11, CW09, HLT11, KA97, LPSSP00, NSW07, SM00]. association [New96]. astrodynamics [ABI15]. Asymptotic [AV15, CL07, FL00, HWZ13, JLZ08, RL07, DHO16, DGS18, LL07a, MWW13]. Asymptotically [TDG16]. Asymptotics [IS13, DS01]. Atmospheric [CLM96a]. atoms [FGO14, Füh16]. attractors [CTZ02]. augmentation [CWX06]. augmented [WLW16, XWL13]. Author [Ano00a, Ano01a, Ano02a, Ano02b, Ano03a, Ano03b, Ano04a, Ano04b]. autocorrelation [BD10, Plo95]. Automated [Nes16]. Automatic [Hag98]. autonomous [MO16]. avalanche [IL16]. average [ABD08, Atr12, Li03, LD07, RJ00]. averaging [MSZ06]. axial [Jia11]. axis [CCH^{+}08]. axisymmetric [BH02, FKS05, KNV01].

B [CHRX06, CM12, FGS16, Flo94, JKLY13, KK99, LM06, MXZ08, MH17, NSS04, Plo95, Str95, Wal95]. B-spline [CHRX06, FGS16, Flo94, KK99, MXZ08, MH17, Wal95]. B-splines [CM12, JKLY13, LM06, NSS04, Plo95, Str95]. backward [AK15, HT11]. backward-Euler [AK15]. balanced [CT17b]. ball [Li08b, Xu04b, DP06]. balls [Noa96]. Banach [CS03, DST04, Tem01, Tem07]. band [MNW04]. band-dominant [MNW04]. banded [AH96]. bandlimited [FPR^{+}12, GP14, Lem09]. bank [CMX07]. banks [EGL13, Jia00, Jia09b]. Barycentric [War96, WSHD07, FHK06, GRB12]. basal [IL16]. based [AS02, AV15, AB97, ABD08, BCM99, Ben97, BS97b, BS97a, BS00, CZS16, DL98, DS10, DGMES13, DAP13, DLK12, Dũn09, FM14, GS13, HL05, Han05, HH15, KA97, Mac94, MSZ06, MW01, MC10, Nai12, Nes16, NBL11, SS16, SS15, TQ17, TSZ98, VK96, Wat96, XZ01, Zha01]. Bases [LW03, LP04, Maz04, DS00, DP06, GJS14, GLT93, HHS04, Hua06, JL06, Li09, LWY06, MS00, QWXZ10, Sab04, TSY10, TYY11]. basis [ASU17, AZW15, AM14, AS05, BCM99, BGH15, Cai02, CP93, CEL15, CL16, DSW05, DP13, Fas99a, Fas99b, FF05, FS98, FW15, Hub12, IUV17, Jüt98, Le 05, LST11, LW14, LM14, LLY06, LL99, LR07, Li05a, Li10a, Li10b, LK05, Low05, Mai05, MRH15, RS99, Rip99, Sch95, SL15, SS08, SRFH12, WH07, Wie15]. be [WS01]. Beamlets [She09]. Bedrosian [TSY10]. behavior [LL07b, RL07]. behaviour [Mur08, Smo07]. Bell [Pit16]. Bell-shaped [Pit16]. below [YLBL12]. Beltrami [KN14, KNQ16, PTC12]. BEM [ADG17, CP15, HS03, MS98]. Bending [Alb15]. Bernstein [Sab04, CMP07a, CP93, Jüt98, LKW17, Sor18, WSM13]. Bernstein-type [Sab04]. Berzolari [CG15].

Bessel [LYY13, Lóp18, Mac94]. Best [BDKY02, Rau05, Li94, Tem98]. between [BWX97, BCE^{+}09, FK10, HR17, PZ16, SM00, She13]. Beyond [GN08, Buh06, Maz01]. Bézier [CC94, DP13, Flo94, LKW17, Sor18]. bi [JS99]. BiCG [Cul96]. bicubic [KP04]. bifurcations [CGS93, ER99]. biharmonic [CL07, CLL08, FKS05, HSY16, Jeo98, JM03, Li10a]. bilateral [CW09]. bilaterally [KPY09]. bilayers [Alb15]. bilinear [BA15, Rei93]. binary [LLY06, Lev99]. binder [SEE96]. Binomial [BMSZ01]. binormal [KKV15]. Biorthogonal [Jia11, CHR00, CL02, Kei95, KKL03, KKLY10, PST08]. bits [MNW04]. Bivariate [MNW96, Rem12, She08, CG04, CGR13, DL98, DS00, DZ04, Flo96b, LS98, LKW17, MH17, NZ04]. Black [PGB15]. blind [Fli18, HR13]. block [AB97, Bai99, sCLC13, CSWP99, HH16, NBL11, TZL16, dCB07]. block-factorized [NBL11]. blossoming [Maz98]. Blossoms [Maz04, Maz99, Maz07]. body [GM14, HKRS14]. Boltzmann [HXC10]. Boor [JS95, Maz07]. both [ZYY16, ZD17]. bound [Bar03]. boundaries [BK96, She13]. Boundary [IMS99, MS99, ABI15, Boc04, CDS02, CWLH95, CP99, CWX12, CT17a, CHS17, CG17, Del17, FLM96, FK98, FZ12, FY13, FH18, GMM09, GM14, Gil17, yGmW98, GH06, GPG15, HH15, HJH12, HR02, HZHL12, JSSE97, Jeo98, JM03, JS05, Jia09a, KP95, Kno09, KS17, Kun95, Lee03, MR96, MB96, Ock96, QC12, RZ98, SW98, TZ07, TSZ98, TGA96, VC00]. bounded [BBR02, CL16, GM17, PST95, San15, TDG16]. Bounds [AS08, Dam07, LB93, Mai05, NG99, NSW16, XHC15, ZXC12]. Boussinesq [Lee03]. box [BS05, CGR13]. box-spline [BS05]. Bracket [Chi03]. branch [IB17]. branched [KC16]. breaking [GH08]. Brunel [New96]. Burridge [XKY15]. BVMs [AB97].

BVODEs [Mui99].

cache [CM99]. cache- [CM99]. Cahn [RWT13]. Calculation [VC00]. calculations [SL15]. camera [FGS13]. can [WS01]. cap [HW12]. capabilities [CLLS06, CLM96b]. capacitance [TW98]. cardiac [GLS15]. Cardinal [HL16a, BS05, Buh06, Led15, RS99]. cascade [CX04, Li04]. case [CM03, Han05, RH12]. Cauchy [Bra01, BR07, CKP13, Cum96, WZ10, ZW12]. Cauchy-type [Cum96]. Causality [GGL07]. cavities [QC12, VW02]. Cayley [DLP98]. cell [ABD08, BEPS96, Dad17, JS05, KHM16, ZQ13, ZG16]. cell-average [ABD08]. cell-centered [KHM16]. center [EHV17]. centered [BEPS96, KHM16]. central [TQ17]. centroids [MFB^{+}11]. ceramics [SEE96]. certain [Sto96]. certified [SFDE15]. CG [ZT14]. CG-like [ZT14]. chain [PR15]. chains [DMMS11, ZXC12]. challenges [Onc14]. channel [Jia00]. characteristic [CCE^{+}12]. Characterization [LLS95, Li04, SZ09, FGG^{+}17, Fus08a, Fus08b, GP95, LWY06, Mon09]. Characterizations [LY07, Far10]. Chebyshev [BR13, Cop98, MV99, Maz99, Maz01, Nee96, SC08, TZL16, WK93, Zhu15]. chemotaxis [CEHK18]. chiral [ZGGW12]. choice [GPT17, JJLR14, Kno09, LLY12, RH12]. Ciarlet [CL07]. circle [CJ07, GVSLN96]. circle-preserving [CJ07]. circuit [SS15]. circular [HO17, MKS02, PW08]. circulation [SI15]. class [ADL11, ABY13, Atr12, Bai99, BHT11, DG13, GMS12, Hak00, HR17, Hub12, LLS94, LSXZ15, MSS02, MN00]. classes [CKP13, Dai16, KL07, Pei05, Peñ11]. classical [Zoz15a]. Classification [CG04, XP10]. classifiers [ZJ06]. Clenshaw [XHC15]. Clifford [CLM02]. Closed [Hub12, Noa96]. Clough [LL98]. co [PR15]. coarse [KXZ04]. coated [CH15]. cocoercive [nCV13]. codes [CHMR97, Hal95, Kir98]. coefficient [GS13, SJJL15]. coefficient-based [GS13]. coefficients [Ehr00, EGL13, Pet02, ZB12, dB00]. cohomology [PR15]. Collocation [AHHR16, BJX09, Cum96, FS98, Gu17, yGqW09, GW14, Han05, Hau97, HS13, HR02, HS16, LLS09, LK05, Liu99, MR96, Nai12, Ram95, RWT13, Sch09, XH13]. colonoscopy [MU14]. combination [FPT06]. combinations [DGMES13, MPS96, MPS97]. Combined [DSS09, BBB^{+}16, HO15, HY18]. common [Len94]. commuting [GLT93]. compact [Gna07, Str13]. Compactly [CHSS03, JJK13, JS99, Wu95, CL16, Hub12, Lai06, LLS95, Len96, Wen95]. comparing [Sch15]. comparison [BBB^{+}16, Zoz15a]. compass [GMB97]. compatible [TDG16]. competition [Ble07, DLK12]. complements [Dai16, Peñ11]. completely [Dou94, Xu14]. completion [CMX07, WLW16]. completions [MRS14, MRS18]. complex [DOZ94, FMR00, FH05, Han10, IL16, MKS02, PC17, SM99]. Complexity [NUWZ17, Tom96, Cho95, Pei05]. components [RST^{+}14]. Composite [yGjW10, HXC10, Pla12]. comprehensive [FGG^{+}17]. compressible [BBB^{+}16]. compressing [MST05]. compression [DPS93, DKLT93, LW14]. compressive [SL12]. compromise [BCE^{+}09]. Comput [Fus08a, Hem96]. Computation [Ano02c, PR15, ZA10, ANSZ17, BPJ02, BH02, BPK96, CX04, ER99, GH99, LD07, MFB^{+}11, NP18]. Computational [Ano01b, Ano03c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano16, Pei05, Sau95, Yoo01, dB00, BOP^{+}15, Lee03, Sch15, NW04]. Computationally [TL18]. computations [ASU17, KP98]. computers [EJ97, vdH93]. Computing [CKP13, GH08, HBM03, Jon17, KL15, BI11, Kir98, LMV11, MXY13, SC08]. concentrated [GG06, RST^{+}14]. Concentration [GZ13]. concepts [WLPV15]. condensed [ASS10]. condition [Awa13, Flo94, QC12, Sch95, SS96, WL05, ZYY16]. Conditional [Xia13]. conditioned [ČF11]. conditioning [Ant18, LG13]. conditions [CDP09, Del17, EHV17, FLM96, FS07, Füh16, GW17, Gil17, HXC10, JM03, Jia09a, KK99, Kun95, MB96, TGA96, ZhXpZ16]. conduction [BDMR10, CN96, HT11]. Cone [Leo02]. configurations [BEG17, CG04]. confluent [DGS18, NP18]. conformal [CWLH95, CL18, HL16b, MKS02, TB16]. Conforming [MN16, ZC14]. conic [Flo96a]. conical [EHS02, PW08]. conjunction [DSS09]. connected [CWLH95, CL18, HL16b, San15, TB16]. connection [BGM93, NSW07, SM00]. connectivity [PŠWX03]. conservation [KK05, MV01, ZQ13]. conserving [ABI15]. consistency [CX06, HFH^{+}07, MNPR06]. consistent [KHM16]. constant [KPY09, LY08, ZB12]. constrained [BHT11, IUV17, KPY09, SW06]. constraint [KXZ04]. constraints [WOBL17]. Constructing [CGM01, FMPS13, GW17, GU02, Lem09, GM14]. Construction [AF96, BD10, ČF11, CHR00, CaL06, DHO16, FOW14, GS10, JMO00, Lai06, OBS15, Rei93, SX07, WOBL17, ABC14, BCZ05, CCH^{+}11, CP14, CMX07, CL02, FHK06, Hua06, KKLY10, LH13, MSS02, PST95, SW98, Str13]. constructions [CHL17]. constructive [CGR13, DG16]. contact [HKRS14]. context [HSV09, Wie15]. continuation [Awa15, Mur08, She13]. continuity [LM04]. Continuous [CS94, DM93, ALZ02, BP93, DY10, DAP13, JL97, Len96, Mui99, WY17, ZT14]. contraction [PR15]. Control [KLNS95, AV15, CBK01, Cas07, CCE^{+}12, GN13, HR01, Lee03, LY01, MRS03, Wil98, Yan03, dLdDSM07]. controls [MS16]. convection [CCE^{+}12, HO17, KY01, KXZ04, LRS12, Mat09, OS09, PP97]. convection-dominated [CCE^{+}12, KY01]. converge [CSW14, Gon93]. Convergence [AK15, BQRB13, CCSS09, FS98, Gom95, HX02, KB15, Li08a, LG13, MNC16, Noe95, PK94, Ryn00, WYW11, WZ10, Yao16, Yin07, CLC16, CDP09, Che06, CKW10, FJ99, FP99, FL00, GH06, Led15, LR07, Li04, Li05b, Li08b, PS98, Tem02]. Convergent [Lóp18, KXZ04, LLS09, WMB13]. conversion [LPY10]. Convex [FPT06, sCLC13, CZS16, CCH^{+}15, FHK06, FK10, HH16, Kva14, LD07, LY01, MR10, MRS03, Pet96, RGB13, War96, WSHD07]. Convexity [CC94, KK99, SS14, Car95, Flo94, LPSSP00]. convolution [Cop98, HJJV97, Li03, YZ13]. convolution-type [YZ13]. Convolutions [CK03b]. Coorbit [DST04]. coordinate [YWC12, Zwi94]. coordinates [FHK06, KB15, RGB13, War96, WSHD07]. corner [Noa98, Sab04]. corner-cutting [Noa98]. correction [KCW17, XH13]. corrections [FXZ96]. corrector [CSWP99, ZD14]. Corrigendum [Hem96]. corrupted [XWL13]. Coulomb [HL08]. Counterexamples [LL07a]. counting [Kot08]. coupled [GLO10, HFH^{+}07, JRS15, MRH15, MS98, OS09]. coupling [ADG17, DGMM16, GMM09]. Courant [CBK01]. Crank [How96, New96, Ock96]. creates [Ito96]. criss [SXYY06]. criss-cross [SXYY06]. criteria [ZV08]. criterion [FBCR13, Tem02]. Critical [Sto96]. cross [HL96, SXYY06]. cross-validation [HL96]. Crouch [JMO00]. cryptographic [Ren09]. crystal [HY14, RH15, Zhu15]. crystals [KK02]. CSOR [HJJV97]. cubature [DG13, GS10, IMS99, Pet00, Xu98, Xu00]. cubic [ANSZ17, CG06, ČF11, Flo96a, JL06, NSZ04, Pet96, Rem12]. cubical [PR15]. cubics [Noa06, PFMS05]. cuboid [HSY16]. CUR [VM17]. curl [BKK17]. Current [MM99, BMSR^{+}16, Cai02]. Curtis [XHC15]. Curvature [RS01, CC94]. Curve [Pot95, CM03, MU14, Pet96]. curved [KB15]. Curves [MP04, BX94, BWX97, BGN17, BC94, CLM02, CGM01, Der04, FS94, FaKS02, Far10, FGS16, FGG^{+}17, Flo96b, HBMY14, JKK^{+}12, KLNS95, KKV15, Mon09, Pot95, PS98, RSA14, Sab04]. curvilinear [LQ14]. cuts [DT12]. cutting [Noa98, Sab04]. Cycloidal [MP04]. cylinder [MKS02]. cylindrical [BGMS07, YDF97].

D [Hem96, HS13, Lee03, CK11, GT14, Han05, NV10, TZ07, Boc04, DL98, FF05, GM14, GL13, Hem95, MU14, VW02, vdHS00, JRS15]. D-problems [Hem96, Hem95]. damped [ADG17]. Darcy [CHM18, MRH15]. Data [CP04, HO15, LW14, Bes03, BE00, CCSS09, CP15, CGW13, DZ04, DSW05, DSS09, FHN94, FM04, Grä12, GL13, KMO^{+}14, Kun09, LW04, LNW02, LP14, MST05, MNW99, NSW98, NG99, Pot04, RS99, WL05, WSM13, Yoo01]. Data-driven [HO15]. data-independent [DSW05]. Daubechies [DKB99, EGL13]. Davidson [TZL16]. deblurring [CZS16, KPY09]. decay [NV10]. decaying [LP10]. December [Ano16]. decomposition [CM99, DT12, FKS05, GL13, Hag98, HKRS14, HFH^{+}07, HS03, Hu07, Las16, LH13, LHY08, PS05, PSNP11, SI15, Sto17]. Decompositions [DS13, AL13, CCN^{+}16, CHRX06, VM17]. deconvolution [Fli18, HR13]. decoupled [CHM18]. Deficits [BCHL03]. defined [FMR00]. definite [Bai99, GS10, LS05, Pin04, SM99, WMY13, WMB13, Wen95, Wu95]. definiteness [Luo99]. degenerate [Gna07, LB93]. degeneration [Bes03]. Degree [PR99, ABC14, BR13, DOZ94, FL00, Mar95, Mon09, PK94, SM00, Wen95]. degree-raising [FL00]. delay [BP93, BPW95, ER99, GH08, WY17]. delay-integro-differential [WY17]. delays [Dad17, XH13]. delta [CDS02, DRS08, LPY10, Wan08]. demixing [Fli18]. Democracy [GHdN12]. denoising [BCE^{+}09, JZ10, MSXZ13]. densely [She09]. densities [DMT03]. density [JJK13]. dependent [LZZ09, SW10, TB16, TGA96, dFGAJN18]. derangements [IS13]. Derivative [FXZ96, FF05, FW15, KS17, MB96]. derivatives [BRSV15, XL10]. design [Pot95]. designing [FGS13]. designs [HL08, Hes09]. desired [Lem09]. Detecting [BLP02, DGMES13]. detection [CGS93, CGW13, ER99, MP00]. determinants [Maz98]. determined [BGM93]. Determining [SF14]. deterministic [WSM13]. detonation [ZG16]. Development [CDP03]. Diagonal [Peñ11, Dai16, FMPS13, Kos15]. diagonalized [LWL17]. Diagonally [vdHS00, EJ97]. dictionary [Mai06]. difference [BEPS96, Che06, CLL08, CEHK18, JRS15, KY01, MB96, NV10, PL05, TZ07, Xu14, ZD16]. different [LQ14]. Differentiability [Sau06]. Differential [KK02, BP93, BPW95, ByLl01, CJX17, CWX06, CN96, DLP98, ER99, FZ12, Fas99b, FY13, GH08, yGS08, yGqW09, Hau97, HBM03, HS08, JLZ08, KA97, LMV11, Liu99, PM17, SW96, She13, lWyG06, WY17, Wat96, ZV08, ZD14, vdHv97]. differential-algebraic [LMV11]. differentiation [WH07, XL10]. Diffraction [BBR02, EHS02]. diffusion [BCE^{+}09, CK11, CLM96a, CLZW13, CHS17, CCE^{+}12, GLO10, HO17, KY01, KXZ04, Kim06, KHM16, KC16, Kun01, LD16, LRS12, MR15, Mat09, MXO13, MNC16, OS09, PP97, TDG16, ZD17, ZC14, ZW12]. digital [LPY10]. digital-to-analog [LPY10]. dilation [Gro13, Han98, Han06]. dilations [CHSS03]. Dimension [CL12, TQ17, KORS17, Mai06, PŠWX03]. Dimension-by-dimension [TQ17]. dimensional [BNR00, BR15, CMP07b, CGH^{+}13, CCH^{+}08, DLS14, HSY16, JN18, KF03, LG13, LZZ09, LL07b, MR15, Nov00, Onc14, SM00, WHS18, XKY15, ZQ13, ZG16]. Dimensions [ZWDD16, AS08, Che02, FP99, Füh16, HW00]. diminishing [CGP95]. DIMSEMs [EJ97]. DIMSEMs-diagonally [EJ97]. Dirac [CDS02, DRS08, yGSlX03]. Direct [AM14, LL99, Che02, DZ04, MS16]. direction [BS05, LLS96, LL98]. directional [CCH^{+}15, CGRS15, GL17, MSS02, NSS04]. directly [TQ17]. Dirichlet [AK01, ACH10, JM03, Lee03, Li05b, WL12]. Dirichlet-to-Neumann [WL12]. discontinuities [AL13, BR15]. Discontinuous [Ye04, AK15, BQRB13, KF03, Pet02, SXZ06]. Discrepancy [Leo13, Pla00, RJ00]. Discrete [CMX02, JSSE97, dLdDSM07, BF11, Ble07, CJX17, CKM99, Dad17, Kir98, LC13, LL13, MZ13, Mui99, Nee96, Rei93, TZ03, ZV08]. discrete-time [BF11]. discretisations [Mat09]. discretization [AS05, Bra01, FW96, HBMY14, HL06, HH15, JRS15, MV01, NV10, RSZ11, TZ07]. discretizations [Ang00, PGB15, XY15, XZ01, ZC14]. Discretized [HT11, CGS93]. discriminating [DGMES13]. disk [CWLH95, CL18]. dispersion [CLMR15]. dispersive [PM17]. dissipation [CLMR15]. distance [PZ16]. Distances [BWX97]. distinct [VK96]. distributed [FYL08, HH16, LY01]. distribution [ZD17]. distributions [CDS02, Li03]. div [BKK17, JJLR14, dFGAJN18]. divergence [Sor18, Urb95]. divergence-free [Urb95]. Divergent [Zoz15b, Zoz15a]. Dixon [Chi03]. Domain [DT12, AS05, BMSR^{+}16, CWLH95, DAP13, EHV17, GM17, Hag98, HKRS14, HFH^{+}07, HO17, HS03, Hu07, LHY08, PS05, PSNP11, Rem12, Sto17, SL05, VW02, ZhXpZ16]. domains [BK96, CLL08, CL16, GH06, yGS08, HL16b, IMS99, KB15, LG13, LHY08, Str13, TDG16]. dominance [Dai16, Kos15, Peñ11]. dominant [MNW04]. dominated [CCE^{+}12, KY01]. dot [LS05]. doubly [Tor16, TB16]. doubly-connected [TB16]. DOUG [Hag98]. Douglas [HH16]. Dr [Sil99]. driven [HO15]. dual [Boc04, CP03, zCpST09, CKK12, CKK16, Chr15, DM15, FGO14, FOW14, JJK13, Jüt98, KPY09, LPY10, Lem09, LKW17, LYY13, nCV13]. Duality [Noa06]. duals [CG10, CSS98]. Dyadic [DM99, CS94]. Dynamic [dCB07]. dynamical [Tan17, WOBL17]. Dynamics [Dad17, Kim06].

earthquakes [XKY15]. eddy [BMSR^{+}16]. Edge [CGW13, PGB10, GL17]. edges [Var96]. Editor [Ano95, MM99]. Editorial [GS12, MM93]. editors [Ano05a, Ano05b, Ano03c, Ano04d, Ano04e, Ano04f]. effective [HKR04]. efficiency [DL04]. Efficient [JGW15, MV01, PS07, VM17, AZW15, CM99, CCE^{+}12, DGMES13, HFH15, IP15, LD07, SW98, TL18]. eigenfunctions [SS16]. eigenproblem [CR99]. eigenproblems [JM00]. Eigenvalue [YLBL12, BPK96, CL07, EJ97, Gna07, HR01, Kos15, LLZ10, MSZ06, Mar94, RSZ11, TZL16, TBD94, XY15]. Eigenvalues [CGS93, Jon17]. eigenvector [BPK96]. eigenvector/eigenvalue [BPK96]. eighth [TBD94]. eighth- [TBD94]. EIM [RH15]. Elastic [BC94, Ble07, BGN17, FE15]. elasticity [AAQ15, Awa13, BQRB13, LLS03, Lee16, MC08, PGM09]. elastomers [Zhu15]. elastostatic [GG16]. electric [DD94]. electrical [HLN12]. electrically [SFDE15]. Electromagnetic [CH15, BH02, Cai02, VW05]. Electromagnetics [Ano02c]. electron [SL15]. electrophysiology [GLS15]. electrostatic [Tor16]. element [AK01, ARPR01, AB11, APV14, AL97, BNP14, BK96, CLLS06, CCG10, CZZ17, CBK01, CDS02, CLZ02, CLZW13, CT17a, CKW10, CG17, DY01, DLZ16, EHS02, FF95, FYL08, FH18, GMM09, GH06, HR01, HX02, Jan98, Jeo98, JM03, JS05, JLZ08, JK08, KORS17, KL00, LL04, LLS03, LB93, LLC16, LL10, LL12b, MR15, MN16, MSZ06, MC08, MS99, MY17, NBL11, PP97, SW96, SXYY06, SW98, VW05, WY16, WLPV15, Xie08, XY15, XZ01, Ye00, ZP06, ZSB01a, Zha14, ZPY15, ZT15, ZSHZ15]. elementary [FH05, Lóp18]. elements [AS02, ASS10, AAQ15, Awa13, BKK17, Bra01, FF95, HS03, HSY16, HY18, LLS94, LLS96, fLL05, LL07a, LPSSP00, SL15, YDF97, YLBL12, Ye04, dFGAJN18]. elevation [PK94]. elimination [WW12]. elliptic [AC05, ACH10, AHC11, CDP03, Cas07, CWX12, CP15, DFR07, FK98, FZ12, FYL08, GM14, GH06, HR01, Hol01, Jan98, Li05a, LMT01, OP11, Pet02, PSNP11, SXZ06, WHS18, XZ03, Yan01, ZB12, Zha14, ZL16]. embedded [BR15, She09]. embeddings [GHdN12]. empirical [CDP09, CHRX06, MNPR06]. Energetic [ADG17]. Energy [ABI15, Ble07, HL08, Hes09, Leo13]. Energy-conserving [ABI15]. Enforcing [HSV09]. engineering [SEE96]. enriched [HT12]. Entropy [CDP09, Mai06]. environment [ZGGW12]. epidemic [Kim06]. Equal [CK03a, Möß10]. Equal-Norm [CK03a]. equation [ADG17, AC00, CLC16, CBK01, Che02, CHS17, CGS93, DD94, DLS14, FZ12, FYL08, GT14, Gil17, yGSlX03, GW14, Han05, HS13, HR02, HY18, HS16, Jeo98, JM03, KP95, KA97, KN14, KNQ16, Kun01, LM14, Li05b, Li08a, Li09, Li10a, LL07b, MS99, MS16, NV10, PM17, PGB15, PTC12, Ryn00, Sch09, SW96, Sto17, TZ07, TGA96, WL12, WZ10, Xu14, ZT14, ZSB01b, ZW12]. equations [AG98, AB02, Ang00, APV14, ACH10, AHC11, Awa15, Bai99, BNP14, BP93, BPW95, ByLl01, BF11, BK16, Boc04, Bra01, Bre95, CDP03, CCG10, CJX17, CMX97, CP99, CWX06, CLL08, CLZW13, CS94, CN96, Cum96, CT17b, DFR07, DM93, DPS93, DS01, DX10, ER99, FH05, Fas99b, FR10, GH99, GLO10, Gu17, GH08, yGS08, yGqW09, Han06, HBMY14, Hau97, HL06, HBM03, HS08, HSY16, HY18, Hua06, HZHL12, Jan98, JJLR14, JSSE97, JLZ08, JK08, JRS15, KNP98, KY01, KL17, KHM16, KC16, LN14, Lee03, Li04, Li08b, LC09, LC13, LD16, LZY98, LMV11, Liu99, LHY08, LY08, LLC16, LLY12, Lut04, MN16, MXO13, MY17, MS98, Nai12, Pen98, Pet02, Pla12, PS98, RZ98, She13, SS15, Sun14, gTpM02, TQ17, TSZ98, VW02, lWyG06, WY16, WY17, XH13, Xie08, XZ03, YZ13, ZVW95, ZV08, ZT15]. equations [ZD14, ZhXpZ16, dH94, dFGAJN18, vdHv97]. equilibrium [sCLC13]. Equivalence [SC08]. Erasures [CK03a]. ergodic [ZXC12]. ERM [ZXC12]. Errata [Ano04c, Ano08]. Erratum [Ano16, Dũn12, Fus08a]. Error [Alz10, Ang00, Ano01b, CCG10, GRB12, yGW07, KF03, LB93, Sch95, YDF97, aKT17, AK01, ARPR01, Bar03, BGH15, Bra01, CLC16, CBK01, CK11, CLZW13, CCE^{+}12, DY01, DLZ16, GN13, HR01, HWZ13, JK08, KY01, KK05, KL17, KNV01, Kun01, LMV11, LMT01, LY01, LL10, LL12b, MSZ06, MRH15, MW01, NSW16, Pet00, Pet02, PS05, RGB13, Sch15, TZ03, War13, WSM13, XHC15, Xu14, Yan01, Yan03, Zha01, ZSB01a, ZC14]. error-optimal [Sch15]. errors [CLMR15, DM15, Nes16, Wan08]. ESDIRK [BBB^{+}16]. estimate [CK11, CLZ02, DY01, LL10]. estimates [ARPR01, AL97, CLC16, CLZW13, DG16, FS98, Füh08, Fus08a, Fus08b, GRB12, Gom95, GZ13, GN13, KY01, KL17, KNV01, LMT01, LY01, Low05, LL12b, MSZ06, PS05, PR99, RGB13, Sch95, Spe17, War13, WSM13, Zha01, ZC14]. Estimation [Ano01b, BL09, Bra01, Ehr00, KF03, Kun01, MRH15, MW01, WK93, aKT17]. estimators [AK01, Pet02, Yan01, Yan03, ZSB01b, ZSB01a]. Euclidean [Bün11, Kun09, NSW98]. Euler [AK15, CLC16, KvL95, XKY15]. evaluating [WZtX16]. Evaluation [CWLH95, GP13, Bar03, CN96, DP13, GG16, IP15, Joh13, Pan95, aKT17]. evolution [GH99, MU14]. evolutionary [BPW95, Xu14]. Ewald [Tor16]. exact [SLN14, ZV08]. exactness [HW12]. example [PW94]. Excesses [BCHL03]. Existence [MSS02, CCH^{+}11, Far10, LM06]. exotic [lZmCRX03]. expansion [CL07, HLN12, LL07a, NP18, aKT17]. expansions [Atr12, CI13, Dam07, DHO16, DGS18, HWZ13, JLZ08, LT13, Lóp18, Tem07, WZL17, vdMNS03]. experimental [CBK01]. Experiments [Wil98, KMO^{+}14]. Explicit [BCZ05, CHL17, CHMR97, CMR07, CLMR10, Hal95, HH15, Nie98, PIZ01, ZP06]. exploitation [AG98]. exploiting [AD08]. Exploring [CDTV99]. exponent [Pla00]. Exponential [CM12, JKLY13, ZPR03, LM06, NV10, WZtX16, ZL17]. exponentials [CCS16, JGW15]. extension [DZ04]. Extensions [HR13, CKK16]. exterior [GMM09, yGS08, yGjW10, LST11, LG13]. extraordinary [Pra98]. Extrapolation [LLZ10, fLL05, Xie08, BRZ94, CL07, HZHL12, JLZ08, RZ98, Sid17]. Extremal [SW04].

factor [Han98]. Factorization [CCS16, AH96, GMRS97, Sau06]. factorized [NBL11]. factors [MN15]. fairness [GU02, KP04]. Families [MH17, CL02, Gro13, Led15]. family [DLUS17, Wat94, Xu98]. Fan [Hof06]. far [SFDE15]. Fast [AZW15, BCM99, IP15, KP98, KN14, MFB^{+}11, AC00, BMB14, BA15, Bra01, Cai02, CJX17, Che02, CLL08, DPS93, DLK12, GG16, Han05, JZ10, Joh13, LT98, LLY12, SFDE15, YZ13, SF14, Ste98]. Fatou [Cho95]. feasibility [sCLC13]. feedforward [Mha93]. FEM [ADG17, MW01, MS98, Yan03, ZSB01b]. FFT [DG16]. FFTs [HKR04, Nes16]. field [DD94, SFDE15]. fields [BH02, GG16, Sor18, TS05]. Filon [XHC15]. filter [CMX07, CW09, EGL13, Jia09b, TL18]. filters [CL02, Jia09b, MWW13]. financial [ALZ02]. finding [Yao16]. Fine [JN18]. Finite [BF03, CG17, EHS02, JK08, LMWX13, NBL11, VW05, AK01, ARPR01, AB11, APV14, AL97, AAQ15, BNP14, BKK17, BEPS96, BK96, Bra01, CL12, CDP03, CLLS06, CCG10, CZZ17, CBK01, CCN^{+}16, CLZ02, Che06, CLL08, CWX12, CGH^{+}13, CLZW13, CEHK18, CKW10, DY01, DLZ16, FF95, FWW06, FYL08, GMM09, Grö03, Hem95, Hem96, HR01, HX02, Jan98, JLZ08, JRS15, KY01, KORS17, KH05, KL00, KHM16, LLS94, LLS96, LLS03, Leo13, LC09, LC13, LL07a, LB93, LLC16, LPSSP00, LL10, LL12b, MR15, MN16, MSZ06, MC08, MY17, NV10, Noe95, PL05, PP97, SL15, SW96, SXYY06, SRFH12, Sun08, TZ05, TZ07, WY16, WZL17, WLPV15, Xie08, XY15, XZ01, YDF97, Ye00, ZSB01a, Zha14, ZPY15, ZT15, ZD16, ZSHZ15, dFGAJN18]. finite-difference [JRS15, KY01]. Finite-element [NBL11]. finite-part [WZL17]. finite-volume [Hem95, Hem96]. finite-volume-finite-difference [CEHK18]. finitely [MNW04]. fire [FPR^{+}12]. first [CLMR10, HR02, HZHL12, Nai12, Pla12]. fitted [KL16, ZL16]. Fitting [JF02, BCM99, DZ04, DSS09, WK93]. Fix [GW17, WL05]. Fixed [CZS16, LSXZ15]. fixed-point [LSXZ15]. flat [SRFH12]. flatness [GL17]. Flexible [HL05]. flights [ZD17]. flow [CLL08, CT17a, HFH15, KH05, LG13, lXyG02, Ye04, ZPY15]. flows [BBB^{+}16, CD15, KvL95, PIZ01, PCM^{+}96, RHZ17, TL18]. fluids [DGMM16]. flux [DLZ16]. focal [PW08]. fold [Jia11]. FOM [Cul96]. Foreword [Buh99, HN96, How96]. form [AM14, FGS16, Hub12, LL99, LKW17, Rei93]. Formal [Sal98]. formed [Mai06, SEE96]. forms [Mar95, SM00]. formula [CG15, Fra99, SS13]. formulae [CG04, GS10, MC10, Xu98, Xu00]. formulas [BB07, FXZ96, Jeo98, SX07, Wal95]. formulation [CL18, DGMM16, HJH12, MS16]. formulations [TW98]. Fortran [Kir98]. four [GMS12, NSS04]. four-directional [NSS04]. four-point [GMS12]. Fourier [SF14, Ste98, CGW13, FQ10, LL07b, LP14, QW11, SL05, SS13, TZ03]. fourth [Li05a]. fourth-order [Li05a]. fractional [GM17, HX17, KS17, LD16, MNC16, PM17, SU12, TDG16, ZD14, ZD16, ZW12]. Frame [CR08, MRS18, CP14, DFR07, Jia09b, KKL03, LL11, MR10, MRS14, PŠWX03]. framelet [CCSS09]. framelets [Lem09]. Frames [BCHL03, BF03, CK03a, CC03, GH03, HM03, HLT11, Jia03, Sun08, BS06, BH15, CL12, CCH^{+}11, CK07, CCN^{+}16, CKK16, Chr15, CSS98, CHSS03, DST04, DM15, Far10, FGG^{+}17, FS07, FOW14, FWW06, Füh08, Grö03, JJK13, LPY10, LMO14, LH13, LL13, MNPW00, Onc14, Søn07, Zha05, CS03]. framework [ABC14, DAP13, WZtX16].

Franklin [NP18]. Fredholm [FR10]. Free [KL04, PCM^{+}96, CG17, DAP13, JW95, LLS03, Ock96, Urb95, ZL16]. Free-knot [KL04]. Freeknot [KL07]. frequency [GG06, KCW17, LHY08, Onc14, RST^{+}14, SF14]. frequency-concentrated [GG06]. friction [HKRS14, XKY15]. Friedman [NP18]. Froissart [BC99]. Full [DS01, ASS10]. fully [LWL17]. Function [HM03, Alz10, AM14, Boc04, BCZ05, CTZ02, CGH^{+}96, DSW05, EGST13, FF05, IB17, JW95, Kot08, LW14, LLY06, LL99, LT13, Mac94, Mai10, MNW99, MP00, NG99, Rip99, Sch95]. functional [Liu99, Sun14, ZD17]. functional-differential [Liu99]. functionals [MR10, SM00]. Functions [Zoz15b, BCM99, BGH15, Bes03, BMB14, BW15, Cai02, Car95, Cas07, CP03, CX04, CSW14, CL16, CTZ02, CG10, Chr15, CSW96, De 06, DG16, DX10, DGS18, Fas99a, Fas99b, FPR^{+}12, FHN94, FS98, Füh08, FW15, GP13, GHdN12, GG06, GGL07, Gro13, GH16, GS10, yGSlX03, GYZ17, Han98, Hub12, JS99, Jüt98, KKLY10, KL04, KL07, Le 05, LST11, LM14, Len96, Li04, Li05a, LK05, LWL17, Lóp18, Low05, LP14, MM00, MST05, MNW04, MS04, NP18, NSW16, PŠWX03, Pet00, Pin04, RZ10, RS99, RY13, Sau06, SW06, Sch11a, SS08, SRFH12, Sto96, SM99, TZ05, War13, WH07, Wen95, Wu95, WFL02, WZtX16, Xu98, Zoz15a]. fundamental [AC05, Ant18, FK98, LL98, Li05b, Li08a, Li08b, WZ10]. furnaces [BGMS07]. fusion [CCH^{+}11, CP14, LH13]. fuzzy [DLK12].

G [DDP14]. G-symplectic [DDP14].

Gabor [CL12, CC03, CKK12, Chr15, DM15, FS07, FGO14, Füh08, JJK13, LL13, LL11, LW03, RST^{+}14, Søn07]. Galerkin [AK15, AS05, BQRB13, CMX97, CMX02, CP15, CT17a, Cop98, CCE^{+}12, DHK98, DS01, FLM96, Gom95, HL06, Hua06, KNP98, KF03, LN14, LLC16, OBS15, Ryn00, SXZ06, WY16, WY17, ZB12]. Galerkin-characteristic [CCE^{+}12]. Gasca [CP07a]. Gauss [Dou94, yGqW09, LT13, SC08, WZL17]. Gaussian [CC94, CSW14, MC10, RS99, SX07, SHTS14]. Gaussians [Xia13, YZ08]. Gegenbauer [ESM13]. general [Alb15, ABC14, BCZ05, BRZ94, FHK06, HL16a, Han06, Kos15, RL07]. Generalization [NG99, ZXC12, CGP95, CK07, CKBP11, JS95, MNPR06]. Generalized [BRS13, Dar03, SL05, WL05, gYyG12, Zoz15b, AB11, CLC16, CG15, sCLC13, CHL17, DP06, GW17, GP13, GRB12, GH99, GW14, IS13, LWL17, MN00, Pen98, SF14, Zoz15a]. generalized-functions [Zoz15a]. generally [Gon93]. generated [CG06, Chr15, CSW96, JS99]. generating [CK07, NSS04]. Generation [FWW06, SR02, CGR11, Cop98]. Generators [Jia03]. geodesic [BMK15, SLN14]. geodesics [SLN14]. Geometric [GPG15, Mai10, MO16, PFMS05, BS06, BI11, BWX97, FH05, JKK^{+}12, LM04, Leo02, LO08]. Geometrical [Ano02d]. geometrically [ZXC12]. geometry [PW08, PGM09, PGB10, RAB^{+}09]. Gersgorin [Hof06, CKBP11]. Gibbs [RY13]. Ginzburg [HX02, LL07b]. Givens [CM99]. Global [Kim06, LSW17, Len94]. globally [BF94]. GMRES [BCM99, Cul96, MS98, Smo07, dH94, dCB07]. GMRES/FOM [Cul96]. good [Rip99, Sch11b, WS01]. Gordon [GW14]. grad [JJLR14, dFGAJN18]. Gradient [DY01, BMK15, DSS09, Wan12, Yan01, Yan03, Zha01]. gradients [YWC12]. Graduated [BCE^{+}09]. Gram [CK07]. graph [DGMES13, DT12]. graphs [KK02]. grating [BBR02, LN14]. Greedy [GN01, Tem01, Tem07, TYY11, DT96, GN08, LT06, LT12, QW11, Tem98, Tem00, Tem02]. green [FY13, Cai02]. Gregory [Ano93]. Greville [CHY11]. grid [BNP14, BMK15, Che06, DG16, Hem95, KHM16, MR15, PGB15, XZ01, Hem96]. grids [ANSZ17, BNR00, FF05, Hag98, KXZ04, LQ14, Noe95, Pla00, PS05, SXYY06, Ste98, Wie97, Zha01]. Grossman [JMO00]. growth [RH15, RS99]. Guaranteed [HFH^{+}07]. Gudkov [Hof06]. Guest [Ano95]. guided [BPJ02]. Gupta [Ock96].

h [WY17]. Haar [ZhXpZ16]. habil [Sil99]. Hadamard [WZL17]. half [GG16, Li94, LWL17, lWyG06, gYyG12]. half-plane [Li94]. half-space [GG16]. halfband [MWW13]. Hamilton [TQ17]. Hamiltonian [ABI15, DDP14, JM00, Mar94, MO16]. Hammerstein [KNP98]. hand [FR10]. Hankel [AM14, Li94]. Hanrahan [Han05]. hardening [GN13]. Hardy [CS06]. harmonic [HS13, HY18, KP98, LHY08, Sch09, WZtX16, Xu04a]. harmonics [FM14]. HCT [LLS94]. HDG [MNC16]. heat [BDMR10, CN96, HT11, HXC10, JGW15, LM14, MS16, PCM^{+}96, TGA96]. heat-conduction [CN96]. heating [BGMS07]. Heaviside [RY13]. helical [Mon09]. Helix [PW94]. Helmholtz [DRS08, DLS14, GMM09, HH15, HJH12, HY18, Kir98, Li08b, Li09, Sto17]. hemivariational [Yao16]. Hermite [ABD08, ABY13, CCS16, DM99, FaKS02, Fas99a, Fra99, Füh08, yGSlX03, Hak00, JL06, KJ09, Luo99, MST05, MS12, Moo17, PFMS05, RSA14, SX95, TQ17]. Hermitian [WMY13, WMB13]. Hessenberg [EGST13]. heterogeneous [TS05]. Heuristic [MN15, DGMES13]. Hexagonal [Jia09b]. hidden [CLM96b]. Hierarchical [Spe17, Han05]. High [BNR00, CEHK18, HBMY14, LD16, NP18, Nov00, ZD16, ZZ09, CLM03, Han10, HKR04, HS03, JMO00, Jia09b, JN18, LL12a, MS04, TL18]. High-order [CEHK18, HBMY14, HKR04, JN18]. high-pass [Jia09b]. High-precision [NP18]. Higher [CWX12, MRS03, CDP03, Füh16, GW17, KL16, Mat09, Noe95, OBS15, WOBL17, Wie97]. Higher-order [CWX12, MRS03, GW17]. highly [DGMES13, XHC15]. Hilbert [BMB14, Chr15, Dam07, HT11, MXY13, RY16, SM99, Yao16, vdMNS03]. histogram [CP07b]. hodograph [AF96, FS94, FaKS02, FGS16, JKK^{+}12, KJ09, PFMS05, Pot95, RSA14]. holonomic [WOBL17]. homoclinic [CS93]. homogeneous [AK01, BC99, DST04, Jia09a, LR07, ZGGW12]. homogenized [Str95]. homological [PR15]. homology [PR15]. Homotopy [PM17, AF96]. Hood [fLL05]. Hopf [CGS93, ER99, Far10]. Householder [FWW06]. Hunting [CS93]. Huygens [IKS10]. Hybrid [CM96, LQ14, BPJ02, sCLC13, CEHK18]. hyper [AZW15]. hyper-reduction [AZW15]. Hyperbolic [Len96, CKM99, KK05, KF03, MV01, ZQ13]. hypergeometric [CDTV99, DGS18, LT13, NP18]. hyperinterpolation [Li08b]. hypersingular [DLS14, HWZ13, MS99].

ideal [She08]. idealized [Jia09b]. ideals [MS04, Xu00]. identical [XP10]. identification [BJX09, FYL08, SJJL15, Tan17]. identity [GMB97, TSY10]. II [CMX97, DPS93, HL06, Hof06, MPS97, SS13, zCFX07]. ill [Ant18, BL09, BE00, LLY12, XY12]. ill-posed [BL09, BE00, LLY12, XY12]. image [BCE^{+}09, CZS16, Del17, DLK12, JZ10, LSXZ15, MSXZ13, RWT13, SX09]. images [RY16, SF14]. imaging [BKN11, DGK15, KMO^{+}14]. immersed [CLLS06, CZZ17, CKW10, GPG15]. impedance [HLN12, QC12]. implementation [DHO16, Jan98, dH94]. Implicit [Wie15, XKY15, BB07, CLC16, CLM03, EJ97, GH08, GPG15, HBM03, Mui99, vdH93, vdHv97, vdHS00]. implicit-time [GPG15]. implicitly [FMR00]. implicitly-defined [FMR00]. imposed [Awa13]. Improved [Fus08a, Fus08b, KL17, FJ99, Sto17]. inclusions [nCV13]. incompressible [Ang00, BQRB13, CLL08, HFH15, KH05, lXyG02, Ye04, ZVW95]. increasing [CSW14]. increasingly [SRFH12]. independence [Ito96]. independent [CTZ02, DS00, DSW05]. Index [Ano02a, Ano02b, Ano03a, Ano03b, Ano04a, Ano04b, Ano00a, Ano01a, Cum96]. indexing [dB00]. indicator [ZQ13, ZG16]. indicators [LQ14]. induced [CGR13]. induction [BGMS07]. Inequalities [GG06, Alz10, BR07, RZ10]. inequality [JRS15, LB93, Yao16]. inf [dFGAJN18]. infinite [BD10, Buh06, FF05, lXyG02, Zhe06]. infinitely [RZ10]. influence [EHV17]. initial [Alb15, vdH93]. initial-value [vdH93]. injectivity [FK10]. inner [Joh13, Pin04]. innovation [Sun08]. input [CP15]. insertion [BGM93, KLNS95, Wal95]. instability [TBD94]. integer [CHSS03]. integrability [RS01]. Integral [KNQ16, AG98, Boc04, CWLH95, CMX97, CP99, Che02, CHS17, Cum96, Dar03, DD94, FR10, Gil17, Gna07, Gu17, HBMY14, HJH12, HR02, Hua06, HZHL12, HS16, JSSE97, Kir98, KN14, LN14, Li09, LZY98, LLY12, MS99, Nai12, Pla12, RZ98, Ryn00, TW98, TSZ98, VC00, WL12, XH13, Xu14, YZ13]. Integrals [Zoz15b, CLMR10, CKP13, DL98, DG13, HWZ13, MV99, NUWZ17, VC00, WZL17, XHC15, Zoz15a]. integrate [FPR^{+}12]. integrated [BPJ02].

Integration [HW00, DDP14, DL98, ESM13, HW12, MO16, Nov00, NSW16, Pei05, SW04, ZB12]. integrators [KL16, OBS15, WOBL17, WZtX16, dLdDSM07]. integro [JLZ08, WY17, ZV08]. integro-differential [JLZ08]. interface [BK96, CZZ17, CKW10, DLZ16, FZ12, KP95, WHS18, ZL16]. interface-fitted [ZL16]. interfaces [BK96]. Interior [AL97, NV10]. interpolant [Dũn09, SW08]. interpolants [ABC14, ABY13, GMB97, MS07, Rem12, Spe17]. Interpolating [CT05, AF96, BL99, CGM01, DL98]. Interpolation [BLS02, CG06, DG13, GY99, Noa02, Pet96, RGB13, dB07, ABD08, AM14, BNR00, BS05, ByLl01, Buh06, Cal05, Car95, CG04, CS06, CM03, DSW05, DS10, DAP13, DM99, FaKS02, FGS13, Fas99a, FP99, FE15, FF05, Fra99, GS00, GRB12, GO05, Hak00, HL16a, JKK^{+}12, KJ09, Kun09, Kva14, LL98, LW04, LM04, Le 00, LLY06, LL99, LR07, LZY98, LL07a, Luo99, LNW02, MS00, NSW98, NZ04, NSZ04, Osw04, PFMS05, RZ10, RS99, Rip99, RSA14, Sau95, SX95, Sch95, Sch11b, SS96, SS15, SRFH12, Ven94, WS01, WL05, Xu00, Xu04b, dB00, dB06, vdMNS03, zCFX07]. interpolation-free [DAP13]. interpolative [VM17]. interpolators [Led15]. Interpolatory [Zhe06, CGR11, CGR13, DG16, DLUS17, JS99, KKLY10, ZA10]. intersecting [LD07]. intersection [BX94]. intersections [HFH^{+}07]. Intertwining [Cal05]. interval [ČF11, JL06, Jia09a, MS99, PST95, gTpM02]. intervals [LMV11]. Introduction [AX98, Bar94, HAS05, Mar95]. invariance [CL12]. Invariant [MT98, CHL17, FaKS02, GU02, HHS04, Mai06, MXZ08, NSW16, Sun10, Wat94, vdMNS03]. invariants [DGMES13]. inverse [BDMR10, CR99, CH15, CHY10, HO15, HXC10, IKS10, QC12, RSZ11, San15, WL12]. inverses [NBL11, SU12]. Inversion [SS13, GPT17, Li03, LL12a]. Inverted [BKK17]. involving [Sto96, Tom96, nCV13]. Irrational [yGS08]. Irregular [CC03, CLL08, FS07, Noe95, Pla95, Zha01, ZhXpZ16]. isometry [SL12]. isomorphism [DGMES13]. isospectral [PIZ01]. isotropic [BCE^{+}09, PGM09, RAB^{+}09]. issue [Ano16, BOP^{+}15, CHY10, XY12]. Issues [BPW95, Cai02]. iterated [BS97b, BS97a, BS00, CTZ02]. iteration [BCM99, RSZ11]. Iterative [Cul96, BBB^{+}16, FP99, GO95, IKS10, LNW02, WW12, WMY13, ZT14, dH94, dCB07].

Jacobi

[ByLl01, CR99, DHO16, TQ17, gYyG12]. Jacobi-type [DHO16]. Jacobian [ZD14]. Jacobian-predictor [ZD14]. John [Ano93, How96, New96]. joint [FRT09, Möß10]. Julia [Cho95]. jump [ZhXpZ16]. junction [WHS18]. Justen [HR13].

Kantorovich [BR07]. Kármán

[MN16, Mur08]. Kelvin [BNP14]. kernel [BJX09, DS10, DL04, FY13, Gna07, HH15, HS08, HT11, HJJV97, JSSE97, JGW15, KNP98, RY16, SS16, Xu14, YZ13, vdMNS03]. kernel-based [DS10, SS16]. kernel-split [HH15]. kernels [BRS13, BRSV15, Cop98, FJ99, FY13, GW17, HS13, LMWX13, LS05, Opf06, SW06, Sch09, SHTS14, ZJ06]. kind [CMX97, CP99, CHS17, HR02, HZHL12, HS16, LL04, Nai12, Pla12, TW98]. Kinematic [SR02]. Klein [GW14]. Knopoff [XKY15]. Knot [BGM93, KLNS95, KL04, LY07, Wal95]. knots [Plo95, Str95]. Kohn [CGH^{+}13, GP14]. Krylov [ZVW95]. KSR [BF94]. KSR-1 [BF94]. Kutta [CSWP99, BP93, BS97b, BS97a, BS00, CHMR97, CLM03, CMR07, CLMR10, CLMR15, Hal95, JMO00, Liu99, Mui99, vdHdS97, vdHS00].

Kutta-collocation [Liu99].

L1 [MSXZ13]. L1/TV [MSXZ13]. lacunary [Ven94]. lag [BP93]. Lagrange [Cal05, CG04, Hu07, Kun95, LL98, NZ04, NSZ04, WLW16]. Lagrangian [XWL13]. Laguerre [DGS18, yGjW10, LWL17, PGM09, PGB10, lWyG06, lXyG02]. Lambert [CGH^{+}96, IB17]. Landau [HX02, LL07b]. Laplace [BW15, HS13, KP95, KN14, LM14, LL12a, PTC12, Sch09, WZ10]. Laplacian [Jon17, SU12, TDG16]. Laplacians [CW16]. large [BF11, SI15, SFDE15, ZPY15]. large-scale [BF11]. laser [GN13]. lattice [FGO14, GH99, NSW16]. Lattices [CC03, CG06]. Laurent [GMRS97, Lev99, MS04]. Lauricella [BW15, TZ05]. laws [KK05, MV01, XKY15, ZQ13]. Lax [GLS15]. layer [CLM96b, IMS99, aKT17]. layered [Cai02, Wat96]. layers [Kno09, LZZ09, ZL17]. Learning [GS13, MNPR06, XP10, YZ08, YWC12, GZ13, GYZ17, LMWX13, LS05, RZ10, SW06]. Least [FHN94, JF02, WK93, CHY11, LK05, Pla00, Sto96, SW10, Tom96, ZPR03]. Least-Squares [JF02, LK05, Sto96]. Lebesgue [Las16]. Leer [CT17b]. Leer-type [CT17b]. Leffler [GP13]. Left [SU12]. Left-inverses [SU12]. Legendre [yGqW09, yGjW10, HKR04, gTpM02, lXyG02]. lemma [HHS04]. level [Bre95, DMMS11, HL06, HT12]. life [FRT09]. like [ZT14]. likelihood [BL09]. Limitations [CLM96b]. line [Dam07, GP13, LWL17, MV99, NUWZ17, VC00, lWyG06, gYyG12]. Linear [GPS10, HL96, AH96, ABY13, Bai99, BKN11, BP11, Cas07, CJ07, CD15, CL18, Dũn09, Han10, Ito96, KA97, KL17, KvD01, LLS03, LY07, LMV11, LY08, MC08, MPS96, MPS97, MO16, MV01, MS16, PGM09, PP97, Sch15, TZL16, VK96, WMB13, Wil98, dH94, dCB07, vdHdS97]. linear-constant [LY08]. linear-phase [Han10]. linearization [HX02]. linearizing [XWL13]. Linearly [SW06, CTZ02, DS00]. lines [BEG17]. link [Len94]. Liouville [GM17, KS17]. Lippmann [Che02]. Lipschitz [GH16]. liquid [HY14, Zhu15]. LLT [zCpST09]. loads [FE15, GM14]. Local [ANSZ17, FGS16, KK05, Mat09, Sun10, XZ01, ARPR01, AZW15, ABC14, APV14, AS05, BCE^{+}09, DL98, DZ04, GY00, Han12, HHS04, HT12, HSV09, HY18, LSW17, Len94, LWY06, MSZ06, MRS18, Spe17, SZ09]. Localization [CP03, Lem09, MSST14]. localizations [Kos15]. Localized [Grö03, Sun14, Rau05, RST^{+}14]. Locally [DS00, CTZ02, GH16, KC16]. locations [DSW05]. locking [LLS03]. locking-free [LLS03]. logarithmic [JSSE97, KNP98]. logarithmic-kernel [JSSE97]. long [LL07b, Xu14]. long-time [LL07b]. Lorentz [DKB99]. loss [GYZ17]. Lotka [SC08]. low [AL97, BH15, CLMR15, KvL95, LHY08]. low-frequency [LHY08]. low-redundancy [BH15]. low-subsonic [KvL95]. lower [Mai05]. lowest [ZT15]. Lyapunov [LMV11, Pen98].

machine [RZ10, SW06]. Machines

[EPP00, CX06]. macro

[AS02, ASS10, LL04]. macro-element [LL04]. macro-elements [AS02, ASS10]. magnetohydrodynamic [RHZ17]. MALDI [KMO^{+}14]. manifold [Mai10]. manifolds [GH16, Hol01, LR07, Mai06, Moo17, Noa14, YZ08]. many [MNW04]. map [Far10, San15, WL12]. mapping [CWLH95, EGST13, TB16]. mappings [FK10]. maps [FPT06, HR17]. marching [Awa15, DD94]. Mariano [CP07a]. market [ALZ02]. Markov [DMMS11, ZXC12]. Masks [Jia03, De 06, Gon93, Han10, LP10, Zhe06]. matched [LZZ09]. Material [XB05]. materials [HXC10]. Matérn [BRS13]. Math. [Fus08a, Hem96]. Mathematica [CKP13]. Mathematical [SEE96, CLM96a, SX09]. mathematics [BOP^{+}15, Ano16, Ano03c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b]. Matrices [BMSZ01, Chi03, AD08, AH96, BGM93, CR99, Dai16, EGL13, FMPS13, GP95, LH13, Möß10, Peñ11, Tom96]. Matrix [Low05, AC00, BD93, BI11, BR07, CMX07, Cop98, DY10, ESM13, EGST13, FKS05, FE15, Fus08a, Fus08b, GM17, Han06, HHS04, JL97, KA97, VM17, WLW16, DPS93]. Matrix-valued [Low05, Fus08a, Fus08b, HHS04]. Matzinger [DKB99]. Maximal [BS05]. maximizing [Bün11]. Maximum [CK11, CHSS03, KL17]. maximum-norm [KL17]. Maxwell [AB02, HY18, KCW17, LHY08, VW02, Xie08]. mean [FK10, RGB13]. means [BI11, Xu04a]. Measurable [GH03]. measure [CMP07a]. measurement [FZ12]. measurements [Zwi94]. measures [GU02, MSST14, MRS03, MT98]. measuring [DL04]. mechanical [HZHL12]. media [Cai02, DGMM16, TS05]. medial [CCH^{+}08]. median [Osw04]. median-interpolation [Osw04]. medians [GY99]. melting [PCM^{+}96]. memoriam [Ano93]. memory [IP15, Jan98]. Merrien [Sab04]. mesh [BS05, DM99, LLS96, LL98, MSS02, NSS04, ZL16]. meshes [AAQ15, DY01, FPT06, GH06, HSY16, KP98, Kun01, LL10, LL12b, Mat09, Pla95, PW08, PGB10, ZWDD16, ZL17, ZC14, ZD16]. Meshfree [TS05]. meshless [AS05, FS98, HS08, HS13, KH05, Sch09, Sch15, ZB12]. Metamorphosis [RY16]. method [AF96, ARPR01, AC05, AB11, Ant18, APV14, AL97, AC00, ACH10, AHC11, AS05, BQRB13, BK16, Boc04, BK96, BS97a, CJX17, CEL15, CDS02, sCLC13, CMX97, CLZ02, CLL08, CZS16, CT17a, CKW10, CL02, CN96, CHY11, CCE^{+}12, DHK98, DLS14, EHV17, FK98, FGO14, FH18, GM17, GG16, Gna07, Gom95, Gu17, GH06, yGW07, yGjW10, GW14, GPG15, Han05, HKRS14, HL06, HH16, HR13, HO17, HAS05, HX17, Hu07, HXC10, HY14, HY18, HS16, IUV17, Jan98, JM03, JS05, JK08, KNP98, KY01, KH05, KHM16, LN14, LLS03, Li05b, Li08a, Li08b, LC09, LG13, LC13, LZY98, LLC16, LWL17, MR15, MR96, MV99, MST05, MY17, MS98, Mur08, OP11, PM17, Pla12, Ryn00, SL15, SXYY06, SW98, TZL16, lWyG06, WY16, WY17, WHS18, WZ10, Wie15, WLPV15, WZtX16, Xie08, lXyG02, Xu14]. method [Yao16, gYyG12, ZT14, ZP06, Zha14, ZPY15, ZT15, ZW12, ZL16, ZhXpZ16, lZmCRX03, ZQ13, Zhu15, ZG16, dCB07]. Methods [Ano01b, ASU17, ABI15, AK15, Awa15, BP93, BCM99, BBB^{+}16, BDMR10, BEPS96, BJX09, BS97b, Bur97, BS00, But97, CDP03, CLM03, CLMR10, CLMR15, CZZ17, CMX02, CWX06, CWX12, CL16, CEHK18, CSWP99, Cul96, Cum96, DDP14, DFR07, DPS93, DMMS11, DS01, DRS08, DLZ16, DT12, EJ97, FLM96, Fas99b, FYL08, FR10, FS98, GMM09, GO95, yGqW09, HBMY14, Hau97, HS08, HR02, Hua06, HZHL12, IKS10, JMO00, JSSE97, JLZ08, KXZ04, KN14, KNQ16, LW04, Lee03, Lee16, Li05a, LK05, LMV11, Liu99, LLY12, LL10, LL12b, LNW02, MN16, MNC16, Nai12, Nie98, PIZ01, PS98, Ram95, RWT13, RZ98, Sch15, SXZ06, SX09, Sid17, gTpM02, TBD94, TGA96, VC00, Wan12, WL12, WMY13, WMB13, Wie15, Wil98, XB05, XZ03, XY12, ZVW95, ZB12, ZD16, vdH93, vdHdS97, vdHv97, vdHS00]. metric [CZS16, LD07]. MFS [FKS05]. Microlocal [GL17]. midpoint [SC08]. Mindlin [AL97, Ye00]. Mini [KL00]. Minimal [Sid17, Xu98, PGM09, Wen95]. minimax [Yao16, Zwi94]. Minimization [MR10, CM96, zCpST09, CZS16, DT12, Fli18, FRT09, GN08, HH16, MNPR06, WFL02]. minimizations [ZYY16]. minimizers [MRS18]. Minimizing [Bün11, Far10, FGS13, FGG^{+}17]. Minimum [Noa14, ABC14]. Minkowski [FMR00, FH05, KJ09]. MIRKs [Ben97]. misclassification [CM96]. missing [CCSS09, Sch11a]. Mittag [GP13]. Mittag-Leffler [GP13]. Mixed [AAQ15, Li05a, lXyG02, AK15, Awa13, CHM18, DGMM16, DLZ16, FF95, GMM09, Gil17, JLZ08, Lee16, LMT01, MS16, Xie08, YDF97]. MLPG [AS05]. mode [CHRX06]. Model [BOP^{+}15, SS15, AV15, AL97, BNP14, BGH15, CHM18, zCpST09, FE15, GLS15, HY14, HX02, JZ10, KORS17, Kim06, KCW17, MSXZ13, Ock96, RH15, SI15, XKY15, Yan03, ZPR03]. Modeling [Ano02c, IL16, BWX97, DAP13, TS05, ZPR03]. modelling [SEE96]. models [ARAK09, CLM96a, CGH^{+}13, CW09, PGB15, RWT13, Wat96]. modes [BPJ02]. modification [FGS16]. Modified [KL00, WMY13, Bai99, Li08b, RJ00, WLW16]. modular [TL18]. Mollification [BB07]. mollifiers [CGW13]. moment [BR07, Füh16, TQ17]. moment-based [TQ17]. moments [CHR00, CHSS03, CaL06, CCH^{+}15, Han10]. Monge [Awa15]. mono [Mui99]. mono-implicit [Mui99]. monomials [dB07]. Monotone [Kva14, ABY13, yGmW98, KL04, KL07, nCV13, Wan12]. monotonic [Xu14]. monotonicity [FP98, FBCR13]. monotonicity-preserving [FBCR13]. MoRePaS [BOP^{+}15]. Morley [HSY16, YLBL12]. morphing [LD07]. mortar [FH18]. mortaring [LRS12]. motion [BNP14]. motions [FGS13]. motivated [Wie97]. moving [BMSR^{+}16, FE15, Ock96, TZ03]. MR0019410 [CN96]. MR1338896 [Hem96]. MR2438345 [Fus08a]. MR2783300 [Dũn12]. MR3428571 [Ano16]. MR3428572 [Ano16]. MR3428573 [Ano16]. MR3428574 [Ano16]. MR3428575 [Ano16]. MR3428576 [Ano16]. MR3428577 [Ano16]. MR3428578 [Ano16]. MR3428579 [Ano16]. MR3428580 [Ano16]. MRA [PŠWX03]. Multi [HL06, LSXZ15, RZ98, BR15, CMR07, CaL06, CHS17, HH16, KCW17, LY07, LMWX13, Onc14, PGB15, gTpM02]. multi-asset [PGB15]. multi-block [HH16]. multi-dimensional [BR15, Onc14]. multi-frequency [KCW17]. multi-interval [gTpM02]. multi-knot [LY07]. Multi-level [HL06]. Multi-parameter [RZ98]. multi-revolution [CMR07]. Multi-step [LSXZ15]. multi-subdomain [CHS17]. multi-task [LMWX13]. multi-wavelets [CaL06]. multibody [FE15]. multicategory [CX06]. multidelay [ZV08]. multidelay-integro-differential [ZV08]. Multidomain [BK16]. multifilter [Jia00]. multiframelets [LYY13]. Multigrid [BH02, BA15, BEPS96, GPG15, Hem95, Hem96, KXZ04, KvL95, SXYY06, ZVW95]. multilayer [WFL02]. multilayered [Mha93]. Multilevel [BMK15, CWX06, KP95, Kun95, XH13, CHM18, DMMS11, Fas99b, GJS14, GO95, HX02, Nie98]. multimedia [Ren09]. Multiobjective [IUV17]. Multiparameter [BJX09]. Multiple [HJH12, CDTV99, GPT17, MXO13, Plo95, Yao16]. multiplication [AC00, BD93]. multiplier [Hu07, WLW16]. multipliers [Kun95]. multiply [HL16b, San15]. multiply-connected [HL16b]. multipoint [DLZ16]. multipole [GG16]. multiquadratic [FHN94]. Multiquadrics [BLS02, Buh06, HL16a]. multiresolution [ADL11, AL13, ABD08, CGRS15, Jia11, KKL03, RAB^{+}09]. Multiscale [CL16, Opf06, BJX09, GL17, Hua06, JK08, KORS17, LW14, LSW17, Li03, LLY12, MR15, Var96]. multisplitting [WMY13]. Multistep [FJ99, BS97b, BS97a, BS00, Wil98]. Multivariate [BL99, Car95, LP04, MS04, SRFH12, BC99, BCZ05, Cal05, CHR00, CMX07, CDTV99, DS13, Lai06, Le 00, Li04, NSW16, Sau95, SX95, Sau06, SS96, SW08, dB00, Sch04]. Multiwavelet [HM03, PS98, LYY13, CMX97]. multiwavelets [MS97]. multiwindow [JJK13]. Mumford [KPY09].

nat [Sil99]. native [Fus08a, Fus08b]. natural [Dou94]. nature [LO08]. Navier [Ang00, APV14, CHM18, CGS93, HL06, JK08, LC13, PCM^{+}96, ZVW95, dFGAJN18]. Near [BDKY02, DSW05, MM00, SFDE15]. near-field-to-far-field [SFDE15]. Near-optimal [DSW05]. necessary [MNPR06]. Nédélec [Xie08]. Nested [ZT14, KP98]. nets [Sto96]. network [CD15, Len96, Mha93]. Networks [EPP00, Noa96, BE00, CLM96b, Len94, MM00, MNW99, Nee96, Wat96]. Neumann [AHC11, LST11, WL12, ZB12]. neural [BE00, CLM96b, Len94, Len96, MM00, Mha93, Nee96, Wat96]. neutral [ZV08]. Newton [CG15, EHV17, HY14]. Newton-penalty [HY14]. Newtonian [DGMM16, Li10b]. Nirenberg [GP14]. Nitsche [LRS12]. Nitsche-mortaring [LRS12]. no [Fus08a, Hem96]. node [Gom95]. noise [LO08]. noisy [BE00, NG99, RH12]. Non [Dũn09, LW03, AK01, ABY13, Atr12, ANSZ17, CJ07, CGR11, DLUS17, DY01, Fra99, JKLY13, JM00, Lev99, LHY08, LP14, MO16, PSNP11, TYY11, Wan12, WMY13, WMB13, XP10, ZWDD16, ZB12, ZD16]. non-autonomous [MO16]. non-constant [ZB12]. non-greedy [TYY11]. non-Hermitian [WMY13, WMB13]. non-homogeneous [AK01]. non-identical [XP10]. Non-linear [Dũn09, ABY13, CJ07]. non-monotone [Wan12]. non-oscillatory [DLUS17]. non-overlapping [LHY08, PSNP11]. non-periodic [LP14]. non-polynomial [Fra99]. non-rectangular [ZWDD16]. non-selfadjoint [JM00]. non-stationary [CGR11, JKLY13]. Non-Uniform [LW03, Atr12, ANSZ17, DY01, Lev99, ZD16]. nonconforming [LLS03, LY08, MC08]. nonconvex [SL12]. nonequispaced [Nes16, Ste98]. nonhomogeneous [AC05, HY18, Li04]. nonintrusive [CEL15]. Nonlinear [AL13, Noa98, ADL11, AZW15, BQRB13, CLC16, DKLT93, GH99, GMS12, yGmW98, GW14, KNV01, LT98, MN15, Osw04, QWXZ10, Sun14, TL18, XZ01, XZ03, ZPR03]. Nonlinearity [Ito96]. nonlocal [FLM96, TDG16]. nonmatching [PS05]. nonnegative [Gon93]. nonregular [LT98]. Nonstandard [MC10]. Nonstationary [DM15, Pit16]. nonsymmetric [dH94]. Norm [CK03a, Bün11, CK11, KL17, Nie98, ZPR03]. normal [GG16]. Normalized [BF03, JKLY13]. norms [MRS18, Wat94]. note [BI11, CLZ02, Dai16, Del17, Kun09, Ste98]. Null [KCW17]. number [RJ00, TL18]. number-theoretic [RJ00]. numbers [Sch95]. Numerical [AG98, CGH^{+}13, CCH^{+}15, Cum96, DDP14, DL98, ER99, FR10, FQ10, GT14, Hau97, HW12, HXC10, Kei95, KMO^{+}14, Mar94, OS09, Pen98, PST08, RHZ17, She13, SJJL15, TBD94, WH07, XY12, ZVW95, ZGGW12, ZB12, ZD17, ADG17, BPW95, BF11, BGMS07, Bes03, Ble07, BP11, Cas07, CX04, CN96, CT17b, DHK98, DD94, DLP98, EJ97, FM14, GVSLN96, HO17, Hol01, KS17, LLS09, LL12a, Lut04, Mur08, OP11, SW04, WHS18, WLPV15, XL10, BMSR^{+}16]. numerically [AD08]. NURBS [BX94]. Nyström [CSWP99, DLS14, HBMY14, HH15, LG13]. Nyström-type [CSWP99].

objects [SF14]. oblique [CG10]. Observability [JRS15]. observable [LSLS14]. Observations [DHK98, GPT17, XWL13]. observers [GT14]. obstacle [BBR02, IKS10, LMT01, Yan01]. obstacles [CH15, ZGGW12]. ocean [SI15]. ODE [AB97, Ben97]. ODEs [Bur97, EJ97, JM00, WK93]. offset [PGB10]. offsets [CCH^{+}08]. One [ZYY16, CLM96b, HT12, KORS17, LZZ09, MS12, XKY15, ZG16]. one-dimensional [LZZ09, XKY15, ZG16]. one-level [HT12]. Online [GYZ17, Yin07]. onset [IL16]. onto [CWLH95]. open [CL18]. operational [ESM13]. operations [PR15]. operator [BMK15, DFR07, FQ10, GM14, Lut04, MC10, RH12, RSZ11, XZ03]. operators [BDD06, CS06, CCS16, DL04, DY10, FY13, Gna07, GLT93, GP14, JL97, KP98, Kir98, KK02, Len96, Li94, LT98, MR10, MXZ08, PL05, nCV13, ZD16]. Optics [Ano02c, BKN11, BPJ02]. Optimal [BEG17, CS06, CKW10, Dũn11, GL13, LO08, MRS14, Maz04, Mui99, CLC16, CLZ02, DSW05, DP13, GHdN12, GH06, GN13, HJJV97, KP95, Lee03, LY01, MRS03, NZ04, RH12, RJ00, Sch15, Yan03, dLdDSM07, Dũn12]. optimality [CP93, MM00]. Optimally [LP04, ČF11, RST^{+}14]. Optimization [SLN14, BHT11, BA15, CD15, ESM13, IUV17, LSXZ15]. Optimized [Onc14]. option [lZmCRX03]. orbits [CS93]. Order [But97, FW15, AL97, BS05, BOP^{+}15, BGH15, BA15, CDP03, CLM03, CWX12, CEHK18, FS98, GW17, GLS15, HBMY14, HKR04, HS03, HLN12, JMO00, JN18, KCW17, KL16, Li05a, LD16, LL12a, MRS03, Mat09, MS04, Möß10, NS04, Noe95, NZ04, OBS15, PM17, PSNP11, RHZ17, SI15, SS15, TBD94, TGA96, WOBL17, Wie97, XL10, Xu14, ZP06, ZPR03, ZT15, ZD16, ZZ09]. Order-preserving [FW15]. ordering [CM99, Dou94]. ordinary [yGqW09]. oriented [BWX97]. originated [MH17, Zhe06]. Orthogonal [Bes03, CMP07a, FM04, LH13, CI13, Dam07, DHO16, JS99, Jia00, MPS96, MPS97, Rei93, Sal98, SI15, SX07, TSY10, Wat94]. Orthogonality [MPS96, MPS97]. orthogonalization [CK07]. orthogonally [Wat94]. orthomaps [ZB99]. Orthonormal [QWXZ10, CaL06, Han98, Han10, HHS04, Lai06]. oscillatory [DLUS17, NUWZ17, XHC15]. Oseen [LLC16]. outflow [Kno09]. overlapping [HS03, LHY08, PTC12, PSNP11].

p [WY17]. package [BD93]. packet [CR08]. Padé [BC99, BCZ05, CDTV99, Dar03]. Padé-type [Dar03]. pages [Ano16]. pair [ZT15]. Pairs [CG10, Chr15, NSS04, CKK16, GLS15, Lem09]. pairwise [GYZ17]. panel [HH15]. panel-based [HH15]. parabolic [AV15, BA15, FLM96, GLO10, KL17, Le 05, Li05a, LB93, gTpM02, VK96]. Parallel [AB97, AH96, BP93, BF94, BS97b, BS97a, Bur97, BS00, Chi03, Jan98, vdHdS97, Ben97, But97, CSWP99, EJ97, Moo17, WMY13, WLPV15, XZ01, ZSHZ15, dH94, vdH93]. Parameter [HO17, GPT17, HO15, JJLR14, Kno09, LLY12, Nes16, OP11, RH12, Rip99, RZ98, Sch11b, She13, Wie15, WK93]. Parameter-uniform [HO17, OP11]. Parameterization [Jia00, CLM02, CL18]. Parameterizations [Jia03, HL16b]. parameterized [BOP^{+}15, CM99, SS15]. parameters [EHV17, FYL08]. Parametric [BGH15, KKV15, CC94, FE15, GPS10, KK99, SFDE15]. parametrized [MN15]. Parseval [CK07]. part [WZL17]. partial [Atr12, CJX17, CN96, yGS08, HS08, PM17, SW96, She13]. particle [KH05, XB05]. partition [CM12, HL05, NSS04, ZSHZ15]. partitioning [Wie15]. partitions [MS07, SS04]. Parzen [ZZ09]. pass [Jia09b]. Past [Ano03c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b]. Patches [Kra02, BLP02, GPS10]. PCA [WW12]. PDE [AV15, BHT11, CW09, DAP13, IUV17, Sch15]. PDE-constrained [BHT11, IUV17]. PDEs [Le 05, MN15, VK96]. penalized [BL09]. penalty [GS13, HY14]. pencils [CG06]. perceptrons [WFL02]. perfect [EGL13]. perfectly [LZZ09]. period [SC08]. Periodic [CDS02, Zha05, AB02, CP03, CG10, DS01, FW15, GY00, LL13, LP14, MP00, She13, Tor16]. periodization [Søn07]. permutation [NSW16]. permutation-invariant [NSW16]. persistence [PZ16]. perturbation [Che06]. perturbations [She08]. Perturbed [Ram95, RSZ11, AB02, BBR02, CK11, GLO10, HO17, Kun01, LRS12, MR15, MR96, MXO13, OS09, OP11, ZC14]. Petrov [AS05, CMX97, CMX02, DS01, Hua06, WY17]. phase [BH15, Han10, KL16]. phases [QWXZ10]. phenomenon [BC99, RY13]. photoacoustic [DGK15]. photographs [ZB99]. photonic [KK02]. physical [TS05]. pi [Len94, Len96]. Piecewise [Gu17, Wen95, ABC14, BGM93, Cas07, HL05, JSSE97, KPY09, LY07, LM06, PR99]. piecewise-polynomial [ABC14]. pipeline [CM99]. pipeline-efficient [CM99]. planar [BLP02, BGN17, DOZ94, LLS03, Zwi94]. plane [GMM09, HBMY14, HY18, Li94, PC17]. plasma [CG17]. plastically [SEE96]. plastically-formed [SEE96]. plate [AL97, FP99, Joh13, Ye00, Yoo01, ZB99]. plus [SS08]. POD [AV15, RH15]. Poincaré [KP98]. point [CZS16, DSW05, DLUS17, Dou94, GMS12, yGmW98, KS17, Leo13, LSXZ15, MR96, SC08]. points [BLP02, Gom95, GH08, IP15, Pan95, Pra98, SW04, Sto96, Ven94, WHS18]. Pointwise [Dam07, NV10]. poised [BH17]. Poisson [ALZ02, AK15, BL09, CBK01, Li05b, Li08a, ZSB01b, ZhXpZ16]. poles [VK96]. polygonal [GH06]. polygons [FHK06, FK10, Jon17, KB15, LD07, RGB13]. polyharmonic [BRS13]. polyhedrons [HR17]. Polynomial [GS00, JZ04, MNPW00, Xu00, Xu04b, zCFX07, ABC14, Bar03, BGM93, BNR00, BR13, CT05, Cho95, DS00, Der04, FGG^{+}17, Fra99, HW12, JKLY13, Le 00, Lev99, MV99, Maz99, MWW13, MS00, Mon09, Nie07, Pan95, PS07, PR99, Sau95, Sid17, Sor18, Wen95, WS01, ZJ06, dB00]. polynomially [LP10]. polynomials [Bün11, CMP07a, CI13, CCS16, DZ04, DHO16, DKB99, DGS18, DS13, Flo96b, GMRS97, IP15, Jüt98, LKW17, MPS96, MPS97, MS04, MN00, Pot04, Rau05, Rei93, Sal98, SX07, SS08, WSM13, Xu04a, dB06]. polytopes [CCH^{+}15, War96]. porous [DGMM16]. posed [BL09, BE00, LLY12, XY12]. Positive [LS05, Bai99, De 06, GP95, GS10, Luo99, Pin04, SM99, WMY13, WMB13, Wen95, Wu95]. positivity [CEHK18, LM06, Maz01]. positivity-preserving [CEHK18]. post [LZY98]. post-processing [LZY98]. postbuckling [Mur08]. postconditioning [FJ99]. Posteriori [Ano01b, Bra01, CBK01, CK11, DY01, KY01, KL17, MSZ06, MRH15, MW01, ZC14, Kun01, AK01, ARPR01, DLZ16, GN13, HR01, KNV01, LY01, LLY12, Pet02, PS05, Yan01, Yan03, Zha01, ZSB01b]. potential [Li09, aKT17]. potentials [IMS99, Li10b]. Powell [AS02, LLS96, LL98, MS07, Rem12, SW08]. power [ASS10, LS98, NSW07, RS99]. power-growth [RS99]. Practical [WMB13, CN96, WZL17]. precise [Jon17]. precision [GPS10, NP18]. Preconditioned [LNW02, BCM99, MS98, RSZ11]. preconditioner [Bre95, FH18, HS03, LK05, PTC12, PL05, Sto17, Zhu15]. preconditioners [Bai99, MS99, NBL11, SW98, TSZ98]. Preconditioning [Wan12, vdH93, BHT11, BPK96, CHY11, JS95, KP95, KvL95, Kun95, MV01, ZT14]. predictive [AV15]. predictor [CSWP99, ZD14]. Preface [AS03, Ano00b, Ano02d, Ano04g, Ano05c, Ano06, BP11, CHY10, Dah95, GOZ03, NW04, Sch04, SX09, Twi96, XY12, BOP^{+}15]. prescribed [CL02, FMPS13, MRS18]. presence [AL13]. Preservation [CK03b, CLMR10, CMP07b, FP98]. preserve [MXZ08]. Preserving [CP04, CP93, Car95, CJ07, CEHK18, CCS16, CGM01, CM03, CP07b, Der04, FBCR13, FW15, GO05, HR17, KNP98, KvD01, SS14]. pressure [BR15]. pricing [lZmCRX03]. primal [DGMM16, KPY09]. primal-dual [KPY09]. primal-mixed [DGMM16]. Prime [LMO14, Kot08]. prime-counting [Kot08]. principal [IB17, SLN14]. principle [IKS10]. principles [Kos15, LSLS14, MSST14, RT14]. problem [AK01, ACH10, AHC11, BGMS07, BMSR^{+}16, BE00, CK11, sCLC13, CH15, CL07, CMX07, CZS16, CL16, CG17, FLM96, FF95, FRT09, GMM09, yGmW98, GN13, HO17, HT11, HXC10, HFH15, KL00, KS17, KCW17, LST11, Li05b, LLZ10, fLL05, MR15, MC08, Mur08, OP11, Pei05, QC12, TW98, Wan08, WZ10, XY15, Yan03, ZB12, ZPY15, ZC14, ZW12, vdH93]. problems [ARPR01, AC05, ABI15, BL09, BKN11, BHT11, Bes03, BDMR10, BA15, BK96, BS97a, BS00, But97, CZZ17, Cas07, sCLC13, Che06, CWX12, CHY10, CP15, CHS17, CHY11, CCE^{+}12, DDP14, DRS08, DLZ16, EHS02, FK98, FKS05, GM17, GM14, Gil17, Gna07, GO95, GH06, yGjW10, Hak00, HKRS14, HH16, HH15, Hem95, Hem96, HR01, HO15, HJH12, HX17, JN18, KORS17, KXZ04, KPY09, KF03, KNV01, Lee03, Li05a, LSXZ15, LRS12, LMT01, LY01, MSZ06, Mar94, MO16, MR96, Mat09, MNC16, OS09, Ock96, PP97, PSNP11, RSZ11, SXZ06, Sto96, Sun14, TZL16, TDG16, Tom96, TBD94, VW05, WL12, WHS18, Wat94, XZ01, XY12, Yan01, Zha14, ZL17, ZL16]. procedure [BRZ94, PSNP11, Sch11b]. process [ALZ02, DAP13]. processes [CDP09, SU12]. processing [Jia11, LSXZ15, LZY98, PZ16, RT14, SX09, Sun14]. processor [AH96]. Procrustes [Wat94]. product [Bün11, CP15, Flo94, FP98, GO95, KK99, LS05, Pin04]. products [FMR00, GY00]. Professor [Sil99]. projection [APV14, CLMR15, HT12, KH05, LH13, Mat09]. projectors [She08]. prolate [LST11]. Proof [FP99]. proper [SI15]. properties [AB11, BS06, CR08, CHL17, CGR11, CL02, GGL07, JL97, Mai10, MPS96, MPS97, Mha93, MNW99, PŠWX03, Pla12, RAB^{+}09]. property [BH17, CGP95, CM12, JKLY13, MP96, Pré94, SL12]. proportional [XH13]. proportionality [CHMR97]. Prössdorf [Sil99]. Proximity [MSXZ13, LSXZ15]. Pseudo [Awa15, Mai06]. Pseudo-dimension [Mai06]. Pseudoaffinity [Maz07]. pseudodifferential [DPS93, DS01, FQ10, GP14, PS98]. pseudoframes [LO08]. pseudospectral [yGSlX03, lWyG06]. Publisher [Ano01c]. Puiseux [WZL17]. Pythagorean [AF96, FS94, FaKS02, FGS16, JKK^{+}12, KJ09, KKV15, PFMS05, Pot95, RSA14]. Pythagorean-hodograph [AF96, FS94, FaKS02, FGS16, JKK^{+}12, PFMS05, Pot95, RSA14].

QCMC [HL16b]. QMR [Cul96].

QMR/BiCG [Cul96]. QTT [KORS17].

QTT-finite-element [KORS17].

quadrangulations [NSZ04]. Quadratic [CLLS06, MP04, MS07, Noa02, CLMR10, Kva14]. Quadratic-Cycloidal [MP04]. quadratically [MN15]. Quadrature [Nai12, BF94, FXZ96, FM14, GVSLN96, HZHL12, MC10, SX07, XHC15, aKT17]. quadratures [LL12a]. quadrilateral [AAQ15, FF95, Jia11, LY08, LPSSP00, LL10, LL12b, MC08]. quadtree [KHM16]. qualocation [JSSE97]. Quantile [SHTS14]. quantiles [Xia13]. quantization [LPY10, Wan08]. quartic [MH17]. Quasi [KKL03, KKLY10, NV10, CP99, CSS98, DL98, DGMM16, Dũn09, HL16b, LM04, MS07, RH12, Rem12, Spe17, WL05]. quasi-affine [CSS98]. Quasi-biorthogonal [KKL03]. quasi-conformal [HL16b]. quasi-interpolant [Dũn09]. quasi-interpolants [MS07, Rem12, Spe17]. quasi-interpolating [DL98]. quasi-interpolation [LM04, WL05]. Quasi-interpolatory [KKLY10]. quasi-Newtonian [DGMM16]. quasi-optimal [RH12]. quasi-wavelet [CP99]. quasigeostrophic [SI15]. quasilinear [LB93]. quasiuniform [GH06]. Quaternion [Far10]. quintic [Far10, FGS16, SS04]. quintics [KJ09]. quotient [MS04].

Rachford [HH16]. Radau [BS97b, BS97a, BS00]. Radial [BRSV15, LR07, Li10a, Li10b, BCM99, BGH15, CL16, CSW96, DSW05, Fas99a, Fas99b, FF05, FS98, FW15, GW17, Hub12, LST11, LW14, LM14, LLY06, Li05a, Li09, LK05, Low05, Mai05, RS99, Rip99, Sch95, SS08, SRFH12, WH07, Wen95, Wu95, Xu98, dB06]. radial-basis [RS99]. Radiation [MKS02]. radiosity [AC00, Han05]. radius [DY10, Möß10]. Radon [CG15]. raising [FL00]. Ramlau [HR13]. random [CP15]. randomized [ZZ09]. Rank [NSW16, LMWX13, Sid17]. Rank- [NSW16]. Rate [Li08b, Sun08]. rates [FJ99, LR07]. Rational [TSY10, CLM02, DP13, DOZ94, EGL13, Far10, FGS13, FGG^{+}17, GMB97, JJK13, JW95, Pot95, VK96, gYyG12].

Raviart [CL07]. ray [GL13]. RBF [Han12, LSW17, Sch11b]. RBF-interpolation [Sch11b]. RBFs [Fus08a, Fus08b, NSW07]. reaction [CK11, GLO10, KC16, Kun01, MR15, MXO13, ZC14]. real [ANSZ17, BMB14, Dam07, FRT09, GP13, IB17, NUWZ17, Pan95, Pin04, VK96]. real-life [FRT09]. real-time [ANSZ17]. real-valued [IB17]. Reconstruction [DMT03, Var96, ARAK09, Atr12, BDMR10, CCH^{+}15, DM15, EGL13, FPR^{+}12, GP14, KL15, SW06, Sun10]. recoverable [KL15]. Recovering [WL12, GN08]. Recovery [FZ12, CCSS09, CS06, DY01, Dũn09, Dũn11, GL13, XWL13, Yan01, Yan03, Zha01, Dũn12]. rectangles [CDP03]. Rectangular [Awa13, DM99, Rem12, ZWDD16]. recurrence [EHV17, MXZ08, RL07]. recursion [Wal95]. recursions [Wal95]. Recursive [Cop98, TZ03]. Reduced [ASU17, GLS15, MRH15, PGB15, AZW15, CEL15, IUV17, RH15, SI15, SL15, Sid17, Wie15]. reduced-basis [IUV17]. reduced-order [SI15]. Reducing [Ant18]. reduction [AZW15, BOP^{+}15, BGH15, FE15, HO15, KCW17, PS07, SS15, Tom96]. redundancy [BH15]. Refinable [Gro13, HM03, CX04, CTZ02, De 06, GGL07, JS99, KKLY10, LLS95, Li04, MS04, Pit16, Sau06]. refined [GJS14, KP95, KP98]. Refinement [DX10, ARPR01, DM93, DY10, Han06, JL97, Li04]. reflection [Wan12]. reflectivity [DMT03]. regard [TYY11]. region [DLK12]. regions [CMP07b, JW95, KvL95, Pet96, San15, TB16]. register [CM99]. register- [CM99]. regression [SHTS14, SW10]. regular [Led15]. Regularity [MS97, MXO13]. Regularization [EPP00, Zoz15b, Zoz15a, BJX09, DSS09, FQ10, GS13, HT11, MZ13, RH12, SHTS14, ZW12]. Regularized [SW10, DGK15, GYZ17, Hu07, WZ10]. Regularizers [MMP13]. regularizing [Pla12]. Reissner [AL97, Ye00]. related [DGS18, Sid17, TDG16]. Relation [PZ16]. relations [EHV17, RL07]. relaxation [HJJV97, vdHv97]. relaxations [Las16]. Reliable [CCE^{+}12]. Remarks [HJJV97, DT96, Dou94]. removal [RY13, SEE96]. representation [BDD06, CWLH95, CLM02, Lev99, MSST14, MNW04, Nie98]. Representations [Ano02d, CP93, Dar03, Dũn09, GN08, GL17, Hub12, TZ05]. reprint [CN96]. Reproducing [FY13, HT11, JKLY13, RY16, vdMNS03]. Reproduction [JZ04]. rer [Sil99]. residual [APV14, MW01]. residual-based [MW01]. resolution [HFH^{+}07]. respect [CMP07a, Rei93]. response [TZL16]. restarted [dH94]. restoration [Del17, FRT09]. Restricted [CMP03, SL12]. Restriction [AD08]. result [JS95]. results [GVSLN96, Kun09, MN00, NSW98]. retarded [Hau97]. RETRACTED [Zoz15b]. retrieval [BH15, Wan12]. reversible [CS93]. revisited [BRZ94]. revolution [CMR07]. Reynolds [TL18]. Riccati [BF11, KA97]. Richardson [JLZ08]. ridge [Mai10, WFL02]. Riemann [Bra01, EGST13, JN18, KS17]. Riemannian [GH16, Noa06, Noa14, WYW11, YZ08]. Riesz [Grö03, HHS04, Leo13]. right [FR10]. right-hand [FR10]. ripplets [Pit16]. risk [MNPR06]. RKDG [ZQ13, ZG16]. RKN [Ram95]. RMS [Nes16]. Robin [PSNP11, SJJL15]. Robin-type [PSNP11]. Robust [Kun01, ZC14]. robustness [ZSB01b, ZYY16]. role [Joh13, Ock96]. Rosenbrock [BBB^{+}16]. rotating [SF14]. rotation [FaKS02, Far10, FGS13, FGG^{+}17]. rotation-invariant [FaKS02]. rotation-minimizing [Far10, FGS13, FGG^{+}17]. rounding [Bar03]. Roundoff [TZ03]. rule [SC08]. Ruled [SR02]. rules [Han10, NSW16, RH12, RJ00, WYW11, WZL17]. Runge [BP93, BS97b, BS97a, BS00, CHMR97, CLM03, CMR07, CLMR10, CLMR15, CSWP99, Hal95, JMO00, Liu99, Mui99, vdHdS97, vdHS00].

Sabin

[AS02, LLS96, LL98, MS07, Rem12, SW08]. Sacker [LMV11]. safe [HKR04]. salesman [Wan08]. Salzman [Han05]. same [CGR11]. sampler [FPR^{+}12]. samples [SW10]. Sampling [RZ10, WSM13, vdMNS03, ARAK09, Atr12, Dũn09, Dũn11, Dũn12, GZ13, JJK13, Søn07, SZ09, Sun10, Sun14, Tan17, XP10, ZZ09]. sampling-reconstruction [ARAK09]. satellite [FM04]. satisfying [TSY10]. SBFs [NSW07]. scalable [WLPV15]. scalar [BR07]. scale [BF11, BRSV15, CS94, Plo95, RH15, SI15]. Scaled [BLS02]. scales [MXO13]. scaling [CP03, Han98]. Scattered [DZ04, FHN94, Grä12, IP15, Kun09, LW04, LNW02, MNW99, NSW98, NG99, Pot04, WL05, WSM13, Yoo01]. scattered-data [NSW98]. scattering [Cai02, Gil17, HR02, IKS10, LN14, MKS02, QC12, VW05, WL12, ZGGW12]. scheme [BNP14, BS05, CLC16, CJ07, CL07, CGRS15, CT17b, FGS13, HFH^{+}07, HH15, LY08, MB96, Wie97]. schemes [ADL11, Ble07, CMR07, CD15, CMX97, CGR11, CGR13, DD94, DLUS17, DS13, Flo96a, GMS12, JKLY13, Jia95, JN18, LLY06, Lev99, LP10, LQ14, LD16, MS12, Mui99, MH17, Nie07, Noe95, SW96, TQ17, TDG16, Zhe06]. Schmidt [CK07, Chr15]. Schoenberg [SS96]. Scholes [PGB15]. Schrödinger [BK16, CLC16]. Schur [Dai16, Peñ11]. Schwarz [Bre95, BR07, PTC12]. Schwinger [Che02]. SDFEM [ZL17]. Second [BA15, CHS17, HLN12, TW98, TGA96, Bra01, CMX97, CP99, HS16, Mar95, PSNP11, RHZ17, XL10, Xu14, ZP06]. Second-kind [TW98]. Second-order [BA15, HLN12, TGA96]. sections [Flo96a]. sectorial [FM04]. segment [MNW96]. segmentation [DLK12, KPY09, RWT13]. segments [Pet96]. Seidel [Dou94]. seismic [Wan12]. selecting [Rip99, Sch11b]. selection [CHMR97]. self [FMPS13, SM00]. self-adjoint [FMPS13]. self-associated [SM00]. selfadjoint [JM00]. Sell [LMV11]. Semi [Buh06, PIZ01, BS05, CLC16, JRS15, LC13, PP97]. semi-cardinal [BS05]. semi-discrete [LC13]. semi-discretization [JRS15]. Semi-explicit [PIZ01]. semi-implicit [CLC16]. Semi-infinite [Buh06]. semi-linear [PP97]. semidefinite [Las16]. semidiscrete [CW16]. semidiscretized [VK96]. semilinear [AV15, Cas07, CK11, FLM96, KL17]. seminorms [AM14, LL99]. sensing [SL12]. Sensitivity [KA97]. separable [CS03]. separating [lZmCRX03]. separation [Leo13]. sequences [BD10, BR07, GGL07, Grö03, LY07, LYY13, MP96, SZ09]. series [CDTV99, LT13, Pré94, QW11]. set [BLP02, FGG^{+}17, GLT93, KPY09, Pan95, XKY15]. set-valued [XKY15]. sets [BH17, BS06, CG15, Cho95, FMR00, FH05, Leo13, LL13, LL11, WSHD07, Wie15]. several [GS00, MS12, Xu00]. Shah [KPY09]. shallow [CT17b, IL16, vdHS00]. Sham [CGH^{+}13]. Shape [CP93, CMP07b, CM03, CP04, CP07b, Der04, GO05, CH15, CGM01, HSV09, KvD01]. Shape-Preserving [CP04, CM03, GO05]. shaped [Pit16]. shapes [Alb15]. shared [Jan98]. shared-memory [Jan98]. shearlet [GL13]. shells [YDF97]. shift [CHL17, HHS04, Sun10]. shift-invariant [CHL17, HHS04, Sun10]. shifted [Yoo01]. Shishkin [Mat09, ZL17]. shooting [CS93]. sides [FR10]. Siegfried [Sil99]. sifting [DAP13]. Sigma [Wan08, LPY10, Len94, Len96]. sigmoidal [Sto96]. Signal [MSST14, PZ16, RT14, Sun14]. signals [Ren09, Var96, Wan12]. signatures [CC94]. simple [CG04, CKBP11, MST05, MY17]. simplices [Wie97]. simplified [HY14, RHZ17]. simply [CWLH95, CL18]. simply-connected [CWLH95, CL18]. Simulation [Zhu15, BGMS07, FE15, XKY15]. simulations [CEL15, ZG16]. Simultaneous [Füh08, LT06, PCM^{+}96]. Single [gTpM02, EJ97]. single-eigenvalue [EJ97]. singly [Tor16]. singular [ByLl01, Bes03, Boc04, Che06, CKP13, FR10, HX17, HS16, Pla12, Smo07]. singularities [HZHL12, MV99, MP00]. Singularity [KNP98, lZmCRX03]. singularity-separating [lZmCRX03]. singularly [CK11, GLO10, HO17, Kun01, LRS12, MR15, MR96, MXO13, OS09, OP11, ZC14]. six [CMP07b]. six-dimensional [CMP07b]. size [BS00]. slits [San15]. small [BGH15, CKK12, NSS04]. SMMP [BD93]. Smolyak [Pet00]. Smooth [AS02, BK96, Car95, HBMY14, JSSE97, LP14, MM00, RZ10]. smoothing [BB07, FJ99, Fas99b, MST05]. Smoothness [Osw04, Pra98, CSW14, FBCR13, Li04, MT98, SRFH12, WYW11]. snapshot [ASU17]. Sobolev [DG16, FY13, Han06, KL07, LYY13, War13]. Sobolev-type [KL07]. Software [WLPV15, CKP13, Mac94]. solidification [PCM^{+}96]. solidification/melting [PCM^{+}96]. Solution [FH05, LST11, ADG17, BPW95, BF11, BA15, Bra01, Cum96, DPS93, DMMS11, DLP98, EHS02, EJ97, GM14, KS17, KC16, Mar94, Pen98, Wat94, ZVW95, ZYY16]. Solutions [Han06, AC05, Ant18, CLZW13, CN96, ER99, FK98, HBM03, JLZ08, JN18, KK05, Kim06, KF03, Li05b, Li08a, Li08b, Li09, Mur08, RL07, Ryn00, She13, WZ10, XH13, Yao16, ZSB01a, ZV08]. Solvability [HS08]. Solvable [Wat96]. solve [ARPR01]. solver [Ben97, Cai02, YZ13]. solvers [AB97, BBB^{+}16, KP95, Sch15, SW96, vdHdS97]. Solving [Fas99b, HS13, JM00, LM14, PC17, Sch09, AC00, sCLC13, Cul96, Li05b, LSXZ15, LLY12, MS00, TQ17, WHS18, ZT14, ZhXpZ16]. Some [CHM18, DT96, Dou94, GVSLN96, KY01, MN00, CLM96a, CKP13, Dai16, DS01, HX17, Le 00, LT98, LZZ09, LPSSP00, Peñ11, RT14, SW98]. space [BLP02, BKK17, CGM01, FS94, Far10, Fus08a, Fus08b, GG16, HT11, KCW17, KJ09, NV10, Pin04, She13, TZ07, Yao16, ZD16]. spaced [LP14]. Spaces [GH03, AS08, BGM93, CLLS06, CMP07b, CS06, CS03, CG10, DST04, DS00, FY13, GHdN12, GJS14, Han06, HHS04, Jia95, LYY13, LT98, Lut04, Maz99, NSW16, RY16, SS16, Spe17, Sun08, Sun10, Tem01, Tem07, WYW11, ZWDD16, dB07, vdMNS03]. spanned [dB07]. Sparse [BD93, CCH^{+}11, CP15, Fli18, Hem95, WW12, BNR00, BF11, KL15, MR15, NBL11, PGB15, Pla00, SU12, SJJL15, Hem96]. Sparse-grid [Hem95, Hem96]. sparsity [BDMR10, FRT09, GN08, MMP13]. Spatial [CP04, FaKS02, Leo02, PFMS05]. special [BOP^{+}15, CHY10, Tom96, XY12]. Spectral [GMRS97, yGSlX03, JL97, Smo07, ACH10, AHC11, BK16, CJX17, CP14, CP15, CGW13, DY10, Gu17, yGW07, yGjW10, HL06, HX17, HY18, HS16, LY07, LMV11, LWL17, LL07b, MFB^{+}11, Möß10, Mur08, lXyG02, gYyG12, ZP06, Zhu15]. spectrum [FMPS13]. sphere [BC94, DST04, DG13, FM14, FM04, HKR04, HR17, KN14, KNQ16, LW14, LM14, Leo13, Leo02, Li08a, MNW99, MNPW00, MNW04, NW04, Pot04, SW04, SS08, SM99, WS01, Xu04a, Xu04b, zCFX07]. Spheres [Noa02, Fas99a, Kun09, Le 05, Luo99, NSW98]. Spherical [FW96, BMK15, DG16, FM14, HL08, Hes09, HW12, IP15, Le 05, NS04, PTC12]. spheroid [CP15, LST11]. spin [CLM02].

Spline [Jia09a, JF02, AS08, BS05, ČF11, CHRX06, DS00, DS01, DX10, FGS16, FP99, Flo94, GJS14, KK99, KvD01, LW04, Leo02, MS07, MXZ08, MXY13, MH17, Rem12, SS96, Sor18, Spe17, SZ09, Wal95, ZWDD16]. splines [AF96, AHHR16, ANSZ17, BR13, BL99, CM12, CGR13, CKM99, DL98, DZ04, DOZ94, GO05, JKLY13, JL06, Joh13, KLNS95, KL04, KL07, KP04, Kva14, LL98, LS98, LLS95, Leo02, LM06, MR96, MSS02, Maz01, MNW96, NS04, NSS04, NZ04, NSZ04, PTC12, Plo95, PW94, Sch04, SS04, SS14, Str95, War13, Yoo01, ZB99, Zhe06, ZA10]. split [HH15]. splits [AS02]. Splitting [HZHL12, XWL13, HH16, nCV13, ZT14, ZPY15]. splittings [GO95, WMB13]. square [SW10]. Squares [JF02, CHY11, FHN94, LK05, Sto96, Tom96, WK93]. SSOR [Bai99]. stabilisation [Mat09]. Stability [CK03b, DS10, NSW98, ZV08, ARAK09, AV15, But97, DHK98, DP13, ER99, Fus08a, Fus08b, JS95, Kei95, Kun09, Low05, MN15, MB96, MNPR06, PST08, XB05, PS98]. stabilization [HT12, JJLR14, NV10, TL18, TZ07, dFGAJN18]. Stabilized [Ye00, BQRB13, LC09, LC13, MU14, SI15, XY15, ZT15]. Stable [BH15, LP04, XL10, GJS14, She13, TGA96, Ye04, dFGAJN18]. Stair [lWyG06]. standard [BDD06]. star [GMB97]. starting [CLM03]. state [Ang00, DMMS11, ER99, HO15]. Stationary [LLY06, Bre95, CGR11, JKLY13, LC09, MS12]. statistical [BEG17]. statistics [HSV09]. steady [Ang00, DMMS11, ER99]. steady-state [Ang00, DMMS11]. steel [GN13]. Stefan [HXC10]. Steklov [KP98, LLZ10]. step [BS00, CLM03, HFH15, LSXZ15]. step-size [BS00]. stepping [KC16, VK96, ZPY15]. Stepsize [CHMR97, Hal95, Wil98]. stiff [Ben97, BS97a, BS00, But97, EJ97]. stochastic [CCG10, CJX17, MM00, SU12, WSM13].

Stokes [Ang00, APV14, Bre95, BR15, CHM18, CCG10, CLL08, CL16, CT17a, CGS93, FF95, HL06, JJLR14, JK08, KL00, LC09, LG13, LC13, fLL05, LY08, MRH15, MY17, PCM^{+}96, WY16, XY15, ZVW95, ZT15, dFGAJN18]. Stokes/Darcy [CHM18]. Straight [San15, DD94]. Strang [GW17, WL05]. strategies [MN15]. strategy [ARPR01, CGR13, GPT17, Hal95, LLY12]. streamline [KY01]. stress [GG16, Lee16]. Strictly [Luo99, Pin04, SM99, GP95]. strip [lXyG02]. strong [DGMM16]. Strongly [GJS14, OS09]. structural [ARPR01, CCN^{+}16]. structural-acoustic [ARPR01]. structure [AB02, Wat96]. Structured [BKN11, GN08, Kim06, MMP13, ZPR03]. structures [JN18, KC16]. Studies [LZZ09]. study [EHV17]. Sturm [GM17]. Subdivision [CD15, Gon93, JZ04, Jia95, LP10, ADL11, CJ07, CGR11, CGR13, CCS16, CGRS15, DM93, De 06, DLUS17, DS13, FBCR13, GO05, GMS12, JKLY13, LLY06, Lev99, MS12, Moo17, MH17, Nie07, Osw04, PK94, Pra98, PR99, RS01, WYW11, Zhe06, ZA10]. subdomain [CHS17]. subgradient [GH16]. subsonic [KvL95]. subspace [BPK96, GO95, XWL13, ZVW95]. subspaces [Atr12, HLT11, LO08, LL11, SZ09]. Sufficient [FS07, MNPR06]. sum [CMP07a, Han10, JGW15, TZ05, WFL02]. sum-of-exponentials [JGW15]. summaries [Joh13]. summation [Pré94]. sums [Tor16]. Sumudu [PM17]. sup [dFGAJN18]. Super [LL13, LT12]. Supercloseness [ZL17]. superconductors [HX02]. Superconvergence [CZZ17, HSY16, LY08, Zha14, Gom95, LLZ10, LL12b, MNC16, ZSB01a]. superposition [Nee96]. SUPG [Kno09]. Support [EPP00, CX06, CKK12, NSS04]. supported [CL16, CKK12, CHSS03, Hub12, JJK13, JS99, Lai06, LLS95, Len96, Wen95, Wu95]. Supports [CTZ02, KL15]. suppression [LO08]. Surface [Kra02, BX94, Ble07, FM14, FHN94, GN13, HFH^{+}07, Jia11, PCM^{+}96, ZL16]. surface/surface [BX94]. Surfaces [JF02, SR02, CW16, CC94, CLM02, CL18, DP13, Flo94, GU02, HFH^{+}07, KK99, KP04, Leo02, LPSSP00, MT98, PGM09, Pra98, PR99, RS01]. survey [CBK01, RT14]. SVM [ZJ06]. sweeping [Sto17]. Sylvester [ZT14]. symbol [GP14, Plo95]. symbols [Nie07]. Symmetric [Han98, Han10, Jia03, Bai99, JKLY13, JS95, Lee16, WYW11]. Symmetric/Antisymmetric [Jia03]. symmetries [GH99]. symmetry [AD08, AG98, Awa13, Jia11]. symplectic [DDP14, Ram95]. synthesis [ZYY16]. System [Tan17, AK15, BR15, GLO10, Gu17, KA97, MRH15, MXO13, MS00, OS09, SC08, vdHdS97]. Systems [CMP03, AZW15, Bai99, BOP^{+}15, Bes03, BKK17, CS93, CEHK18, CTZ02, CR08, CKK16, CHL17, Dad17, DLP98, FE15, Hol01, KKLY10, LZZ09, Liu99, Mar94, MV01, MS98, SXYY06, SW04, Smo07, Tor16, WMY13, WMB13, WOBL17, dH94, dCB07]. systolic [AH96]. Szego [CMP07a].

T [ZWDD16]. T-meshes [ZWDD16]. Table [Mac94]. Table-based [Mac94]. tangentially [MU14]. task [LMWX13]. Taylor [Del17, fLL05]. Tchebycheffian [PW94]. technique [BHT11, CS93, FP99, Gil17, KH05, Ren09, SFDE15]. techniques [Sor18]. tempered [LD16]. tension [CKM99, MR96, ZL16]. Tensioned [LM04]. Tensor [FP98, GO95, CP15, Flo94, KK99]. Tensor-product [FP98, Flo94, KK99]. tenth [TBD94]. tenth- [TBD94]. term [RL07, Tem98]. terms [CC94, Lóp18, XZ03]. tessellations [GS10]. tesseral [FM04]. tests [Mac94]. tetrahedral [FPT06, Kun01, SS04]. tetris [CP14]. their [ByLl01, CLLS06, CSS98, yGS08, HLT11, JGW15, NSW07, PŠWX03, RSA14, She08, Wat96]. themes [Hof06]. theorem [CKBP11]. theorems [GP14]. theoretic [RJ00]. theoretically [Wie97]. theory [AM14, LL99, LS05, MNPR06, Ock96, dLdDSM07]. thermal [TBD94]. thermoelectrical [BGMS07]. thin [DD94, FP99, Joh13, Ryn00, Yoo01, ZB99]. thin-plate [Joh13, Yoo01]. third [PM17, SM00]. three [BS05, HSY16, LLS96, LL98, LG13, LL07b, MSS02, RL07, WHS18]. three-dimensional [LG13, WHS18]. three-direction [BS05]. thresholding [AL13]. Tight [BF03, BS06, CK03a, Jia03, PŠWX03, CCSS09, CHSS03, FWW06, Jia09b, LMO14]. Tikhonov [HT11]. tilings [CTZ02, FPT06]. Time [DD94, TB16, VK96, ALZ02, ANSZ17, Awa15, BF11, GG06, GPG15, HY18, KC16, Lem09, LZZ09, LHY08, LL07b, MNC16, Onc14, PM17, Rau05, RST^{+}14, gTpM02, TGA96, VW02, Xu14, ZPY15, ZW12, dFGAJN18, HL06]. time-concentrated [GG06]. Time-dependent [TB16, LZZ09, TGA96, dFGAJN18]. time-domain [VW02]. time-fractional [PM17]. time-frequency [Onc14, RST^{+}14]. time-harmonic [HY18, LHY08]. Time-marching [DD94]. Time-stepping [VK96]. timer [ALZ02]. TM [LN14]. Tocher [LL98]. Toeplitz [WLW16]. tolerance [CHMR97]. tomographic [GL13]. tomography [HLN12]. tool [Sch15]. Top [DMMS11]. Top-level [DMMS11]. topography [CT17b, IL16]. topological [Cho95, HLN12]. topology [HFH^{+}07]. Toric [Kra02]. torus [Pot04]. Total [BL09, LM06, BCE^{+}09, DGK15, DT12, FRT09, JZ10, KPY09, Maz01, PZ16, ZPR03]. totally [GP95]. traces [HJH12]. tractability [NSW16]. Training [BE00, WFL02]. transaction [ALZ02]. transfer [HXC10, PCM^{+}96]. transform [ANSZ17, BMB14, CCH^{+}08, DLP98, FW96, LL12a, MXY13, PM17, RST^{+}14, SL05, SS13, TZ03, SF14]. transformation [SFDE15]. transformations [FWW06]. transforms [ADL11, DMT03, Dam07, HKR04, Kei95, LM14, PST08, Ste98]. Transient [BGMS07, Awa15, BMSR^{+}16, CT17a, LC13, VW05]. translates [AM14, CL12, LL99]. translation [LSLS14, MXZ08, vdMNS03]. transmission [DRS08, HJH12]. transport [Moo17]. trapezoidal [Pla12]. traps [ZD17]. traveling [Wan08]. travelling [She13]. treatment [Bes03, Hau97, Hol01]. Tree [BDKY02]. Tresca [HKRS14]. triadic [ZA10]. triangle [AS02, yGW07]. Triangular [LLS94, CLZW13, DP13, SL05, ZL17]. triangulated [NSZ04]. triangulations [FPT06, NS04, Rem12]. Tribute [Sil99]. trigonometric [DS13, KLNS95, LN14, LWY06, Pot04, Rau05, RSA14]. trimmed [HFH^{+}07]. triple [WHS18]. triply [Tor16]. trivariate [AS08, LL04]. troubled [ZQ13, ZG16]. troubled-cell [ZQ13, ZG16]. truncated [CGW13]. tuning [Nes16]. TV [MSXZ13, RWT13]. twelfth [TBD94]. twelfth-order [TBD94]. Two [ASS10, CLM03, CCH^{+}08, Jia03, Plo95, BNP14, Bre95, Bün11, Che02, CS94, Dad17, DLS14, FP99, yGmW98, HKRS14, HT12, HFH15, JN18, KS17, KF03, LD07, MR15, MR96, OS09, OP11, RH15, SM00, Sun14, XZ01, ZQ13]. two-body [HKRS14]. two-cell [Dad17]. Two-dimensional [CCH^{+}08, JN18, KF03, MR15, SM00, ZQ13]. two-grid [BNP14, XZ01]. two-level [Bre95, HT12]. two-point [yGmW98, KS17, MR96]. Two-scale [Plo95, CS94, RH15]. Two-step [CLM03, HFH15]. type [Awa15, BR07, BS97b, BS97a, BS00, Chr15, CKP13, CSWP99, CN96, Cum96, CT17b, Dar03, DHO16, DOZ94, DY01, Fra99, GO95, JMO00, KL07, LLS94, LLS96, LT12, LL07b, MSZ06, Maz98, PSNP11, RWT13, Sab04, SS04, SS96, WZL17, Xu14, Yan01, Yan03, YZ13]. type-4 [SS04]. types [LL98].

ultra [Jon17]. ultra-precise [Jon17]. unbounded [GZ13, LHY08]. Uncertainty [GY00, LSLS14, MSST14, RT14]. Unconditional [CLC16]. unevenly [LP14]. unified [Bar03, Grä12, Lee16, Wal95]. Uniform [Che06, CLZW13, DGS18, LW03, TZ07, Atr12, ANSZ17, DG16, DY01, HO17, Lev99, OP11, Rem12, ZD16]. uniform-grid [DG16]. Uniformity [LW03]. Uniformly [KXZ04]. unimodular [BD10]. unions [Grö03]. Uniqueness [GP14, ZYY16]. unisolvent [Cal05]. unit [CWLH95, GVSLN96, LM14, Leo13, Li08a, Li08b, Xu04a, Xu04b, zCFX07]. unitarily [vdMNS03]. unitary [DLP98, GLT93]. unity [CM12, HL05, NSS04, ZSHZ15]. univariate [Fra99, MH17]. University [New96]. unknown [Wie15, ZPR03]. unlearning [HL96]. unorganised [BLP02]. unstable [She13]. unsteady [BBB^{+}16]. unstructured [Hag98, Wie97]. unsymmetric [LLS09]. updates [AZW15]. upwind [Ang00, Wie97]. upwind-discretizations [Ang00]. use [RSA14]. Using [Cas07, Lev99, Zoz15b, AH96, BGH15, CGW13, Dou94, EHV17, FF95, FGS16, FHN94, FS98, GO05, yGSlX03, HFH^{+}07, IUV17, Le 05, LW14, LLY06, Li05a, LZY98, LWL17, MNW99, MNW04, PTC12, RWT13, Sch09, SS08, WL12, WZL17, YLBL12, Yoo01, ZB99, ZPR03, ZT15, Zhu15].

validated [HBM03]. validation [HL96]. value [ABI15, CWX12, FK98, FK10, GM14, yGmW98, HH15, KS17, MR96, RGB13, Rip99, Sch11b, vdH93]. valued [Füh08, Fus08a, Fus08b, GMB97, HHS04, IB17, Low05, XKY15]. values [CLM03, MC10]. Vandermonde [Maz98, Tom96]. Vanishing [Füh16, BP93, CHR00, CHSS03, CaL06, GT14]. Variable [BR13, Bra01, BS00, CT17b]. variables [GS00, MS12, Xu00]. Variants [HH15]. variation [BL09, BCE^{+}09, CGP95, DGK15, DT12, FRT09, JZ10, KPY09, PZ16, QW11]. variation-penalized [BL09]. Variational [CW16, JK08, KL16, LB93, OBS15, WOBL17, dLdDSM07]. variations [Hof06]. varying [Xia13]. Vector [ABD08, EPP00, AC00, CX06, Füh08, GO05, GMB97, Han06, Rob98, SM00, Sal98, Sor18]. vector-valued [Füh08, GMB97]. vectorizing [Dou94]. Vectors [HM03]. version [GH06, Jan98, WY17]. versus [Cul96]. vertex [RJ00]. vertex-modified [RJ00]. via [BRSV15, BJX09, CP14, Dam07, FY13, FGO14, GS10, GL13, HS13, Kos15, LM04, Las16, LM14, MS04, Moo17, NP18, PR15, SF14, WFL02, ZA10]. vibration [ARPR01]. virtual [MU14]. viscoelastic [ZPY15]. viscosity [GT14, ZPY15]. viscosity-splitting [ZPY15]. Vlasov [AK15]. Voigt [BNP14]. Volterra [BJX09, Gu17, HS16, Pla12, SW96, SC08, WY17, XH13]. volume [Ano16, CDP03, CLZ02, CWX12, CEHK18, Hem95, Hem96, HR17, KH05, KHM16, LN14, LC09, LC13, LL10, LL12b, Noe95, Zha14, ZT15]. Voronoi [GS10].

Wachspress [DOZ94, FK10, KB15]. Warping [ZB99]. water [CT17b, vdHS00]. watermarking [Ren09]. wave [ADG17, CR08, GT14, HY18, HX02, JRS15, NV10, She13, TZ07, ZG16]. Waveform [vdHv97, HJJV97]. Wavelet [DPS93, DKLT93, GLT93, HR02, JL06, TSZ98, ANSZ17, BS06, BDMR10, CP99, CMX02, CP03, CKK16, DMT03, Dũn09, Ehr00, EGL13, FOW14, FW96, Füh16, Gom95, Kei95, KKLY10, LT98, PL05, PST08, RSZ11, RWT13, XZ03, Zha05]. Wavelet-based [TSZ98]. wavelet-Galerkin [Gom95]. Wavelets [Pla95, CT05, ČF11, CHR00, CS94, CSW96, CaL06, DKB99, FM04, GY00, Han98, Han10, JS99, Jia09a, Jia11, KKL03, Lai06, PŠWX03, PST95, Rau05, Str13, Urb95, ZhXpZ16]. waves [ZP06]. wavetrain [She13]. Weak [GN01, Tem00, Flo94, LLC16, Tem02, WY16]. weakly [Awa13, HS16, Lee16, Pla12]. WEB [AHHR16, HAS05]. Web-method [HAS05]. WEB-splines [AHHR16]. weight [Xu98]. Weighted [CP14, Dam07, Kva14, MS16]. well [CT17b]. well-balanced [CT17b]. Wendland [CSW14, Sch11a, War13]. WENO [JN18, LQ14, TQ17]. which [MXZ08]. while [Dou94]. Whitney [SS96]. Whittle [BRS13]. whole [BKK17]. wideband [DMT03, DG13]. widths [DKLT93]. window [TZ03]. windowed [SS13]. windows [CKK12, ZZ09]. wire [DD94, Ryn00]. wirebasket [FH18]. within [Jon17]. without [KXZ04]. work [CP07a].

X [GL13]. X-ray [GL13].

Young [MRS03]. Yukawa [KNQ16].

Zero [JW95, BD10, Flo96b, GG16]. Zerofree [JW95]. zonal [FM04, MNW99]. Zooming [LSW17].

References

Arnold:2015:MFE

[AAQ15] Douglas N. Arnold, Gerard Awanou, and Weifeng Qiu. Mixed finite elements for elasticity on quadrilateral meshes. Advances in Computational Mathematics, 41(3):553–572, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9376-x.

Amodio:1997:POS

[AB97] Pierluigi Amodio and Luigi Brugnano. Parallel ODE solvers based on block BVMs. Advances in Computational Mathematics, 7(1–2):5–26, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018926314954. Par-

allel methods for ODEs.

Ammari:2002:MEP

[AB02] Habib Ammari and Gang Bao. Maxwell’s equations in a perturbed periodic structure. Advances in Computational Mathematics, 16(2–3):99–112, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014402300371. Mod-

eling and computation in optics and electromagnetics.

Anitescu:2011:APG

[AB11] C. Anitescu and U. Banerjee. Approximation properties of the generalized finite element method. Advances in Computational Mathematics, 34(4):369–390, May 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9159-y.

Antonelli:2014:GFC

[ABC14] Michele Antonelli, Carolina Vittoria Beccari, and Giulio Casciola. A general framework for the construction of piecewisepolynomial local interpolants of minimum degree. Advances in Computational Mathematics, 40(4):945–976, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9335-y.

Arandiga:2008:VCA

[ABD08] Francesc Aràndiga, Antonio Baeza, and Rosa Donat. Vector cell-average multiresolution based on Hermite interpolation. Advances in Computational Mathematics, 28(1):1–22, January 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-9007-7.

Amodio:2015:ECM

[ABI15] Pierluigi Amodio, Luigi Brugnano, and Felice Iavernaro. Energy-conserving methods for Hamiltonian boundary value problems and applications in astrodynamics. Advances in Computational Mathematics, 41(4):881–905, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9390-z.

Arandiga:2013:NCN

[ABY13] Francesc Aràndiga, Antonio Baeza, and Dionisio F. Yáñez. A new class of non-linear monotone Hermite interpolants. Advances in Computational Mathematics, 39(2):289–309, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-012-9280-1.

Atkinson:2000:FMV

[AC00] Kendall Atkinson and David Da-Kwun Chien. A fast matrix– vector multiplication method for solving the radiosity equation. Advances in Computational Mathematics, 12(2– 3):151–174, February 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018961001209.

Alves:2005:NMF

[AC05] C. J. S. Alves and C. S. Chen. A new method of fundamental solutions applied to nonhomogeneous elliptic problems. Advances in Computational Mathematics, 23(1–2):125–142, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1833-5.

Atkinson:2010:SME

[ACH10] Kendall Atkinson, David Chien, and Olaf Hansen. A spectral method for elliptic equations: the Dirichlet problem. Advances in Computational Mathematics, 33(2):169–189, August 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9125-8.

Aimi:2008:RMN

[AD08] A. Aimi and M. Diligenti. Restriction matrices for numerically exploiting symmetry. Advances in Computational Mathematics, 28(3):201–235, April 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-006-9019-y.

Aimi:2017:EBF

[ADG17] A. Aimi, M. Diligenti, and C. Guardasoni. Energetic BEM– FEM coupling for the numerical solution of the damped wave equation. Advances in Computational Mathematics, 43(3):627–651, June 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9500-1.

Amat:2011:ACN

[ADL11] S. Amat, K. Dadourian, and J. Liandrat. Analysis of a class of nonlinear subdivision schemes and associated multiresolution transforms. Advances in Computational Mathematics, 34(3):253–277, April 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-010-9151-6.

Albrecht:1996:CPH

[AF96] Gudrun Albrecht and Rida T. Farouki. Construction of C^{2} Pythagorean-hodograph interpolating splines by the homotopy method. Advances in Computational Mathematics, 5 (1):417–442, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124754.

Allgower:1998:NES

[AG98] Eugene L. Allgower and Kurt Georg. Numerical exploitation of symmetry in integral equations. Advances in Computational Mathematics, 9 (1–2):1–20, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018929204901. Nu-

merical treatment of boundary integral equations.

Anderson:1996:PFB

[AH96] I. J. Anderson and S. K. Harbour. Parallel factorization of banded linear matrices using a systolic array processor. Advances in Computational Mathematics, 5(1):1–14, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124732.

Atkinson:2011:SME

[AHC11] Kendall Atkinson, Olaf Hansen, and David Chien. A spectral method for elliptic equations: the Neumann problem. Advances in Computational Mathematics, 34(3):295–317, April 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9154-3.

Apprich:2016:CWS

[AHHR16] Christian Apprich, Klaus Höllig, Jörg Hörner, and Ulrich Reif. Collocation with WEB-splines. Advances in Computational Mathematics, 42(4):823–842, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9444-x.

Ainsworth:2001:PEE

[AK01] Mark Ainsworth and Donald W. Kelly. A posteriori error estimators and adaptivity for finite element approximation of the non-homogeneous Dirichlet problem. Advances in Computational Mathematics, 15(1–4):3–23, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014240508621.

Asadzadeh:2015:CAB

[AK15] Mohammad Asadzadeh and Piotr Kowalczyk. Convergence analysis for backward-Euler and mixed discontinuous Galerkin methods for the Vlasov–Poisson system. Advances in Computational Mathematics, 41(4):833–852, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9388-6.

afKlinteberg:2017:EEQ

[aKT17] Ludvig af Klinteberg and Anna-Karin Tornberg. Error estimation for quadrature by expansion in layer potential evaluation. Advances in Computational Mathematics, 43 (1):195–234, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9484-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        016-9484-x.pdf.

Arnold:1997:IEL

[AL97] Douglas N. Arnold and Xiaobo Liu. Interior estimates for a low order finite element method for the Reissner– Mindlin plate model. Advances in Computational Mathematics, 7(3):337–360, May 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018907205385.

Amat:2013:NTM

[AL13] Sergio Amat and Jacques Liandrat. Nonlinear thresholding of multiresolution decompositions adapted to the presence of discontinuities. Advances in Computational Mathematics, 38(1):133–146, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9231-2.

Alben:2015:BBG

[Alb15] Silas Alben. Bending of bilayers with general initial shapes. Advances in Computational Mathematics, 41(1):1–22, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9347-2.

Albeverio:2002:CTF

[ALZ02] Sergio Albeverio, Lan-Jun Lao, and Xue-Lei Zhao. A continuous time financial market with a Poisson process as transaction timer. Advances in Computational Mathematics, 16(4):305–330, May 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014566730969.

Alzer:2010:EFI

[Alz10]  Horst Alzer.  Error func-

tion inequalities. Advances in Computational Mathematics, 33(3):349–379, October 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9139-2.

Arteaga:2014:DFS

[AM14] Cristian Arteaga and Isabel Marrero. Direct form seminorms arising in the theory of interpolation by Hankel translates of a basis function. Advances in Computational Mathematics, 40(1):167–183, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9303-6.

Angermann:2000:EAU

[Ang00] Lutz Angermann. Error analysis of upwind-discretizations for the steady-state incompressible Navier–Stokes equations. Advances in Computational Mathematics, 13(2):167–198, May 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018902210295.

Anonymous:1993:MJG

[Ano93]  Anonymous.  In memoriam

John Gregory. Advances in Computational Mathematics, 1(1):iii, February 1993.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02070818.

Anonymous:1995:GE

[Ano95] Anonymous. The Guest Editor. Advances in Computational Mathematics, 4(1–2):i, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02123469.

Anonymous:2000:AI

[Ano00a] Anonymous. Author index. Advances in Computational Mathematics, 13(4):417–418, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016615818384.

Anonymous:2000:P

[Ano00b] Anonymous. Preface. Advances in Computational Mathematics, 12(4):i–ii, March 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018925420773.

Anonymous:2001:AI

[Ano01a] Anonymous. Author index. Advances in Computational Mathematics, 14(4):393–394, May 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016819732496.

Anonymous:2001:PEE

[Ano01b] Anonymous. A posteriori error estimation and adaptive computational methods. Advances in Computational Mathematics, 15(1–4):1–2, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016858615238.

Anonymous:2001:PA

[Ano01c] Anonymous. Publisher’s announcement. Advances in Computational Mathematics, 14(1): 1–6, January 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1023/A:1016633820248.

Anonymous:2002:AIa

[Ano02a] Anonymous. Author index. Advances in Computational Mathematics, 16(4):389–390, May 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016822023144.

Anonymous:2002:AIb

[Ano02b] Anonymous. Author index. Advances in Computational Mathematics, 17(4):397–398, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016232931523.

Anonymous:2002:MCO

[Ano02c]  Anonymous.  Modeling and

computation in optics and electromagnetics. Advances in Computational Mathematics, 16(2–3):97–98, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016809007214.

Anonymous:2002:PAG

[Ano02d]  Anonymous.  Preface; ad-

vances in geometrical algorithms and representations. Advances in Computational Mathematics, 17(1–2):1–3, July 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016806617055.

Anonymous:2003:AIa

[Ano03a] Anonymous. Author index. Advances in Computational Mathematics, 18(2–4):431–432, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021307330201.

Anonymous:2003:AIb

[Ano03b] Anonymous. Author index. Advances in Computational Mathematics, 19(4):413–414, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1024388324172.

Anonymous:2003:PEA

[Ano03c]  Anonymous.  Past Editors

of Advances in Computational Mathematics. Advances in Computational Mathematics, 19(1–3):337, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1022848030422.

Anonymous:2004:AIa

[Ano04a] Anonymous. Author index. Advances in Computational Mathematics, 20(4):427–428, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1027303623367.

Anonymous:2004:AIb

[Ano04b] Anonymous. Author index. Advances in Computational Mathematics, 21(3–4):449–450, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/B:ACOM.0000032143.
                                        19990.2b.

Anonymous:2004:E

[Ano04c] Anonymous. Errata. Advances in Computational Mathematics, 21(3–4):445, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/B:ACOM.0000032141.
                                        51652.c6;   http://link.springer.

com/content/pdf/10.1023/B:
ACOM.0000032141.51652.c6.
pdf.

Anonymous:2004:PEAa

[Ano04d] Anonymous. Past Editors of Advances in Computational Mathematics. Advances in Computational Mathematics, 20(4):425, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1027315116562.

Anonymous:2004:PEAb

[Ano04e]  Anonymous.  Past Editors

of Advances in Computational Mathematics. Advances in Computational Mathematics, 21(1–2):219, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000016693.
57844.49.

Anonymous:2004:PEAc

[Ano04f]  Anonymous.  Past Editors

of Advances in Computational Mathematics. Advances in Computational Mathematics, 21(3–4):447, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000032142.
27408.87.

Anonymous:2004:P

[Ano04g] Anonymous. Preface. Advances in Computational Mathematics, 20(1–3):1–3, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1025814926647.

Anonymous:2005:PEAa

[Ano05a]  Anonymous.  Past editors

of Advances in Computational Mathematics. Advances in Computational Mathematics, 22(1):101, January 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-005-5323-1.

Anonymous:2005:PEAb

[Ano05b]  Anonymous.  Past editors

of Advances in Computational Mathematics. Advances in Computational Mathematics, 22(2):199, February 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-6475-0.

Anonymous:2005:P

[Ano05c] Anonymous. Preface. Advances in Computational Mathematics, 23(1–2):1–3, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-2232-7.

Anonymous:2006:P

[Ano06] Anonymous. Preface. Advances in Computational Mathematics, 24(1–4):1–3, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7205-3.

Anonymous:2008:E

[Ano08] Anonymous. Errata. Advances in Computational Mathematics, 29(1):91–100, July 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9090-7.

Anonymous:2016:EAC

[Ano16] Anonymous. Erratum to: Advances in Computational Mathematics, volume 41 December 2015, issue 6, pages 1391– 1641 [MR3428571; MR3428572; MR3428573; MR3428574; MR3428575; MR3428576; MR3428577; MR3428578; MR3428579; MR3428580]. Advances in Computational Mathematics, 42(4):1013–1014, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9464-1;
 http://link.springer.com/
content/pdf/10.1007/s10444-
016-9464-1.pdf. Due to an

oversight during issue building, the page numbers of issue 6 overlaps issue 5. Issue 5 start and end page is 955–1389, while issue 6 start and end page is 955– 1205. Corrected page numbers are supplied here.

Averbuch:2017:LCS

[ANSZ17] Amir Averbuch, Pekka Neittaanmäki, Etay Shefi, and Valery Zheludev. Local cubic splines on non-uniform grids and real-time computation of wavelet transform. Advances in Computational Mathematics, 43(4):733–758, August 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9504-x.

Antunes:2018:RIC

[Ant18] Pedro R. S. Antunes. Reducing the ill conditioning in the method of fundamental solutions. Advances in Computational Mathematics, 44 (1):351–365, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9548-6.

Araya:2014:ARL

[APV14] Rodolfo Araya, Abner H. Poza, and Frédéric Valentin. An adaptive residual local projection finite element method for the Navier–Stokes equations. Advances in Computational Mathematics, 40(5– 6):1093–1119, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9343-6.

Acosta-Reyes:2009:SSR

[ARAK09] Ernesto Acosta-Reyes, Akram Aldroubi, and Ilya Krishtal. On stability of samplingreconstruction models. Advances in Computational Mathematics, 31(1–3):5–34, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9083-6.

Alonso:2001:PEE

[ARPR01] Ana Alonso, Anahí Dello Russo, Claudio Padra, and Rodolfo Rodríguez. A posteriori error estimates and a local refinement strategy for a finite element method to solve structural-acoustic vibration problems. Advances in Computational Mathematics, 15 (1–4):25–59, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014243118190. A

posteriori error estimation and adaptive computational methods.

Alfeld:2002:SME

[AS02] Peter Alfeld and Larry L.

Schumaker. Smooth macroelements based on Powell– Sabin triangle splits. Advances in Computational Mathematics, 16(1):29–46, January 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1014299228104.

Aldroubi:2003:P

[AS03] Akram Aldroubi and Qiyu Sun. Preface. Advances in Computational Mathematics, 18 (2–4):91–92, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1021340203871.

Atluri:2005:BMD

[AS05] Satya N. Atluri and Shengping Shen. The basis of meshless domain discretization: the meshless local Petrov–Galerkin (MLPG) method. Advances in Computational Mathematics, 23(1–2):73–93, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1813-9.

Alfeld:2008:BDT

[AS08] Peter Alfeld and Larry L. Schumaker. Bounds on the dimensions of trivariate spline spaces. Advances in Computational Mathematics, 29 (4):315–335, November 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9051-6.

Alfeld:2010:TCM

[ASS10] Peter Alfeld, Larry L. Schumaker, and Tatyana Sorokina.

Two condensed macro-elements with full approximation power. Advances in Computational Mathematics, 32(4):381–391, May 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9110-7.

Ali:2017:RBM

[ASU17] Mazen Ali, Kristina Steih, and Karsten Urban. Reduced basis methods with adaptive snapshot computations. Advances in Computational Mathematics, 43(2):257–294, April 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9485-9.

Atreas:2012:CNU

[Atr12] Nikolaos D. Atreas. On a class of non-uniform average sampling expansions and partial reconstruction in subspaces of L_{2}(R). Advances in Computational Mathematics, 36(1):21–38, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9177-4.

Alla:2015:ASP

[AV15] Alessandro Alla and Stefan Volkwein. Asymptotic stability of POD based model predictive control for a semilinear parabolic PDE. Advances in Computational Mathematics, 41(5):1073–1102, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9381-0.

Awanou:2013:RME

[Awa13] Gerard Awanou. Rectangular mixed elements for elasticity with weakly imposed symmetry condition. Advances in Computational Mathematics, 38(2):351–367, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9240-1.

Awanou:2015:PTC

[Awa15] Gerard Awanou. Pseudo transient continuation and time marching methods for Monge– Ampère type equations. Advances in Computational Mathematics, 41(4):907–935, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9391-y.

Atkinson:1998:I

[AX98] Kendall Atkinson and Yuesheng Xu. Introduction. Advances in Computational Mathematics, 9(1–2):0, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018968919014.

Amsallem:2015:FLR

[AZW15] David Amsallem, Matthew J.

Zahr, and Kyle Washabaugh. Fast local reduced basis updates for the efficient reduction of nonlinear systems with hyper-reduction. Advances in Computational Mathematics, 41 (5):1187–1230, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9409-0.

Borzi:2015:SOA

[BA15] Alfio Borzì and Sergio González Andrade. Second-order approximation and fast multigrid solution of parabolic bilinear optimization problems. Advances in Computational Mathematics, 41(2):457–488, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9369-9.

Bai:1999:CMB

[Bai99] Zhong-Zhi Bai. A class of modified block SSOR preconditioners for symmetric positive definite systems of linear equations. Advances in Computational Mathematics, 10(2):169–186, February 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018974514896.

Barry:1994:I

[Bar94] Phillip Barry. Introduction. Advances in Computational Mathematics, 2(1):ix–x, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02519032.

Barrio:2003:URE

[Bar03] Roberto Barrio. A unified rounding error bound for polynomial evaluation. Advances in Computational Mathematics, 19(4):385–399, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1024203520270.

Beatson:2007:MFI

[BB07] R. K. Beatson and H.-Q. Q. Bui. Mollification formulas and implicit smoothing. Advances in Computational Mathematics, 27(2):125–149, August 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-005-7512-3.

Blom:2016:CRE

[BBB^{+}16] David S. Blom, Philipp Birken, Hester Bijl, Fleur Kessels, Andreas Meister, and Alexander H. van Zuijlen. A comparison of Rosenbrock and ES-DIRK methods combined with iterative solvers for unsteady compressible flows. Advances in Computational Mathematics, 42(6):1401–1426, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9468-x;
 http://link.springer.com/
content/pdf/10.1007/s10444-
016-9468-x.pdf.

Bonnet-Bendhia:2002:DAG

[BBR02] Anne-Sophie Bonnet-Bendhia and Karim Ramdani. Diffraction by an acoustic grating perturbed by a bounded obstacle. Advances in Computational Mathematics, 16(2–3):113–138, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014437813575. Mod-

eling and computation in optics and electromagnetics.

Brunnett:1994:ECS

[BC94] Guido Brunnett and Peter E.

Crouch. Elastic curves on the sphere. Advances in Computational Mathematics, 2(1):23– 40, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/BF02519034.

Becuwe:1999:FPM

[BC99] Stefan Becuwe and Annie Cuyt.

On the Froissart phenomenon in multivariate homogeneous Padé approximation. Advances in Computational Mathematics, 11(1):21–40, July 1999.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018911623074.

Bollt:2009:GAI

[BCE^{+}09] Erik M. Bollt, Rick Chartrand, Selim Esedoğlu, Pete Schultz, and Kevin R. Vixie. Graduated adaptive image denoising: local compromise between total variation and isotropic diffusion. Advances in Computational Mathematics, 31(1–3):61–85, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9082-7.

Balan:2003:DEF

[BCHL03] Radu Balan, Peter G. Casazza, Christopher Heil, and Zeph Landau. Deficits and excesses of frames. Advances in Computational Mathematics, 18(2–4):93–116, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1021360227672. Frames.

Beatson:1999:FFR

[BCM99] R. K. Beatson, J. B. Cherrie, and C. T. Mouat. Fast fitting of radial basis functions: methods based on preconditioned GMRES iteration. Advances in Computational Mathematics, 11(2–3):253–270, November 1999. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018932227617. Ra-

dial basis functions and their applications.

Borwein:2005:ECG

[BCZ05] P. B. Borwein, A. Cuyt, and P. Zhou. Explicit construction of general multivariate Padé approximants to an Appell function. Advances in Computational Mathematics, 22(3):249–273, April 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-2600-8.

Bank:1993:SMM

[BD93] Randolph E. Bank and Craig C.

Douglas. Sparse matrix multiplication package (SMMP). Advances in Computational Mathematics, 1(1):127–137, February 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02070824.

Benedetto:2010:CIU

[BD10] John J. Benedetto and Somantika Datta. Construction of infinite unimodular sequences with zero autocorrelation. Advances in Computational Mathematics, 32(2):191–207, February 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9100-9.

Barinka:2006:AAO

[BDD06] Arne Barinka, Stephan Dahlke, and Wolfgang Dahmen. Adaptive application of operators in standard representation. Advances in Computational Mathematics, 24(1–4):5–34, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7631-2.

Baraniuk:2002:NBT

[BDKY02] R. G. Baraniuk, R. A. DeVore, G. Kyriazis, and X. M. Yu. Near best tree approximation. Advances in Computational Mathematics, 16(4):357–373, May 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1014554317692.

Bonesky:2010:AWM

[BDMR10] Thomas Bonesky, Stephan Dahlke, Peter Maass, and Thorsten Raasch. Adaptive wavelet methods and sparsity reconstruction for inverse heat conduction problems. Advances in Computational Mathematics, 33(4):385–411, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-010-9147-2.

Burger:2000:TNN

[BE00] Martin Burger and Heinz W.

Engl. Training neural networks with noisy data as an ill-posed problem. Advances in Computational Mathematics, 13(4):335–354, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016641629556.

Bachoc:2017:OCL

[BEG17] François Bachoc, Martin Ehler, and Manuel Gräf. Optimal configurations of lines and a statistical application. Advances in Computational Mathematics, 43(1):113–126, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9478-8;
http://link.springer.com/
content/pdf/10.1007/s10444-
016-9478-8.pdf.

Bendtsen:1997:PSO

[Ben97]  Claus Bendtsen.  A paral-

lel stiff ODE solver based on MIRKs. Advances in Computational Mathematics, 7(1–2):27–36, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018978331792. Par-

allel methods for ODEs.

Bramble:1996:AMA

[BEPS96] James H. Bramble, Richard E. Ewing, Joseph E. Pasciak, and Jian Shen. The analysis of multigrid algorithms for cell centered finite difference methods. Advances in Computational Mathematics, 5(1):15–29, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124733.

Bespalov:2003:OSS

[Bes03]  Alexei Bespalov.  Orthogo-

nal systems of singular functions and numerical treatment of problems with degeneration of data. Advances in Computational Mathematics, 19(1–3):159–182, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1022862704316. Chal-

lenges in computational mathematics (Pohang, 2001).

Bull:1994:PGA

[BF94] J. M. Bull and T. L. Freeman. Parallel globally adaptive quadrature on the KSR-1. Advances in Computational Mathematics, 2(4):357–373, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02521604.

Benedetto:2003:FNT

[BF03] John J. Benedetto and Matthew Fickus. Finite normalized tight frames. Advances in Computational Mathematics, 18 (2–4):357–385, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021323312367. Frames.

Benner:2011:NSL

[BF11] Peter Benner and Heike Faßbender. On the numerical solution of large-scale sparse discrete-time Riccati equations. Advances in Computational Mathematics, 35(2– 4):119–147, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9174-7.

Benner:2015:PMO

[BGH15] Peter Benner, Sara Grundel, and Nils Hornung. Parametric model order reduction with a small H_{∈}-error using radial basis functions. Advances in Computational Mathematics, 41 (5):1231–1253, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9410-7.

Barry:1993:KIA

[BGM93] Phillip J. Barry, Ronald N.

Goldman, and Charles A. Micchelli. Knot insertion algorithms for piecewise polynomial spaces determined by connection matrices. Advances in Computational Mathematics, 1(2):139–171, June 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02071383.

Bermudez:2007:TNS

[BGMS07] A. Bermúdez, D. Gómez, M. C. Muñiz, and P. Salgado. Transient numerical simulation of a thermoelectrical problem in cylindrical induction heating furnaces. Advances in Computational Mathematics, 26(1–3):39–62, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7470-9.

Brander:2017:APE

[BGN17] David Brander, Jens Gravesen, and Toke Bjerge Nørbjerg. Approximation by planar elastic curves. Advances in Computational Mathematics, 43(1):25–43, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9474-z.

Borm:2002:MCA

[BH02] S. Börm and R. Hiptmair. Multigrid computation of axisymmetric electromagnetic fields.

Advances in Computational Mathematics, 16(4):331–356, May 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014533409747.

Bodmann:2015:SPR

[BH15] Bernhard G. Bodmann and Nathaniel Hammen. Stable phase retrieval with lowredundancy frames. Advances in Computational Mathematics, 41(2):317–331, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9359-y.

Bayramyan:2017:NPP

[BH17] Vahagn Bayramyan and Hakop Hakopian. On a new property of n-poised and GC_{n} sets. Advances in Computational Mathematics, 43(3):607–626, June 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9499-3.

Benzi:2011:PTC

[BHT11] Michele Benzi, Eldad Haber, and Lauren Taralli. A preconditioning technique for a class of PDE-constrained optimization problems. Advances in Computational Mathematics, 35(2–4):149–173, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9173-8.

Bini:2011:NCM

[BI11] Dario Andrea Bini and Bruno Iannazzo. A note on computing matrix geometric means. Advances in Computational Mathematics, 35(2–4):175–192, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
010-9165-0.

Brenner:2009:MRV

[BJX09] Marty Brenner, Ying Jiang, and Yuesheng Xu. Multiparameter regularization for Volterra kernel identification via multiscale collocation methods. Advances in Computational Mathematics, 31(4):421–455, November 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9077-4.

Bramble:1996:FEM

[BK96] James H. Bramble and J. Thomas King. A finite element method for interface problems in domains with smooth boundaries and interfaces. Advances in Computational Mathematics, 6 (1):109–138, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127700.

Birem:2016:MSM

[BK16] Mira Birem and Christian Klein.

Multidomain spectral method for Schrödinger equations. Advances in Computational Mathematics, 42(2):395–423, April 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9429-9.

Boulmezaoud:2017:IFE

[BKK17] Tahar Z. Boulmezaoud, Keltoum Kaliche, and Nabil Kerdid. Inverted finite elements for div–curl systems in the whole space. Advances in Computational Mathematics, 43 (6):1469–1489, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9532-1.

Bardsley:2011:SLA

[BKN11] Johnathan M. Bardsley, Sarah Knepper, and James Nagy. Structured linear algebra problems in adaptive optics imaging. Advances in Computational Mathematics, 35(2– 4):103–117, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9172-9.

Bouhamidi:1999:MIS

[BL99] A. Bouhamidi and A. Le Méhauté. Multivariate interpolating (m, l, s)-splines. Advances in Computational Mathematics, 11(4):287–314, December 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018940429434.

Bardsley:2009:TVP

[BL09] Johnathan M. Bardsley and Aaron Luttman. Total variationpenalized Poisson likelihood estimation for ill-posed problems. Advances in Computational Mathematics, 31 (1–3):35–59, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9081-8.

Blesgen:2007:CEE

[Ble07] T. Blesgen. On the competition of elastic energy and surface energy in discrete numerical schemes. Advances in Computational Mathematics, 27(2):179–194, August 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9028-5.

Bercovier:2002:DPP

[BLP02] Michel Bercovier, Moshe Luzon, and Elan Pavlov. Detecting planar patches in an unorganised set of points in space. Advances in Computational Mathematics, 17(1–2):153–166, July 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1015285722950. Ad-

vances in geometrical algorithms and representations.

Bozzini:2002:AIS

[BLS02] Mira Bozzini, Licia Lenarduzzi, and Robert Schaback. Adaptive interpolation by scaled multiquadrics. Advances in Computational Mathematics, 16(4): 375–387, May 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1023/A:1014584220418.

Bilato:2014:AFH

[BMB14] Roberto Bilato, Omar Maj, and Marco Brambilla. An algorithm for fast Hilbert transform of real functions. Advances in Computational Mathematics, 40(5–6):1159–1168, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9345-4.

Behera:2015:MAG

[BMK15] Ratikanta Behera, Mani Mehra, and N. K.-R. Kevlahan. Multilevel approximation of the gradient operator on an adaptive spherical geodesic grid. Advances in Computational Mathematics, 41(3):663–689, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9382-z.

Bermudez:2016:TEC

[BMSR^{+}16] A. Bermúdez, R. Muñoz-Sola, C. Reales, R. Rodríguez, and P. Salgado. A transient eddy current problem on a moving domain. Numerical analysis. Advances in Computational Mathematics, 42(4):757–789, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-015-9441-0.

Boyd:2001:BM

[BMSZ01] Geoff Boyd, Charles A. Micchelli, Gilbert Strang, and Ding-Xuan Zhou. Binomial matrices. Advances in Computational Mathematics, 14(4):379–391, May 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1012207124894.

Bajpai:2014:TGF

[BNP14] Saumya Bajpai, Neela Nataraj, and Amiya K. Pani. On a two-grid finite element scheme for the equations of motion arising in Kelvin–Voigt model. Advances in Computational Mathematics, 40(5– 6):1043–1071, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9340-1.

Barthelmann:2000:HDP

[BNR00] Volker Barthelmann, Erich Novak, and Klaus Ritter. High dimensional polynomial interpolation on sparse grids. Advances in Computational Mathematics, 12(4):273–288, March 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018977404843. Mul-

tivariate polynomial interpolation.

Bochniak:2004:DSF

[Boc04] Marius Bochniak. The dual singular function method for 2D boundary integral equations. Advances in Computational Mathematics, 20(4):293–310, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1027370428226.

Benner:2015:MOR

[BOP^{+}15] Peter Benner, Mario Ohlberger, Anthony T. Patera, Gianluigi Rozza, Danny C. Sorensen, and Karsten Urban. Model order reduction of parameterized systems (MoRePaS): Preface to the special issue of advances in computational mathematics. Advances in Computational Mathematics, 41(5):955–960, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9443-y;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        015-9443-y.pdf.

Baker:1993:PCR

[BP93] Christopher T. H. Baker and Christopher A. H. Paul. Parallel continuous Runge–Kutta methods and vanishing lag delay differential equations. Advances in Computational Mathematics, 1(3):367–394, October 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02072017.

Bru:2011:PNA

[BP11] Rafael Bru and Juan Manuel Peña. Preface: numerical and applied linear algebra. Advances in Computational Mathematics, 35(2–4):99–102, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9170-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        011-9170-y.pdf.

Bermudez:2002:HAC

[BPJ02] A. Bermúdez, D. Gómez Pedreira, and P. Joly. A hybrid approach for the computation of guided modes in integrated optics. Advances in Computational Mathematics, 16(2–3):229–261, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014454317209. Mod-

eling and computation in optics and electromagnetics.

Bramble:1996:SPA

[BPK96] James H. Bramble, Joseph E.

Pasciak, and Andrew V. Knyazev. A subspace preconditioning algorithm for eigenvector/eigenvalue computation. Advances in Computational Mathematics, 6 (1):159–189, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127702.

Baker:1995:INS

[BPW95] C. T. H. Baker, C. A. H.

Paul, and D. R. Willé. Issues in the numerical solution of evolutionary delay differential equations. Advances in Computational Mathematics, 3 (1–2):171–196, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02988625.

Baroli:2013:CSD

[BQRB13] Davide Baroli, Alfio Quarteroni, and Ricardo Ruiz-Baier. Convergence of a stabilized discontinuous Galerkin method for incompressible nonlinear elasticity. Advances in Computational Mathematics, 39(2):425–443, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-012-9286-8.

Brezinski:2007:CSK

[BR07] C. Brezinski and M. Raydan. Cauchy–Schwarz and Kantorovich type inequalities for scalar and matrix moment sequences. Advances in Computational Mathematics, 26(1–3):71–80, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7646-8.

Bosner:2013:VDP

[BR13] Tina Bosner and Mladen Rogina. Variable degree polynomial splines are Chebyshev splines. Advances in Computational Mathematics, 38 (2):383–400, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9242-z.

Buscaglia:2015:AMD

[BR15] Gustavo Buscaglia and Vitoriano Ruas. Approximation of the multi-dimensional Stokes system with embedded pressure discontinuities. Advances in Computational Mathematics, 41(3):599–634, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9378-8.

Brandts:2001:CRE

[Bra01] Jan Brandts. The Cauchy– Riemann equations: discretization by finite elements, fast solution of the second variable, and a posteriori error estimation. Advances in Computational Mathematics, 15 (1–4):61–77, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014217225870. A

posteriori error estimation and adaptive computational methods.

Brenner:1995:TLA

[Bre95] Susanne C. Brenner. A twolevel additive Schwarz preconditioner for the stationary Stokes equations. Advances in Computational Mathematics, 4(1– 2):111–126, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02123475.

Bozzini:2013:GWM

[BRS13] Mira Bozzini, Milvia Rossini, and Robert Schaback. Generalized Whittle–Matérn and polyharmonic kernels. Advances in Computational Mathematics, 39(1):129–141, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-012-9277-9.

Bozzini:2015:RKS

[BRSV15] Mira Bozzini, Milvia Rossini, Robert Schaback, and Elena Volontè. Radial kernels via scale derivatives. Advances in Computational Mathematics, 41(2):277–291, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9366-z.

Brezinski:1994:GEP

[BRZ94] C. Brezinski and M. Redivo-Zaglia. A general extrapolation procedure revisited. Advances in Computational Mathematics, 2(4):461–477, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02521609.

Burrage:1997:PIMb

[BS97a] K. Burrage and H. Suhartanto. Parallel iterated method based on multistep Runge– Kutta of Radau type for stiff problems. Advances in Computational Mathematics, 7(1–2):59–77, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018982432701. Par-

allel methods for ODEs.

Burrage:1997:PIMa

[BS97b] K. Burrage and H. Suhartanto.

Parallel iterated methods based on multistep Runge–Kutta methods of Radau type. Advances in Computational Mathematics, 7(1–2):37–57, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018930415863. Par-

allel methods for ODEs.

Burrage:2000:PIM

[BS00] K. Burrage and H. Suhartanto.

Parallel iterated methods based on variable step-size multistep Runge–Kutta methods of Radau type for stiff problems. Advances in Computational Mathematics, 13(3):257–270, July 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018906311204.

Bejancu:2005:MAO

[BS05] Aurelian Bejancu and Malcolm A. Sabin. Maximal approximation order for a boxspline semi-cardinal interpolation scheme on the threedirection mesh. Advances in Computational Mathematics, 22(3):275–298, April 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-2601-7.

Benedetto:2006:TFG

[BS06] John J. Benedetto and Songkiat Sumetkijakan. Tight frames and geometric properties of wavelet sets. Advances in Computational Mathematics, 24(1–4):35–56, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7623-2.

Buhmann:1999:F

[Buh99] Martin Buhmann. Foreword. Advances in Computational Mathematics, 11(2–3):i, November 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018963607144.

Buhmann:2006:SIC

[Buh06] M. D. Buhmann. Semi-infinite cardinal interpolation with multiquadrics and beyond. Advances in Computational Mathematics, 24(1–4):57–80, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7612-5.

Bunger:2011:MME

[Bün11] F. Bünger. Minimizing and maximizing the Euclidean norm of the product of two polynomials. Advances in Computational Mathematics, 35(2– 4):193–215, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9158-z.

Burrage:1997:PMO

[Bur97] Kevin Burrage. Parallel methods for ODEs. Advances in Computational Mathematics, 7(1–2):1–3, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018997130884.

Butcher:1997:OSP

[But97] J. C. Butcher. Order and stability of parallel methods for stiff problems. Advances in Computational Mathematics, 7(1–2):79–96, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018934516771. Par-

allel methods for ODEs.

Butler:2015:LAL

[BW15] Ronald W. Butler and Andrew T. A. Wood. Laplace approximation of Lauricella functions F_{A} and F_{D}. Advances in Computational Mathematics, 41(6):1451–1473, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9397-5.

Bogacki:1997:DBO

[BWX97] Przemysław Bogacki, Stanley Weinstein, and Yuesheng Xu. Distances between oriented curves in geometric modeling. Advances in Computational Mathematics, 7(4):593–621, August 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018923609019.

Bajaj:1994:NAS

[BX94] Chandrajit L. Bajaj and Guo Liang Xu. NURBS approximation of surface/surface intersection curves. Advances in Computational Mathematics, 2(1):1–21, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02519033.

Ben-yu:2001:JIA

[ByLl01] Guo Ben-yu and Wang Lilian. Jacobi interpolation approximations and their applications to singular differential equations. Advances in Computational Mathematics, 14(3):227–276, April 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1016681018268.

Cai:2002:AIE

[Cai02]  Wei Cai.  Algorithmic is-

sues for electromagnetic scattering in layered media: Green’s functions, current basis, and fast solver. Advances in Computational Mathematics, 16(2–3):157–174, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014429611757. Mod-

eling and computation in optics and electromagnetics.

Calvi:2005:IUA

[Cal05] Jean-Paul Calvi. Intertwining unisolvent arrays for multivariate Lagrange interpolation. Advances in Computational Mathematics, 23(4):393–414, November 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1840-6.

Chui:2006:COM

[CaL06] Charles K. Chui and Jian ao Lian. Construction of orthonormal multi-wavelets with additional vanishing moments. Advances in Computational Mathematics, 24(1–4):239–262, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
004-7610-7.

Carnicer:1995:MCP

[Car95] J. M. Carnicer. Multivariate convexity preserving interpolation by smooth functions. Advances in Computational Mathematics, 3(4):395–404, May 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02432005.

Casas:2007:UPL

[Cas07] Eduardo Casas. Using piecewise linear functions in the numerical approximation of semilinear elliptic control problems. Advances in Computational Mathematics, 26(1–3):137–153, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-4142-0.

Carstensen:2001:ESP

[CBK01] Carsten Carstensen, Sören Bartels, and Roland Klose. An experimental survey of a posteriori Courant finite element error control for the Poisson equation. Advances in Computational Mathematics, 15 (1–4):79–106, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1014269318616. A

posteriori error estimation and adaptive computational methods.

Cheng:1994:CPB

[CC94] Zheng Xing Cheng and Charles K. Chui. Convexity of parametric Bézier surfaces in terms of Gaussian curvature signatures. Advances in Computational Mathematics, 2(4):437–459, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02521608.

Casazza:2003:GFI

[CC03] Peter G. Casazza and Ole Christensen. Gabor frames over irregular lattices. Advances in Computational Mathematics, 18(2–4):329–344, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021356503075. Frames.

Cui:2012:REE

[CCE^{+}12] Ming Cui, Zhangxin Chen, Richard E. Ewing, Guan Qin, and Hongsen Chen. Reliable and efficient error control for an adaptive Galerkin-characteristic method for convection-dominated diffusion problems. Advances in Computational Mathematics, 37(3):319–353, October 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9210-7.

Cao:2010:EAF

[CCG10] Yanzhao Cao, Zheng Chen, and Max Gunzburger. Error analysis of finite element approximations of the stochastic Stokes equations. Advances in Computational Mathematics, 33(2):215–230, August 2010. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9127-6.

Choi:2008:TDO

[CCH^{+}08] Hyeong In Choi, Sung Woo Choi, Chang Yong Han, Tae wan Kim, Song-Hwa Kwon, Hwan Pyo Moon, Kyeong Hah Roh, and Nam-Sook Wee. Twodimensional offsets and medial axis transform. Advances in Computational Mathematics, 28(2):171–199, February 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9036-5.

Calderbank:2011:SFF

[CCH^{+}11] Robert Calderbank, Peter G. Casazza, Andreas Heinecke, Gitta Kutyniok, and Ali Pezeshki. Sparse fusion frames: existence and construction. Advances in Computational Mathematics, 35(1):1–31, July 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-010-9162-3.

Collowald:2015:NRC

[CCH^{+}15] Mathieu Collowald, Annie Cuyt, Evelyne Hubert, Wen-Shin Lee, and Oliver Salazar Celis. Numerical reconstruction of convex polytopes from directional moments. Advances in Computational Mathematics, 41(6):

1515–1535, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9401-0.

Chan:2016:SDF

[CCN^{+}16] Alice Z.-Y. Chan, Martin S.

Copenhaver, Sivaram K. Narayan, Logan Stokols, and Allison Theobold. On structural decompositions of finite frames. Advances in Computational Mathematics, 42(3):721–756, June 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9440-1.

Conti:2016:FHS

[CCS16] Costanza Conti, Mariantonia Cotronei, and Tomas Sauer. Factorization of Hermite subdivision operators preserving exponentials and polynomials. Advances in Computational Mathematics, 42(5):1055–1079, October 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9453-4.

Cai:2009:CAT

[CCSS09] Jian-Feng Cai, Raymond H.

Chan, Lixin Shen, and Zuowei Shen. Convergence analysis of tight framelet approach for missing data recovery. Advances in Computational Mathematics, 31(1–3):87–113, October 2009. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9084-5.

Charina:2015:SSN

[CD15] Maria Charina and Geir Dahl. Subdivision schemes, network flows and linear optimization. Advances in Computational Mathematics, 41(3):507–528, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9371-2.

Cai:2003:DAH

[CDP03] Zhiqiang Cai, Jim Douglas, Jr., and Moongyu Park. Development and analysis of higher order finite volume methods over rectangles for elliptic equations. Advances in Computational Mathematics, 19(1–3):3–33, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1022841012296. Chal-

lenges in computational mathematics (Pohang, 2001).

Caponnetto:2009:ECC

[CDP09] Andrea Caponnetto, Ernesto De Vito, and Massimiliano Pontil. Entropy conditions for L_{r}-convergence of empirical processes. Advances in Computational Mathematics, 30(4):355–373, May 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9072-9.

Celorrio:2002:PDD

[CDS02] Ricardo Celorrio, Víctor Domínguez, and Francisco-Javier Sayas. Periodic Dirac delta distributions in the boundary element method. Advances in Computational Mathematics, 17(3):211–236, October 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016002302671.

Cuyt:1999:EMP

[CDTV99] Annie Cuyt, Kathy Driver, Jieqing Tan, and Brigitte Verdonk. Exploring multivariate Padé approximants for multiple hypergeometric series. Advances in Computational Mathematics, 10(1):29–49, January 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018918429917.

Chertock:2018:HOP

[CEHK18] Alina Chertock, Yekaterina Epshteyn, Hengrui Hu, and Alexander Kurganov. Highorder positivity-preserving hybrid finite-volume-finite-difference methods for chemotaxis systems. Advances in Computational Mathematics, 44 (1):327–350, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9545-9.

Casenave:2015:NRB

[CEL15] Fabien Casenave, Alexandre Ern, and Tony Lelièvre. A nonintrusive reduced basis method applied to aeroacoustic simulations. Advances in Computational Mathematics, 41 (5):961–986, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9365-0.

Cerna:2011:COC

[ČF11] Dana Černá and Václav Finěk. Construction of optimally conditioned cubic spline wavelets on the interval. Advances in Computational Mathematics, 34(2):219–252, February 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-010-9152-5.

Carnicer:2004:CBC

[CG04] J. M. Carnicer and M. Gasca. Classification of bivariate configurations with simple Lagrange interpolation formulae. Advances in Computational Mathematics, 20(1–3):5–16, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/

10.1023/A:1025823727706. Ap-

proximation and applications.

Carnicer:2006:ILG

[CG06] J. M. Carnicer and M. Gasca.

Interpolation on lattices generated by cubic pencils. Advances in Computational Mathematics, 24(1–4):113–130, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7638-8.

Christensen:2010:POD

[CG10] Ole Christensen and Say Song Goh. Pairs of oblique duals in spaces of periodic functions. Advances in Computational Mathematics, 32(3):353–379, April 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9115-x.

Carnicer:2015:NFG

[CG15] J. M. Carnicer and C. Godés.

A Newton formula for generalized Berzolari–Radon sets. Advances in Computational Mathematics, 41(2):373–386, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9362-3.

Cui:2017:FEA

[CG17] Jintao Cui and Thirupathi Gudi. Finite element approximation of a free boundary plasma problem. Advances in Computational Mathematics, 43(3):517–535, June 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9495-7.

Corless:1996:LF

[CGH^{+}96] R. M. Corless, G. H. Gonnet, D. E. G. Hare, D. J. Jeffrey, and D. E. Knuth. On the Lambert W function. Advances in Computational Mathematics, 5(1):329–359, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124750;   https:
//nickhigham.wordpress.com/
2016/08/23/corless-knuth-
and-lambert-w/.

Chen:2013:NAF

[CGH^{+}13] Huajie Chen, Xingao Gong, Lianhua He, Zhang Yang, and Aihui Zhou. Numerical analysis of finite dimensional approximations of Kohn–Sham models. Advances in Computational Mathematics, 38 (2):225–256, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9235-y.

Costantini:2001:CSP

[CGM01] P. Costantini, T. N. T. Goodman, and C. Manni. Constructing C^{3} shape preserving interpolating space curves. Advances in Computational Mathematics, 14(2):103–127, February 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016664630563.

Carnicer:1995:GVD

[CGP95] J. M. Carnicer, T. N. T. Goodman, and J. M. Peña. A generalization of the variation diminishing property. Advances in Computational Mathematics, 3(4):375–394, May 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02432004.

Conti:2011:AIN

[CGR11] Costanza Conti, Luca Gemignani, and Lucia Romani. From approximating to interpolatory non-stationary subdivision schemes with the same generation properties. Advances in Computational Mathematics, 35(2–4):217–241, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9175-6.

Conti:2013:CAS

[CGR13] Costanza Conti, Luca Gemignani, and Lucia Romani. A constructive algebraic strategy for interpolatory subdivision schemes induced by bivariate box splines. Advances in Computational Mathematics, 39(2):395–424, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-012-9285-9.

Cotronei:2015:ADM

[CGRS15] Mariantonia Cotronei, Daniele Ghisi, Milvia Rossini, and Tomas Sauer. An anisotropic directional multiresolution and subdivision scheme. Advances in Computational Mathematics, 41(3):709–726, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9384-x.

Cliffe:1993:EDN

[CGS93] K. A. Cliffe, T. J. Garratt, and A. Spence. Eigenvalues of the discretized Navier–Stokes equation with application to the detection of Hopf bifurcations. Advances in Computational Mathematics, 1(3):337–356, October 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02072015.

Cochran:2013:EDT

[CGW13] Doug Cochran, Anne Gelb, and Yang Wang. Edge detection from truncated Fourier data using spectral mollifiers. Advances in Computational Mathematics, 38(4):737–762, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9258-4.

Chaulet:2015:EIS

[CH15] N. Chaulet and H. Haddar.

Electromagnetic inverse shape problem for coated obstacles. Advances in Computational Mathematics, 41(6): 1615–1641, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9406-3.

Chen:2002:FDA

[Che02] Yu Chen. A fast, direct algorithm for the Lippmann– Schwinger integral equation in two dimensions. Advances in Computational Mathematics, 16(2–3):175–190, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014450116300. Mod-

eling and computation in optics and electromagnetics.

Chen:2006:UCA

[Che06] Yanping Chen. Uniform convergence analysis of finite difference approximations for singular perturbation problems on an adapted grid. Advances in Computational Mathematics, 24(1–4):197–212, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7641-0.

Chionh:2003:PDM

[Chi03]  Eng-Wee Chionh.  Parallel

Dixon matrices by bracket. Advances in Computational Mathematics, 19(4):373–383, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1024264830187.

Christensen:2017:ECP

[CHL17] Ole Christensen, Marzieh Hasannasab, and Jakob Lemvig. Explicit constructions and properties of generalized shift-invariant systems in L^{2}(R). Advances in Computational Mathematics, 43(2):443–472, April 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9492-x.

Cai:2018:SMD

[CHM18] Mingchao Cai, Peiqi Huang, and Mo Mu. Some multilevel decoupled algorithms for a mixed Navier–Stokes/ Darcy model. Advances in Computational Mathematics, 44 (1):115–145, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9537-9.

Calvo:1997:SST

[CHMR97] M. Calvo, D. J. Higham, J. I.

Montijano, and L. Rández. Stepsize selection for tolerance proportionality in explicit Runge–Kutta codes. Advances in Computational Mathematics, 7(3):361–382, May 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018959222223.

Chong:1995:PTC

[Cho95] C. T. Chong. The polynomial topological complexity of Fatou–Julia sets. Advances in Computational Mathematics, 3(4):369–374, May 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02432003.

Chen:2000:CMB

[CHR00] Di-Rong Chen, Bin Han, and Sherman D. Riemenschneider. Construction of multivariate biorthogonal wavelets with arbitrary vanishing moments. Advances in Computational Mathematics, 13(2):131–165, May 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018950126225.

Christiansen:2015:PDG

[Chr15] Lasse Hjuler Christiansen. Pairs of dual Gabor frames generated by functions of Hilbert– Schmidt type. Advances in Computational Mathematics, 41 (6):1537–1554, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9402-7.

Chen:2006:BSA

[CHRX06] Qiuhui Chen, Norden Huang, Sherman Riemenschneider, and Yuesheng Xu. A B-spline approach for empirical mode decompositions. Advances in Computational Mathematics, 24 (1–4):171–195, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7614-3.

Claeys:2017:SKB

[CHS17] X. Claeys, R. Hiptmair, and E. Spindler. Second kind boundary integral equation for multi-subdomain diffusion problems. Advances in Computational Mathematics, 43 (5):1075–1101, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9517-0.

Chui:2003:CST

[CHSS03] Charles K. Chui, Wenjie He, Joachim Stöckler, and Qiyu Sun. Compactly supported tight affine frames with integer dilations and maximum vanishing moments. Advances in Computational Mathematics, 18(2–4):159–187, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021318804341. Frames.

Cheng:2010:PSI

[CHY10] Jin Cheng, Benny Yiu Chung Hon, and Masahiro Yamamoto. Preface to the special issue on inverse problems. Advances in Computational Mathematics, 33(4):381–383, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9166-z;
http://link.springer.com/
content/pdf/10.1007/s10444-
010-9166-z.pdf.

Cui:2011:GMP

[CHY11] Xiaoke Cui, Ken Hayami, and Jun-Feng Yin. Greville’s method for preconditioning least squares problems. Advances in Computational Mathematics, 35(2–4):243–269, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9171-x.

Cantero:2013:EOP

[CI13] María José Cantero and Arieh Iserles. On expansions in orthogonal polynomials. Advances in Computational Mathematics, 38(1):35–61, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9225-0.

Chalmoviansky:2007:NLC

[CJ07] Pavel Chalmovianský and Bert Jüttler. A non-linear circlepreserving subdivision scheme. Advances in Computational Mathematics, 27(4):375–400, November 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
005-9011-y.

Cao:2017:FDS

[CJX17] Yanzhao Cao, Ying Jiang, and Yuesheng Xu. A fast discrete spectral method for stochastic partial differential equations. Advances in Computational Mathematics, 43 (5):973–998, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9513-4.

Casazza:2003:ENT

[CK03a] Peter G. Casazza and Jelena Kovačević. Equal-norm tight frames with erasures. Advances in Computational Mathematics, 18(2–4):387–430, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021349819855. Frames.

Chae:2003:PSU

[CK03b] Gee Ju Chae and Hong Oh Kim.

On the preservation of stability under convolutions. Advances in Computational Mathematics, 19(1–3):183–193, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1022811229621. Chal-

lenges in computational mathematics (Pohang, 2001).

Casazza:2007:GGS

[CK07] Peter G. Casazza and Gitta Kutyniok. A generalization of Gram–Schmidt orthogonalization generating all Parseval frames. Advances in Computational Mathematics, 27(1):65–78, July 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7478-1.

Chadha:2011:MNP

[CK11] Naresh M. Chadha and Natalia Kopteva. Maximum norm a posteriori error estimate for a 3 d singularly perturbed semilinear reaction–diffusion problem. Advances in Computational Mathematics, 35(1):33–55, July 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9163-2.

Cvetkovic:2011:SGG

[CKBP11] Ljiljana Cvetković, Vladimir Kostić, Rafael Bru, and Francisco Pedroche. A simple generalization of Geršgorin’s theorem. Advances in Computational Mathematics, 35(2– 4):271–280, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9143-6.

Christensen:2012:GWS

[CKK12] Ole Christensen, Hong Oh Kim, and Rae Young Kim. Gabor windows supported on [−1, 1] and dual windows with small support. Advances in Computational Mathematics, 36(4):525–545, May 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9189-0.

Christensen:2016:EWS

[CKK16] Ole Christensen, Hong Oh Kim, and Rae Young Kim. On extensions of wavelet systems to dual pairs of frames. Advances in Computational Mathematics, 42(2):489–503, April 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-015-9432-1.

Costantini:1999:DHT

[CKM99] Paolo Costantini, Boris I.

Kvasov, and Carla Manni. On discrete hyperbolic tension splines. Advances in Computational Mathematics, 11 (4):331–354, December 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018988312596.

Conceicao:2013:CSC

[CKP13] Ana C. Conceição, Viktor G.

Kravchenko, and José C. Pereira. Computing some classes of Cauchy type singular integrals with Mathematica software. Advances in Computational Mathematics, 39(2):273–288, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-012-9279-7.

Chou:2010:OCA

[CKW10] So-Hsiang Chou, Do Y. Kwak, and K. T. Wee. Optimal convergence analysis of an immersed interface finite element method. Advances in Computational Mathematics, 33(2):149–168, August 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9122-y.

Cotronei:2002:MCF

[CL02] Mariantonia Cotronei and Maria Laura Lo Cascio. A method for the construction of families of biorthogonal filters with prescribed properties. Advances in Computational Mathematics, 17(3):199–210, October 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1016040516242.

Chen:2007:AEE

[CL07] Wei Chen and Qun Lin. Asymptotic expansion and extrapolation for the eigenvalue approximation of the biharmonic eigenvalue problem by Ciarlet– Raviart scheme. Advances in Computational Mathematics, 27(1):95–106, July 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9031-x.

Cahill:2012:DIF

[CL12] Jameson Cahill and Shidong Li. Dimension invariance of finite frames of translates and Gabor frames. Advances in Computational Mathematics, 37 (4):505–520, November 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9222-3.

Chernih:2016:MMC

[CL16] A. Chernih and Q. T. Le Gia. Multiscale methods with compactly supported radial basis functions for the Stokes problem on bounded domains. Advances in Computational Mathematics, 42(5):1187–1208, October 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9458-z.

Choi:2018:LFD

[CL18] Gary Pui-Tung Choi and Lok Ming Lui. A linear formulation for disk conformal parameterization of simplyconnected open surfaces. Advances in Computational Mathematics, 44(1):87–114, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9536-x.

Cai:2016:UCO

[CLC16] Wentao Cai, Jian Li, and Zhangxin Chen. Unconditional convergence and optimal error estimates of the Euler semiimplicit scheme for a generalized nonlinear Schrödinger equation. Advances in Computational Mathematics, 42(6): 1311–1330, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9463-2.

Chen:2008:FFD

[CLL08] Guo Chen, Zhilin Li, and Ping Lin. A fast finite difference method for biharmonic equations on irregular domains and its application to an incompressible Stokes flow. Advances in Computational Mathematics, 29(2):113–133, August 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9043-6.

Camp:2006:QIF

[CLLS06] Brian Camp, Tao Lin, Yanping Lin, and Weiwei Sun. Quadratic immersed finite element spaces and their approximation capabilities. Advances in Computational Mathematics, 24(1–4):81–112, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-4139-8.

Chatwin:1996:ADS

[CLM96a] P. C. Chatwin, D. M. Lewis, and N. Mole. Atmospheric diffusion: some new mathematical models. Advances in Computational Mathematics, 6 (1):227–242, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127705.  John

Crank 80th birthday special issue.

Chui:1996:LAC

[CLM96b] C. K. Chui, Xin Li, and H. N. Mhaskar. Limitations of the approximation capabilities of neural networks with one hidden layer. Advances in Computational Mathematics, 5(1):233–243, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124745.

Choi:2002:CAS

[CLM02] Hyeong In Choi, Doo Seok Lee, and Hwan Pyo Moon. Clifford algebra, spin representation, and rational parameterization of curves and surfaces. Advances in Computational Mathematics, 17(1–2):5–48, July 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1015294029079. Ad-

vances in geometrical algorithms and representations.

Calvo:2003:TSH

[CLM03] M. Calvo, M. P. Laburta, and J. I. Montijano. Two-step high order starting values for implicit Runge–Kutta methods. Advances in Computational Mathematics, 19(4):401–412, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1024281802353.

Calvo:2010:APQ

[CLMR10] M. Calvo, M. P. Laburta, J. I.

Montijano, and L. Rández.

Approximate preservation of quadratic first integrals by explicit Runge–Kutta methods. Advances in Computational Mathematics, 32(3):255–274, April 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9105-4.

Calvo:2015:RKP

[CLMR15] M. Calvo, M. P. Laburta, J. I. Montijano, and L. Rández. Runge–Kutta projection methods with low dispersion and dissipation errors. Advances in Computational Mathematics, 41(1):231–251, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9355-2.

Chen:2002:NOE

[CLZ02] Zhongying Chen, Ronghua Li, and Aihui Zhou. A note on the optimal L^{2}-estimate of the finite volume element method. Advances in Computational Mathematics, 16(4):291–303, May 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014577215948.

Chen:2013:UEE

[CLZW13] Hongtao Chen, Qun Lin, Junming Zhou, and Hong Wang. Uniform error estimates for triangular finite element solutions of advection– diffusion equations. Advances in Computational Mathematics, 38(1):83–100, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9228-x.

Chen:1996:HMM

[CM96] Chunhui Chen and O. L. Mangasarian. Hybrid misclassification minimization. Advances in Computational Mathematics, 5(1):127–136, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124738.

Carrig:1999:POC

[CM99] James J. Carrig, Jr. and Gerard G. L. Meyer. A parameterized ordering for cache-, register- and pipeline-efficient Givens QR decomposition. Advances in Computational Mathematics, 10(1):97–113, January 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018970413988.

Costantini:2003:SPI

[CM03] P. Costantini and C. Manni.

Shape-preserving C^{3} interpolation: the curve case. Advances in Computational Mathematics, 18(1):41–63, January 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1021270530342.

Christensen:2012:EBS

[CM12] Ole Christensen and Peter Massopust. Exponential Bsplines and the partition of unity property. Advances in Computational Mathematics, 37(3):301–318, October 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9208-1.

Carnicer:2003:RS

[CMP03] J. M. Carnicer, E. Mainar, and J. M. Peña. Restricted systems. Advances in Computational Mathematics, 18(1):79–90, January 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1021267126479.

Cachafeiro:2007:OPR

[CMP07a] A. Cachafeiro, F. Marcellán, and C. Pérez. Orthogonal polynomials with respect to the sum of an arbitrary measure and a Bernstein–Szegő measure. Advances in Computational Mathematics, 26(1–3):81–104, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7644-x.

Carnicer:2007:SPR

[CMP07b] J. M. Carnicer, E. Mainar, and J. M. Peña. Shape preservation regions for sixdimensional spaces. Advances in Computational Mathematics, 26(1–3):121–136, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7505-2.

Calvo:2007:EMR

[CMR07] M. Calvo, J. I. Montijano, and L. Rández. On explicit multi-revolution Runge– Kutta schemes. Advances in Computational Mathematics, 26 (1–3):105–120, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7209-z.

Chen:1997:PGM

[CMX97] Zhongying Chen, Charles A.

Micchelli, and Yuesheng Xu. The Petrov–Galerkin method for second kind integral equations. II. Multiwavelet schemes. Advances in Computational Mathematics, 7(3):199–233, May 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018994802659.

Chen:2002:DWP

[CMX02] Zhongying Chen, Charles A.

Micchelli, and Yuesheng Xu.

Discrete wavelet Petrov–Galerkin methods. Advances in Computational Mathematics, 16(1): 1–28, January 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1023/A:1014273420351.

Chen:2007:MCP

[CMX07] Qiuhui Chen, Charles A. Micchelli, and Yuesheng Xu. On the matrix completion problem for multivariate filter bank construction. Advances in Computational Mathematics, 26 (1–3):173–204, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7535-9.

Crank:1996:PMN

[CN96] J. Crank and P. Nicolson. A practical method for numerical evaluation of solutions of partial differential equations of the heat-conduction type [reprint of MR0019410 (8,409b)]. Advances in Computational Mathematics, 6(1):207–226, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127704.  John

Crank 80th birthday special issue.

Cope:1998:RGG

[Cop98]  Davis K. Cope.  Recursive

generation of the Galerkin– Chebyshev matrix for convolution kernels. Advances in Computational Mathematics, 9 (1–2):21–35, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018933305810. Nu-

merical treatment of boundary integral equations.

Carnicer:1993:SPR

[CP93] J. M. Carnicer and J. M.

Peña. Shape preserving representations and optimality of the Bernstein basis. Advances in Computational Mathematics, 1(2):173–196, June 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02071384.

Chen:1999:QWA

[CP99] Han-Lin Chen and Si-Long Peng. A quasi-wavelet algorithm for second kind boundary integral equations. Advances in Computational Mathematics, 11(4):355–375, December 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018992413504.

Chen:2003:LDP

[CP03] Han-Lin Chen and Si-Long Peng. Localization of dual periodic scaling and wavelet functions. Advances in Computational Mathematics, 19(1–3):

195–210, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1023/A:1022882423107.

Challenges in computational mathematics (Pohang, 2001).

Costantini:2004:SPA

[CP04] Paolo Costantini and Francesca Pelosi. Shape-preserving approximation of spatial data. Advances in Computational Mathematics, 20(1–3):25–51, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1025803122254. Ap-

proximation and applications.

Carnicer:2007:WMG

[CP07a] J. M. Carnicer and J. M. Peña. The work of Mariano Gasca. Advances in Computational Mathematics, 26(1–3):1–8, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-006-9024-1.

Costantini:2007:SPH

[CP07b] Paolo Costantini and Francesca Pelosi. Shape preserving histogram approximation. Advances in Computational Mathematics, 26(1–3):205–230, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-8008-2.

Casazza:2014:WFF

[CP14] Peter G. Casazza and Jesse Peterson. Weighted fusion frame construction via spectral tetris. Advances in Computational Mathematics, 40(2):335–351, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9310-7.

Chernov:2015:STP

[CP15] Alexey Chernov and Duong Pham. Sparse tensor product spectral Galerkin BEM for elliptic problems with random input data on a spheroid. Advances in Computational Mathematics, 41(1):77–104, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9350-7.

Calvetti:1999:IEJ

[CR99] D. Calvetti and L. Reichel.

On an inverse eigenproblem for Jacobi matrices. Advances in Computational Mathematics, 11(1):11–20, July 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018951304418.

Christensen:2008:FPW

[CR08] Ole Christensen and Asghar Rahimi. Frame properties of wave packet systems in L^{2}(R^{d}). Advances in Computational Mathematics, 29 (2):101–111, August 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9038-3.

Champneys:1993:HHO

[CS93] Alan R. Champneys and Alastair Spence. Hunting for homoclinic orbits in reversible systems: a shooting technique. Advances in Computational Mathematics, 1(1):81–108, February 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02070822.

Chui:1994:CTS

[CS94] Charles K. Chui and Xian Liang Shi. Continuous two-scale equations and dyadic wavelets. Advances in Computational Mathematics, 2(2):185–213, March 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02521107.

Christensen:2003:FSB

[CS03] Ole Christensen and Diana T. Stoeva. p-Frames in separable Banach spaces. Advances in Computational Mathematics, 18(2–4):117–126, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021364413257. Frames.

Cavaretta:2006:ORI

[CS06] A. S. Cavaretta and A. Sharma.

Optimal recovery of interpolation operators in Hardy spaces. Advances in Computational Mathematics, 24 (1–4):131–141, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-4134-0.

Chui:1998:AFQ

[CSS98] Charles K. Chui, Xianliang Shi, and Joachim Stöckler. Affine frames, quasi-affine frames, and their duals. Advances in Computational Mathematics, 8(1–2):1–17, January 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018975725857.

Chui:1996:AWG

[CSW96] C. K. Chui, J. Stöckler, and J. D. Ward. Analytic wavelets generated by radial functions. Advances in Computational Mathematics, 5 (1):95–123, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124736.

Chernih:2014:WFI

[CSW14] A. Chernih, I. H. Sloan, and R. S. Womersley. Wendland functions with increasing smoothness converge to a Gaussian. Advances in Computational Mathematics, 40 (1):185–200, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9304-5.

Cong:1999:RKN

[CSWP99] Nguyen Huu Cong, Karl Strehmel, Rüdiger Weiner, and Helmut Podhaisky. Runge– Kutta–Nyström-type parallel block predictor–corrector methods. Advances in Computational Mathematics, 10 (2):115–133, February 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018930732643.

Capobianco:2005:IPW

[CT05] M. R. Capobianco and W. Themistoclakis. Interpolating polynomial wavelets on [−1, 1]. Advances in Computational Mathematics, 23(4):353–374, November 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1828-2.

Choi:2017:GBE

[CT17a] Young Ok Choi and Johannes Tausch. The Galerkin boundary element method for transient Stokes flow. Advances in Computational Mathematics, 43(3):473–493, June 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9493-9.

Cuong:2017:WBV

[CT17b] Dao Huy Cuong and Mai Duc Thanh. A well-balanced van Leer-type numerical scheme for shallow water equations with variable topography. Advances in Computational Mathematics, 43(5):1197–1225, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9521-4.

Cheung:2002:SLL

[CTZ02] Hoi Ling Cheung, Canqin Tang, and Ding-Xuan Zhou. Supports of locally linearly independent M-refinable functions, attractors of iterated function systems and tilings. Advances in Computational Mathematics, 17(3):257–268, October 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1011906532250.

Cullum:1996:IMS

[Cul96] Jane Cullum. Iterative methods for solving Ax = b, GMRES/ FOM versus QMR/BiCG. Advances in Computational Mathematics, 6(1):1–24, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02127693.

Cuminato:1996:NSC

[Cum96] Jose A. Cuminato. Numerical solution of Cauchy-type integral equations of index −1 by collocation methods. Advances in Computational Mathematics, 6(1):47–64, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02127695.

Chui:2009:PMA

[CW09] Charles K. Chui and Jianzhong Wang. PDE models associated with the bilateral filter. Advances in Computational Mathematics, 31(1–3):131–156, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9095-2.

Carl:2016:VLS

[CW16] Wolfgang Carl and Johannes Wallner. Variational Laplacians for semidiscrete surfaces. Advances in Computational Mathematics, 42(6): 1491–1509, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9472-1;
http://link.springer.com/
content/pdf/10.1007/s10444-
016-9472-1.pdf.

Chandler-Wilde:1995:EBI

[CWLH95] S. N. Chandler-Wilde, J. Levesley, and D. M. Hough. Evaluation of a boundary integral representation for the conformal mapping of the unit disk onto a simply-connected domain. Advances in Computational Mathematics, 3(1–2):115–135, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF03028363.

Chen:2006:MAM

[CWX06] Zhongying Chen, Bin Wu, and Yuesheng Xu. Multilevel augmentation methods for differential equations. Advances in Computational Mathematics, 24(1–4):213–238, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-4092-6.

Chen:2012:HOF

[CWX12] Zhongying Chen, Junfeng Wu, and Yuesheng Xu. Higherorder finite volume methods for elliptic boundary value problems. Advances in Computational Mathematics, 37 (2):191–253, August 2012.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9201-8.

Chen:2004:CAN

[CX04] Qiuhui Chen and Yuesheng Xu. The cascade algorithm for the numerical computation of refinable functions. Advances in Computational Mathematics, 20(1–3):17–24, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1025844410775. Ap-

proximation and applications.

Chen:2006:CMS

[CX06] Di-Rong Chen and Dao-Hong Xiang. The consistency of multicategory support vector machines. Advances in Computational Mathematics, 24 (1–4):155–169, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7207-1.

Chen:2016:FPA

[CZS16] Dai-Qiang Chen, Yan Zhou, and Li-Juan Song. Fixed point algorithm based on adapted metric method for convex minimization problem with application to image deblurring. Advances in Computational Mathematics, 42(6):1287–1310, December 2016. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9462-3.

Cao:2017:SIF

[CZZ17] Waixiang Cao, Xu Zhang, and Zhimin Zhang. Superconvergence of immersed finite element methods for interface problems. Advances in Computational Mathematics, 43(4):795–821, August 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9507-7.

Dadi:2017:DTC

[Dad17] Z. Dadi. Dynamics of two-cell systems with discrete delays. Advances in Computational Mathematics, 43(3):653–676, June 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9501-0.

Dahmen:1995:P

[Dah95] Wolfgang Dahmen. Preface. Advances in Computational Mathematics, 4(1–2):iii–iv, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02123470.

Dai:2016:NDD

[Dai16] Ping-Fan Dai. A note on diagonal dominance, Schur complements and some classes of Hmatrices and P -matrices. Advances in Computational Mathematics, 42(1):1–4, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9375-y.

Damelin:2007:PBO

[Dam07]  S. B. Damelin.  Pointwise

bounds of orthogonal expansions on the real line via weighted Hilbert transforms. Advances in Computational Mathematics, 26(4):451–471, May 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7472-7.

Diop:2013:PBI

[DAP13] El Hadji S. Diop, Radjesvarane Alexandre, and Valérie Perrier. A PDE based and interpolation-free framework for modeling the sifting process in a continuous domain. Advances in Computational Mathematics, 38(4):801–835, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9260-x.

Daras:2003:GPT

[Dar03] Nicholas J. Daras. Generalized Padé-type approximation and integral representations. Advances in Computational Mathematics, 18(1):1–39, January 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021273702443.

deBoor:2000:CAM

[dB00] Carl de Boor. Computational aspects of multivariate polynomial interpolation: indexing the coefficients. Advances in Computational Mathematics, 12(4):289–301, March 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018929521682. Mul-

tivariate polynomial interpolation.

deBoor:2006:IRP

[dB06] Carl de Boor. On interpolation by radial polynomials. Advances in Computational Mathematics, 24(1–4):143–153, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7630-3.

deBoor:2007:ISS

[dB07]  C. de Boor.  Interpolation

from spaces spanned by monomials. Advances in Computational Mathematics, 26 (1–3):63–70, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7538-6.

daCunha:2007:DBG

[dCB07] R. D. da Cunha and D. Becker. Dynamic block GMRES: an iterative method for block linear systems. Advances in Computational Mathematics, 27 (4):423–448, November 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-006-9012-5.

Davies:1994:TMN

[DD94] Penny J. Davies and Dugald B. Duncan. Time-marching numerical schemes for the electric field integral equation on a straight thin wire. Advances in Computational Mathematics, 2(3):279–317, April 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02521113.

DAmbrosio:2014:NIH

[DDP14] Raffaele D’Ambrosio, Giuseppe De Martino, and Beatrice Paternoster. Numerical integration of Hamiltonian problems by G-symplectic methods. Advances in Computational Mathematics, 40(2):553–575, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9318-z.

DeVilliers:2006:RFS

[De 06] Johan De Villiers. On refinable functions and subdivision with positive masks. Advances in Computational Mathematics, 24(1–4):281–295, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7628-x.

DellAcqua:2017:NTB

[Del17] Pietro Dell’Acqua. A note on Taylor boundary conditions for accurate image restoration. Advances in Computational Mathematics, 43(6): 1283–1304, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9525-0.

Derriennic:2004:SPP

[Der04] Marie Madeleine Derriennic.

Shape preserving polynomial curves. Advances in Computational Mathematics, 20(4): 273–292, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1023/A:1027303701445.

deFrutos:2018:AGD

[dFGAJN18] Javier de Frutos, Bosco García-Archilla, Volker John, and Julia Novo. Analysis of the grad– div stabilization for the timedependent Navier–Stokes equations with inf–sup stable finite elements. Advances in Computational Mathematics, 44 (1):195–225, February 2018.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9540-1.

Dahlke:2007:AFM

[DFR07] Stephan Dahlke, Massimo Fornasier, and Thorsten Raasch. Adaptive frame methods for elliptic operator equations. Advances in Computational Mathematics, 27(1):27–63, July 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-005-7501-6.

Dominguez:2013:ICA

[DG13] V. Domínguez and M. Ganesh. Interpolation and cubature approximations and analysis for a class of wideband integrals on the sphere. Advances in Computational Mathematics, 39 (3–4):547–584, October 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-012-9293-9.

Dominguez:2016:SEC

[DG16] V. Domínguez and M. Ganesh. Sobolev estimates for constructive uniform-grid FFT interpolatory approximations of spherical functions. Advances in Computational Mathematics, 42(4):843–887, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9445-9.

Dong:2015:ATV

[DGK15] Yiqiu Dong, Torsten Görner, and Stefan Kunis. An algorithm for total variation regularized photoacoustic imaging. Advances in Computational Mathematics, 41(2):423–438, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9364-1.

Dehmer:2013:EHA

[DGMES13] Matthias Dehmer, Martin Grabner, Abbe Mowshowitz, and Frank Emmert-Streib. An efficient heuristic approach to detecting graph isomorphism based on combinations of highly discriminating invariants. Advances in Computational Mathematics, 39(2):311–325, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-012-9281-0.

Dominguez:2016:PMF

[DGMM16] Sebastián Domínguez, Gabriel N.

Gatica, Antonio Márquez, and Salim Meddahi. A primal-mixed formulation for the strong coupling of quasi-Newtonian fluids with porous media. Advances in Computational Mathematics, 42(3):675–720, June 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-015-9439-7.

Dunster:2018:UAE

[DGS18] T. M. Dunster, A. Gil, and J. Segura. Uniform asymptotic expansions for Laguerre polynomials and related confluent hypergeometric functions. Advances in Computational Mathematics, ??(??):1–34, January 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

daCunha:1994:PIR

[dH94] Rudnei Dias da Cunha and Tim Hopkins. A parallel implementation of the restarted GMRES iterative algorithm for nonsymmetric systems of linear equations. Advances in Computational Mathematics, 2(3):261–277, April 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02521112.

Dallas:1998:ONS

[DHK98] Allan G. Dallas, G. C. Hsiao, and R. E. Kleinman. Observations on the numerical stability of the Galerkin method. Advances in Computational Mathematics, 9(1–2):37–67, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018941607627. Nu-

merical treatment of boundary integral equations.

Deano:2016:CIA

[DHO16] Alfredo Deaño, Daan Huybrechs, and Peter Opsomer. Construction and implementation of asymptotic expansions for Jacobi-type orthogonal polynomials. Advances in Computational Mathematics, 42(4):791–822, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9442-z.

Drouiche:1999:DMW

[DKB99] Karim Drouiche, Djalil Kateb, and Daniel Boichu. Daubechies– Matzinger wavelets and Lorentz polynomials. Advances in Computational Mathematics, 10(3–4):239–260, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018934800784.

DeVore:1993:WCN

[DKLT93] R. A. DeVore, G. Kyriazis, D. Leviatan, and V. M. Tikhomirov. Wavelet compression and nonlinear n-widths. Advances in Computational Mathematics, 1(2):197–214, June 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02071385.

Dagnino:1998:NID

[DL98] C. Dagnino and P. Lamberti. Numerical integration of 2-D integrals based on local bivariate C^{1} quasiinterpolating splines. Advances in Computational Mathematics, 8(1–2):19–31, January 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018927809928.

Dekel:2004:MEK

[DL04] S. Dekel and D. Leviatan. On measuring the efficiency of kernel operators in L_{p}(R^{d}). Advances in Computational Mathematics, 20(1–3):53–65, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1025838129042. Ap-

proximation and applications.

deLeon:2007:DVI

[dLdDSM07] Manuel de León, David Martín de Diego, and Aitor Santamaría-Merino. Discrete variational integrators and optimal control theory. Advances in Computational Mathematics, 26(1–3):251–268, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-4093-5.

Dong:2012:FAI

[DLK12] Fangfang Dong, Chunxiao Liu, and De-Xing Kong. A fast algorithm for image segmentation based on fuzzy region competition. Advances in Computational Mathematics, 37 (4):521–542, November 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9221-4.

Diele:1998:CTN

[DLP98] F. Diele, L. Lopez, and R. Peluso. The Cayley transform in the numerical solution of unitary differential systems. Advances in Computational Mathematics, 8(4):317–334, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018908700358.

Dominguez:2014:NMT

[DLS14] Víctor Domínguez, Sijiang L.

Lu, and Francisco-Javier Sayas. A Nyström method for the two dimensional Helmholtz hypersingular equation. Advances in Computational Mathematics, 40(5–6):1121–1157, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9344-5.

Donat:2017:FNO

[DLUS17] Rosa Donat, Sergio López-Ureña, and Maria Santágueda. A family of non-oscillatory 6-point interpolatory subdivision schemes. Advances in Computational Mathematics, 43(4):849–883, August 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9509-5.

Du:2016:PEA

[DLZ16] Shaohong Du, Runchang Lin, and Zhimin Zhang. A posteriori error analysis of multipoint flux mixed finite element methods for interface problems. Advances in Computational Mathematics, 42(4):921–945, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-015-9447-7.

Dahmen:1993:CRE

[DM93] Wolfgang Dahmen and Charles A. Micchelli. Continuous refinement equations and subdivision. Advances in Computational Mathematics, 1(1):1–37, February 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02070819.

Dubuc:1999:DHI

[DM99] Serge Dubuc and Jean-Louis Merrien. Dyadic Hermite interpolation on a rectangular mesh. Advances in Computational Mathematics, 10(3–4):343–365, May 1999. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018943002601.

Dorfler:2015:NGF

[DM15] Monika Dörfler and Ewa Matusiak. Nonstationary Gabor frames — approximately dual frames and reconstruction errors. Advances in Computational Mathematics, 41(2):293–316, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9358-z.

DeSterck:2011:TLA

[DMMS11] H. De Sterck, K. Miller, T. Manteuffel, and G. Sanders. Toplevel acceleration of adaptive algebraic multilevel methods for steady-state solution to Markov chains. Advances in Computational Mathematics, 35(2– 4):375–403, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9168-x.

Dahlke:2003:RWR

[DMT03] Stephan Dahlke, Peter Maass, and Gerd Teschke. Reconstruction of wideband reflectivity densities by wavelet transforms. Advances in Computational Mathematics, 18 (2–4):189–209, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1021303718373. Frames.

Douglas:1994:SRC

[Dou94] Craig C. Douglas. Some remarks on completely vectorizing point Gauss–Seidel while using the natural ordering. Advances in Computational Mathematics, 2(2):215–222, March 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02521108.

Dikshit:1994:WTR

[DOZ94] H. P. Dikshit, A. Ojha, and R. A. Zalik. Wachspress type rational complex planar splines of degree (3, 1). Advances in Computational Mathematics, 2(2):235–249, March 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02521110.

Delgado:2006:GBB

[DP06] Jorge Delgado and Juan Manuel Peña. On the generalized Ball bases. Advances in Computational Mathematics, 24 (1–4):263–280, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7636-x.

Delgado:2013:ERT

[DP13] Jorge Delgado and Juan Manuel Peña. On the evaluation of rational triangular Bézier surfaces and the optimal stability of the basis. Advances in Computational Mathematics, 38(4):701–721, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9256-6.

Dahmen:1993:WAM

[DPS93] W. Dahmen, S. Prössdorf, and R. Schneider. Wavelet approximation methods for pseudodifferential equations. II. Matrix compression and fast solution. Advances in Computational Mathematics, 1(3):259–335, October 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02072014.

Dominguez:2008:DDM

[DRS08] V. Domínguez, M.-L. Rapún, and F.-J. Sayas. Dirac delta methods for Helmholtz transmission problems. Advances in Computational Mathematics, 28(2):119–139, February 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-006-9015-2.

Davydov:2000:LLI

[DS00] Oleg Davydov and Larry L.

Schumaker. Locally linearly independent bases for bivariate polynomial spline spaces. Advances in Computational Mathematics, 13(4):355–373, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016626526861.

Dominguez:2001:FAS

[DS01] Víctor Domínguez and Francisco-Javier Sayas. Full asymptotics of spline Petrov–Galerkin methods for some periodic pseudodifferential equations. Advances in Computational Mathematics, 14(1):75–101, January 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016620816451.

DeMarchi:2010:SKB

[DS10] Stefano De Marchi and Robert Schaback. Stability of kernelbased interpolation. Advances in Computational Mathematics, 32(2):155–161, February 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9093-4;
 http://link.springer.com/
content/pdf/10.1007/s10444-
008-9093-4.pdf.

Dyn:2013:DTP

[DS13] Nira Dyn and Maria Skopina. Decompositions of trigonometric polynomials with applications to multivariate subdivision schemes. Advances in Computational Mathematics, 38 (2):321–349, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9239-7.

Didas:2009:CDG

[DSS09] Stephan Didas, Simon Setzer, and Gabriele Steidl. Combined l_{2} data and gradient fitting in conjunction with l_{1} regularization. Advances in Computational Mathematics, 30(1):79–99, January 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9061-4.

Dahlke:2004:CSB

[DST04] Stephan Dahlke, Gabriele Steidl, and Gerd Teschke. Coorbit spaces and Banach frames on homogeneous spaces with applications to the sphere. Advances in Computational Mathematics, 21(1–2):147–180, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000016435.
42220.fa.

DeMarchi:2005:NOD

[DSW05] Stefano De Marchi, Robert Schaback, and Holger Wendland. Near-optimal data-independent point locations for radial basis function interpolation. Advances in Computational Mathematics, 23(3):317–330, October 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1829-1.

DeVore:1996:SRG

[DT96] R. A. DeVore and V. N. Temlyakov. Some remarks on greedy algorithms. Advances in Computational Mathematics, 5(1):173–187, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124742.

Duan:2012:DDM

[DT12] Yuping Duan and Xue-Cheng Tai. Domain decomposition methods with graph cuts algorithms for total variation minimization. Advances in Computational Mathematics, 36 (2):175–199, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9213-4.

Dung:2009:NLS

[Dũn09] Dinh Dũng. Non-linear sampling recovery based on quasiinterpolant wavelet representations. Advances in Computational Mathematics, 30(4): 375–401, May 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
008-9074-7.

Dung:2011:OAS

[Dũn11] Dinh Dũng. Optimal adaptive sampling recovery. Advances in Computational Mathematics, 34(1):1–41, January 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9140-9.

See erratum [Dũn12].

Dung:2012:EOA

[Dũn12] Dinh Dũng. Erratum to: Optimal adaptive sampling recovery [mr2783300]. Advances in Computational Mathematics, 36(4):605–606, May 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9198-z;
 http://link.springer.com/
content/pdf/10.1007/s10444-
011-9198-z.pdf. See [Dũn11].

Dubickas:2010:RES

[DX10] Artūras Dubickas and Zhiqiang Xu. Refinement equations and spline functions. Advances in Computational Mathematics, 32(1):1–23, January 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9079-2.

Du:2001:GRT

[DY01] Liu Du and Ningning Yan. Gradient recovery type a posteriori error estimate for finite element approximation on nonuniform meshes. Advances in Computational Mathematics, 14(2):175–193, February 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1016676917360.

Didenko:2010:SRM

[DY10] Victor Didenko and Wee Ping Yeo. The spectral radius of matrix continuous refinement operators. Advances in Computational Mathematics, 33(1):113–127, July 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9124-9.

Davydov:2004:SDF

[DZ04] Oleg Davydov and Frank Zeilfelder. Scattered data fitting by direct extension of local polynomials to bivariate splines. Advances in Computational Mathematics, 21(3–4):223–271, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/B:ACOM.0000032041.
                                        68678.fa.

Ephremidze:2013:ADW

[EGL13] L. Ephremidze, A. Gamkrelidze, and E. Lagvilava. An approximation of Daubechies wavelet matrices by perfect reconstruction filter banks with rational coefficients. Advances in Computational Mathematics, 38(1):147–158, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9232-1.

Escribano:2013:HMR

[EGST13] Carmen Escribano, Antonio Giraldo, M. Asunción Sastre, and Emilio Torrano. The Hessenberg matrix and the Riemann mapping function. Advances in Computational Mathematics, 39(3–4):525–545, October 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-012-9291-y.

Ehrich:2000:EWC

[Ehr00] Sven Ehrich. On the estimation of wavelet coefficients. Advances in Computational Mathematics, 13(2):105–129, May 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018998009387.

Elschner:2002:FES

[EHS02] Johannes Elschner, Rainer Hinder, and Gunther Schmidt. Finite element solution of conical diffraction problems. Advances in Computational Mathematics, 16(2–3):139–156, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014456026778. Mod-

eling and computation in optics and electromagnetics.

Ezquerro:2017:SIC

[EHV17] J. A. Ezquerro and M. A. Hernández-Verón. A study of the influence of center conditions on the domain of parameters of Newton’s method by using recurrence relations. Advances in Computational Mathematics, 43(5):1103–1129, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9518-z.

Enenkel:1997:DDI

[EJ97] Robert F. Enenkel and Kenneth R. Jackson. DIMSEMsdiagonally implicit single-eigenvalue methods for the numerical solution of stiff ODEs on parallel computers. Advances in Computational Mathematics, 7(1–2):97–133, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018986500842. Par-

allel methods for ODEs.

Evgeniou:2000:RNS

[EPP00] Theodoros Evgeniou, Massimiliano Pontil, and Tomaso Poggio. Regularization networks and support vector machines. Advances in Computational Mathematics, 13(1):1–50, April 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018946025316.

Engelborghs:1999:NCS

[ER99] Koen Engelborghs and Dirk Roose. Numerical computation of stability and detection of Hopf bifurcations of steady state solutions of delay differential equations. Advances in Computational Mathematics, 10(3–4):271–289, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018986817622.

Elgindy:2013:OGO

[ESM13] Kareem T. Elgindy and Kate A.

Smith-Miles. On the optimization of Gegenbauer operational matrix of integration. Advances in Computational Mathematics, 39(3–4):511–524, October 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-012-9289-5.

Farouki:2002:HIR

[FaKS02] Rida T. Farouki, Mohammad al Kandari, and Takis Sakkalis. Hermite interpolation by rotation-invariant spatial Pythagorean-hodograph curves. Advances in Computational Mathematics, 17 (4):369–383, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1016280811626.

Farouki:2010:QHM

[Far10] Rida T. Farouki. Quaternion and Hopf map characterizations for the existence of rational rotation-minimizing frames on quintic space curves. Advances in Computational Mathematics, 33(3):331–348, October 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9138-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        009-9138-3.pdf.

Fasshauer:1999:HIR

[Fas99a] Gregory E. Fasshauer. Hermite interpolation with radial basis functions on spheres. Advances in Computational Mathematics, 10(1):81–96, January 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018914229009.

Fasshauer:1999:SDE

[Fas99b] Gregory E. Fasshauer. Solving differential equations with radial basis functions: multilevel methods and smoothing. Advances in Computational Mathematics, 11(2– 3):139–159, November 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018919824891. Ra-

dial basis functions and their applications.

Floater:2013:SCM

[FBCR13] M. Floater, C. Beccari, T. Cashman, and L. Romani. A smoothness criterion for monotonicitypreserving subdivision. Advances in Computational Mathematics, 39(1):193–204, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-012-9275-y.

Fischer:2015:APM

[FE15] Michael Fischer and Peter Eberhard. Application of parametric model reduction with matrix interpolation for simulation of moving loads in elastic multibody systems. Advances in Computational Mathematics, 41(5):1049–1072, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9379-7.

Farhloul:1995:MFE

[FF95] Mohamed Farhloul and Michel Fortin. A mixed finite element for the Stokes problem using quadrilateral elements. Advances in Computational Mathematics, 3(1–2):101–113, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02431998.

Fornberg:2005:ARB

[FF05] Bengt Fornberg and Natasha Flyer. Accuracy of radial basis function interpolation and derivative approximations on 1-D infinite grids. Advances in Computational Mathematics, 23(1–2):5–20, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1812-x.

Farouki:2017:CCS

[FGG^{+}17] Rida T. Farouki, Graziano Gentili, Carlotta Giannelli, Alessandra Sestini, and Caterina Stoppato. A comprehensive characterization of the set of polynomial curves with rational rotation-minimizing frames. Advances in Computational Mathematics, 43(1):1–24, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9473-0.

Feichtinger:2014:ADG

[FGO14] Hans G. Feichtinger, Anna Grybos, and Darian M. Onchis. Approximate dual Gabor atoms via the adjoint lattice method. Advances in Computational Mathematics, 40(3):651–665, June 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9324-1.

Farouki:2013:ISD

[FGS13] Rida T. Farouki, Carlotta Giannelli, and Alessandra Sestini. An interpolation scheme for designing rational rotationminimizing camera motions. Advances in Computational Mathematics, 38(1):63–82, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9226-z.

Farouki:2016:LMP

[FGS16] Rida T. Farouki, Carlotta Giannelli, and Alessandra Sestini. Local modification of Pythagorean-hodograph quintic spline curves using the B-spline form. Advances in Computational Mathematics, 42 (1):199–225, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9419-y.

Farouki:2005:SEE

[FH05] Rida T. Farouki and Chang Yong Han. Solution of elementary equations in the Minkowski geometric algebra of complex sets.

Advances in Computational Mathematics, 22(4):301–323, May 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-2605-3.

Fuhrer:2018:WPM

[FH18] Thomas Führer and Norbert Heuer. A wirebasket preconditioner for the mortar boundary element method. Advances in Computational Mathematics, 44(1):23–49, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9534-z.

Floater:2006:GCB

[FHK06] Michael S. Floater, Kai Hormann, and Géza Kós. A general construction of barycentric coordinates over convex polygons. Advances in Computational Mathematics, 24 (1–4):311–331, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7611-6.

Franke:1994:LSS

[FHN94] Richard Franke, Hans Hagen, and Gregory M. Nielson. Least squares surface approximation to scattered data using multiquadratic functions. Advances in Computational Mathematics, 2(1):81–99, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02519037.

Fasshauer:1999:MAA

[FJ99] Gregory E. Fasshauer and Joseph W. Jerome. Multistep approximation algorithms: improved convergence rates through postconditioning with smoothing kernels. Advances in Computational Mathematics, 10(1):1–27, January 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018962112170.

Fairweather:1998:MFS

[FK98] Graeme Fairweather and Andreas Karageorghis. The method of fundamental solutions for elliptic boundary value problems. Advances in Computational Mathematics, 9 (1–2):69–95, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018981221740. Nu-

merical treatment of boundary integral equations.

Floater:2010:IWM

[FK10] Michael S. Floater and Jiří Kosinka. On the injectivity of Wachspress and mean value mappings between convex polygons. Advances in Computational Mathematics, 32 (2):163–174, February 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9098-z.

Fairweather:2005:MDM

[FKS05] Graeme Fairweather, Andreas Karageorghis, and Yiorgos-Sokratis Smyrlis. A matrix decomposition MFS algorithm for axisymmetric biharmonic problems. Advances in Computational Mathematics, 23(1–2):55–71, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1808-6.

Floater:2000:ACD

[FL00] Michael S. Floater and Tom Lyche. Asymptotic convergence of degree-raising. Advances in Computational Mathematics, 12(2–3):175–187, February 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018913118047.

Flinth:2018:SBD

[Fli18] Axel Flinth. Sparse blind deconvolution and demixing through ℓ_{1}, 2-minimization. Advances in Computational Mathematics, 44(1):1–21, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9533-0.

Lin:2005:EHT

[fLL05] Jia fu Lin and Qun Lin.

Extrapolation of the Hood– Taylor elements for the Stokes problem. Advances in Computational Mathematics, 22 (2):115–123, February 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1089-0.

Fairweather:1996:GMS

[FLM96] G. Fairweather and J. C.

López-Marcos. Galerkin methods for a semilinear parabolic problem with nonlocal boundary conditions. Advances in Computational Mathematics, 6 (1):243–262, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127706.  John

Crank 80th birthday special issue.

Floater:1994:WCC

[Flo94] Michael S. Floater. A weak condition for the convexity of tensor-product Bézier and B-spline surfaces. Advances in Computational Mathematics, 2(1):67–80, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02519036.

Floater:1996:ACA

[Flo96a] Michael S. Floater. An analysis of cubic approximation schemes for conic sections. Advances in Computational Mathematics, 5(1):361–379, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124751.

Floater:1996:ZCB

[Flo96b] Michael S. Floater. On zero curves of bivariate polynomials. Advances in Computational Mathematics, 5 (1):399–415, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124753.

Freeden:2004:OZT

[FM04] W. Freeden and V. Michel. Orthogonal zonal, tesseral and sectorial wavelets on the sphere for the analysis of satellite data. Advances in Computational Mathematics, 21(1–2):181–217, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/B:ACOM.0000016429.
                                        07498.77.

Fornberg:2014:SHB

[FM14] Bengt Fornberg and Jordan M. Martel. On spherical harmonics based numerical quadrature over the surface of a sphere. Advances in Computational Mathematics, 40(5– 6):1169–1184, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9346-3.

Fickus:2013:CAS

[FMPS13] Matthew Fickus, Dustin G.

Mixon, Miriam J. Poteet, and Nate Strawn. Constructing all self-adjoint matrices with prescribed spectrum and diagonal. Advances in Computational Mathematics, 39 (3–4):585–609, October 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9298-z.

Farouki:2000:AMP

[FMR00] Rida T. Farouki, Hwan Pyo Moon, and Bahram Ravani. Algorithms for Minkowski products and implicitly-defined complex sets. Advances in Computational Mathematics, 13(3):199–229, July 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018910412112.

Feichtinger:2014:CAD

[FOW14] Hans G. Feichtinger, Darian M.

Onchis, and Christoph Wiesmeyr. Construction of approximate dual wavelet frames. Advances in Computational Mathematics, 40(1):273–282, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9307-2.

Floater:1998:TPM

[FP98] Michael S. Floater and J. M. Peña. Tensor-product monotonicity preservation. Advances in Computational Mathematics, 9(3–4):353–362, November 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018906027191.

Faul:1999:PCI

[FP99] A. C. Faul and M. J. D. Powell. Proof of convergence of an iterative technique for thin plate spline interpolation in two dimensions. Advances in Computational Mathematics, 11(2–3):183–192, November 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018923925800. Ra-

dial basis functions and their applications.

Feichtinger:2012:ARB

[FPR^{+}12] Hans G. Feichtinger, José C. Príncipe, José Luis Romero, Alexander Singh Alvarado, and Gino Angelo Velasco. Approximate reconstruction of bandlimited functions for the integrate and fire sampler. Advances in Computational Mathematics, 36(1):67–78, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9180-9.

Floater:2006:CCM

[FPT06] Michael S. Floater and Valérie Pham-Trong. Convex combination maps over triangulations, tilings, and tetrahedral meshes. Advances in Computational Mathematics, 25 (4):347–356, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7620-5.

Fu:2010:NPO

[FQ10] Chu-Li Fu and Zhi Qian. Numerical pseudodifferential operator and Fourier regularization. Advances in Computational Mathematics, 33 (4):449–470, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9136-5.

Fermo:2010:NMF

[FR10] L. Fermo and M. G. Russo.

Numerical methods for Fredholm integral equations with singular right-hand sides. Advances in Computational Mathematics, 33(3):305–330, October 2010. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9137-4.

Franssens:1999:NNP

[Fra99] Ghislain Franssens. A new nonpolynomial univariate interpolation formula of Hermite type. Advances in Computational Mathematics, 10(3–4):367–388, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018947103510.

Fornasier:2009:AJS

[FRT09] Massimo Fornasier, Ronny Ramlau, and Gerd Teschke. The application of joint sparsity and total variation minimization algorithms to a reallife art restoration problem. Advances in Computational Mathematics, 31(1–3):157–184, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9103-6.

Farouki:1994:PHS

[FS94] Rida T. Farouki and Takis Sakkalis. Pythagorean-hodograph space curves. Advances in Computational Mathematics, 2(1): 41–66, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1007/BF02519035.

Franke:1998:COE

[FS98] Carsten Franke and Robert Schaback. Convergence order estimates of meshless collocation methods using radial basis functions. Advances in Computational Mathematics, 8(4):381–399, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018916902176.

Feichtinger:2007:SCI

[FS07] Hans G. Feichtinger and Wenchang Sun. Sufficient conditions for irregular Gabor frames. Advances in Computational Mathematics, 26(4):403–430, May 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7210-6.

Fuhr:2008:SEV

[Füh08] H. Führ. Simultaneous estimates for vector-valued Gabor frames of Hermite functions. Advances in Computational Mathematics, 29(4):357–373, November 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9053-4.

Fuhr:2016:VMC

[Füh16] Hartmut Führ. Vanishing moment conditions for wavelet atoms in higher dimensions. Advances in Computational Mathematics, 42(1):127–153, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9414-3.

Fuselier:2008:EIS

[Fus08a]  Edward J. Fuselier.  Erra-

tum: “Improved stability estimates and a characterization of the native space for matrixvalued RBFs” [Adv. Comput. Math. 29 (2008), no. 3, 269– 290; MR2438345]. Advances in Computational Mathematics, 29(3):311–313, October 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9091-6;
 http://link.springer.com/
content/pdf/10.1007/s10444-
008-9091-6.pdf. See [Fus08b].

Fuselier:2008:ISE

[Fus08b] Edward J. Fuselier. Improved stability estimates and a characterization of the native space for matrix-valued RBFs. Advances in Computational Mathematics, 29(3):269–290, October 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9046-3.

See erratum [Fus08a].

Freeden:1996:SWT

[FW96] Willi Freeden and Ulrich Windheuser. Spherical wavelet transform and its discretization. Advances in Computational Mathematics, 5(1):51–94, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124735.

Fuselier:2015:OPD

[FW15] Edward Fuselier and Grady B.

Wright. Order-preserving derivative approximation with periodic radial basis functions. Advances in Computational Mathematics, 41(1):23– 53, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
014-9348-1.

Feng:2006:GFT

[FWW06] De-Jun Feng, Long Wang, and Yang Wang. Generation of finite tight frames by Householder transformations. Advances in Computational Mathematics, 24(1–4):297–309, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7637-9.

Ford:1996:DCQ

[FXZ96] William F. Ford, Yuesheng Xu, and Yunhe Zhao. Derivative corrections for quadrature formulas. Advances in Computational Mathematics, 6 (1):139–157, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02127701.

Fasshauer:2013:RKS

[FY13] Gregory E. Fasshauer and Qi Ye. Reproducing kernels of Sobolev spaces via a green kernel approach with differential operators and boundary operators. Advances in Computational Mathematics, 38(4):891–921, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9264-6.

Feng:2008:AFE

[FYL08] Tao Feng, Ningning Yan, and Wenbin Liu. Adaptive finite element methods for the identification of distributed parameters in elliptic equation. Advances in Computational Mathematics, 29(1):27–53, July 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9035-6.

Fang:2012:RIB

[FZ12] Weifu Fang and Suxing Zeng. Recovery of an interface from boundary measurement in an elliptic differential equation. Advances in Computational Mathematics, 36(2):201–220, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9204-5.

Goh:2006:ITC

[GG06] Say Song Goh and Tim N. T.

Goodman. Inequalities on time-concentrated or frequencyconcentrated functions. Advances in Computational Mathematics, 24(1–4):333–351, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-4145-x.

Gimbutas:2016:FMM

[GG16] Zydrunas Gimbutas and Leslie Greengard. A fast multipole method for the evaluation of elastostatic fields in a half-space with zero normal stress. Advances in Computational Mathematics, 42(1):175–198, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9416-1.

Goh:2007:CPR

[GGL07] Say Song Goh, Tim N. T. Goodman, and S. L. Lee. Causality properties of refinable functions and sequences. Advances in Computational Mathematics, 26(1–3):231–250, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-8007-3.

Goktas:1999:ACG

[GH99] Ünal Göktaş and Willy Hereman. Algorithmic computation of generalized symmetries of nonlinear evolution and lattice equations. Advances in Computational Mathematics, 11(1):55–80, July 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018955405327.

Gabardo:2003:FAM

[GH03] Jean-Pierre Gabardo and Deguang Han. Frames associated with measurable spaces. Advances in Computational Mathematics, 18(2–4):127–147, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1021312429186. Frames.

Guo:2006:OCV

[GH06] Benqi Guo and Norbert Heuer. The optimal convergence of the h–p version of the boundary element method with quasiuniform meshes for elliptic problems on polygonal domains. Advances in Computational Mathematics, 24(1–4):353–374, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7618-z.

Guglielmi:2008:CBP

[GH08] Nicola Guglielmi and Ernst Hairer. Computing breaking points in implicit delay differential equations. Advances in Computational Mathematics, 29(3):229–247, October 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9044-5.

Grohs:2016:SAL

[GH16] P. Grohs and S. Hosseini. εsubgradient algorithms for locally Lipschitz functions on Riemannian manifolds. Advances in Computational Mathematics, 42(2):333–360, April 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9426-z.

Garrigos:2012:DFO

[GHdN12] Gustavo Garrigós, Eugenio Hernández, and Maria de Natividade. Democracy functions and optimal embeddings for approximation spaces. Advances in Computational Mathematics, 37(2):255–283, August 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9197-0.

Gillman:2017:IET

[Gil17] Adrianna Gillman. An integral equation technique for scattering problems with mixed boundary conditions. Advances in Computational Mathematics, 43(2):351–364, April 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9488-6.

Giannelli:2014:SSB

[GJS14] Carlotta Giannelli, Bert Jüttler, and Hendrik Speleers. Strongly stable bases for adaptively refined multilevel spline spaces. Advances in Computational Mathematics, 40(2):459–490, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9315-2.

Guo:2013:ORX

[GL13] Kanghui Guo and Demetrio Labate. Optimal recovery of 3D X-ray tomographic data via shearlet decomposition. Advances in Computational Mathematics, 39(2):227–255, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-012-9276-x.

Guo:2017:MAE

[GL17] Kanghui Guo and Demetrio Labate. Microlocal analysis of edge flatness through directional multiscale representations. Advances in Computational Mathematics, 43(2):295–318, April 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9486-8.

Gracia:2010:CSS

[GLO10] J. L. Gracia, F. J. Lisbona, and E. O’Riordan. A coupled system of singularly perturbed parabolic reaction– diffusion equations. Advances in Computational Mathematics, 32(1):43–61, January 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9086-3.

Gerbeau:2015:ROM

[GLS15] Jean-Frédéric Gerbeau, Damiano Lombardi, and Elisa Schenone. Reduced order model in cardiac electrophysiology with approximated Lax pairs. Advances in Computational Mathematics, 41 (5):1103–1130, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9393-9.

Goodman:1993:WBS

[GLT93] T. N. T. Goodman, S. L. Lee, and W. S. Tang. Wavelet bases for a set of commuting unitary operators. Advances in Computational Mathematics, 1(1):109–126, February 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02070823.

Gillman:2014:ACS

[GM14] Adrianna Gillman and Per-Gunnar Martinsson. An O(N) algorithm for constructing the solution operator to 2D elliptic boundary value problems in the absence of body loads. Advances in Computational Mathematics, 40(4):773–796, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9326-z.

Ghelardoni:2017:MMF

[GM17] Paolo Ghelardoni and Cecilia Magherini. A matrix method for fractional Sturm– Liouville problems on bounded domain. Advances in Computational Mathematics, 43(6): 1377–1401, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9529-9.

Graves-Morris:1997:CSI

[GMB97] P. R. Graves-Morris and B. Beckermann. The compass (star) identity for vector-valued rational interpolants. Advances in Computational Mathematics, 7(3):279–294, May 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018951020406.

Gatica:2009:NCM

[GMM09] Gabriel N. Gatica, Antonio Márquez, and Salim Meddahi. A new coupling of mixed finite element and boundary element methods for an exterior Helmholtz problem in the plane. Advances in Computational Mathematics, 30(3):281–301, April 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9068-5.

Goodman:1997:SFL

[GMRS97] Tim N. T. Goodman, Charles A.

Micchelli, Giuseppe Rodriguez, and Sebastiano Seatzu. Spectral factorization of Laurent polynomials. Advances in Computational Mathematics, 7(4):429–454, August 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018915407202.

Guessab:2012:CNF

[GMS12] Allal Guessab, María Moncayo, and Gerhard Schmeisser. A class of nonlinear four-point subdivision schemes. Advances in Computational Mathematics, 37(2):151–190, August 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9199-y.

Gribonval:2001:AWG

[GN01] R. Gribonval and M. Nielsen. Approximate weak greedy algorithms. Advances in Computational Mathematics, 14(4): 361–378, May 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1023/A:1012255021470.

Gribonval:2008:BSR

[GN08] Rémi Gribonval and Morten Nielsen. Beyond sparsity: recovering structured representations by ℓ^{1} minimization and greedy algorithms. Advances in Computational Mathematics, 28(1):23–41, January 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-005-9009-5.

Gupta:2013:PEE

[GN13] Nupur Gupta and Neela Nataraj. A posteriori error estimates for an optimal control problem of laser surface hardening of steel. Advances in Computational Mathematics, 39(1):69–99, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9270-8.

Gnaneshwar:2007:DKM

[Gna07] N. Gnaneshwar. A degenerate kernel method for eigenvalue problems of compact integral operators. Advances in Computational Mathematics, 27(3):339–354, October 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-9005-9.

Griebel:1995:TPT

[GO95] M. Griebel and P. Oswald.

Tensor product type subspace splittings and multilevel iterative methods for anisotropic problems. Advances in Computational Mathematics, 4(1– 2):171–206, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02123478.

Goodman:2005:SPI

[GO05] T. N. T. Goodman and B. H.

Ong. Shape-preserving interpolation by splines using vector subdivision. Advances in Computational Mathematics, 22(1):49–77, January 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-1174-z.

Gomes:1995:CEW

[Gom95]  Sônia M. Gomes.  Conver-

gence estimates for the wavelet-Galerkin method: superconvergence at the node points. Advances in Computational Mathematics, 4(1–2):261–282, December 1995. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF03177516.

Gonsor:1993:SAN

[Gon93] Daniel E. Gonsor. Subdivision algorithms with nonnegative masks generally converge. Advances in Computational Mathematics, 1(2):215–221, June 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02071386.

Guo:2003:P

[GOZ03] Benqi Guo, Hae-Soo Oh, and Zhimin Zhang. Preface. Advances in Computational Mathematics, 19(1–3):1–2, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1022844220807.

Gasca:1995:CAS

[GP95] M. Gasca and J. M. Peña. On the characterization of almost strictly totally positive matrices. Advances in Computational Mathematics, 3(3):239–250, April 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02432001.

Garrappa:2013:EGM

[GP13] Roberto Garrappa and Marina Popolizio. Evaluation of generalized Mittag-Leffler functions on the real line. Advances in Computational Mathematics, 39(1):205–225, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-012-9274-z.

Grochenig:2014:URT

[GP14] Karlheinz Gröchenig and Elmar Pauwels. Uniqueness and reconstruction theorems for pseudodifferential operators with a bandlimited Kohn– Nirenberg symbol. Advances in Computational Mathematics, 40(1):49–63, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9297-0.

Guy:2015:GMI

[GPG15] Robert D. Guy, Bobby Philip, and Boyce E. Griffith. Geometric multigrid for an implicit-time immersed boundary method. Advances in Computational Mathematics, 41(3):635–662, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9380-1.

Garcia-Puente:2010:LPP

[GPS10] Luis David Garcia-Puente and Frank Sottile. Linear precision for parametric patches. Advances in Computational Mathematics, 33(2):191–214, August 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9126-7.

Gerhards:2017:PCS

[GPT17] Christian Gerhards, Sergiy Pereverzyev, Jr., and Pavlo Tkachenko. A parameter choice strategy for the inversion of multiple observations. Advances in Computational Mathematics, 43(1):101–112, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9477-9.

Graf:2012:UAS

[Grä12] Manuel Gräf. A unified approach to scattered data approximation on S^{3} and SO(3). Advances in Computational Mathematics, 37(3):379–392, October 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9214-3.

Gillette:2012:EEG

[GRB12] Andrew Gillette, Alexander Rand, and Chandrajit Bajaj. Error estimates for generalized barycentric interpolation. Advances in Computational Mathematics, 37(3):417–439, October 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9218-z.

Grochenig:2003:LFF

[Grö03] Karlheinz Gröchenig. Localized frames are finite unions of Riesz sequences. Advances in Computational Mathematics, 18(2–4):149–157, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021368609918. Frames.

Grohs:2013:RFD

[Gro13] Philipp Grohs. Refinable functions for dilation families. Advances in Computational Mathematics, 38(3):531–561, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9248-6.

Gasca:2000:PIS

[GS00] Mariano Gasca and Thomas Sauer. Polynomial interpolation in several variables. Advances in Computational Mathematics, 12(4):377–410, March 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018981505752. Mul-

tivariate polynomial interpolation.

Guessab:2010:CPD

[GS10] Allal Guessab and Gerhard Schmeisser. Construction of positive definite cubature formulae and approximation of functions via Voronoi tessellations. Advances in Computational Mathematics, 32(1):25– 41, January 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
008-9080-9.

Greengard:2012:E

[GS12] Leslie Greengard and Michael Shelley. Editorial. Advances in Computational Mathematics, 36(1):1, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9267-3;
 http://link.springer.com/
content/pdf/10.1007/s10444-
011-9267-3.pdf.

Guo:2013:LCB

[GS13] Zheng-Chu Guo and Lei Shi. Learning with coefficient-based regularization and ℓ^{1}-penalty. Advances in Computational Mathematics, 39(3–4):493–510, October 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
012-9288-6.

Garcia:2014:NOV

[GT14] Galina C. García and Takéo Takahashi. Numerical observers with vanishing viscosity for the 1 d wave equation. Advances in Computational Mathematics, 40(4):711–745, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9320-5.

Tang:2002:SMI

[gTpM02] Jian guo Tang and He ping Ma. Single and multi-interval Legendre τ-methods in time for parabolic equations. Advances in Computational Mathematics, 17(4):349–367, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016273820035.

Gravesen:2002:CIF

[GU02] Jens Gravesen and Michael Ungstrup. Constructing invariant fairness measures for surfaces. Advances in Computational Mathematics, 17(1–2): 67–88, July 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1023/A:1015229622042.

Advances in geometrical algorithms and representations.

Gu:2017:PSC

[Gu17]  Zhendong Gu.  Piecewise

spectral collocation method for system of Volterra integral equations. Advances in Computational Mathematics, 43(2):385–409, April 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9490-z.

Gonzalez-Vera:1996:SRA

[GVSLN96] P. González-Vera, J. C. Santos-León, and O. Njåstad. Some results about numerical quadrature on the unit circle. Advances in Computational Mathematics, 5(1):297–328, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124749.

Guo:2014:CMG

[GW14] Ben-Yu Guo and Zhong-Qing Wang. A collocation method for generalized nonlinear Klein– Gordon equation. Advances in Computational Mathematics, 40(2):377–398, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9312-5.

Gao:2017:CRK

[GW17] Wenwu Gao and Zongmin Wu. Constructing radial kernels with higher-order generalized Strang–Fix conditions. Advances in Computational Mathematics, 43(6): 1355–1375, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9528-x.

Goodman:1999:IM

[GY99] Tim N. T. Goodman and Thomas P. Y. Yu. Interpolation of medians. Advances in Computational Mathematics, 11(1):1–10, July 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018959506235.

Goh:2000:UPL

[GY00] Say Song Goh and Chee Heng Yeo. Uncertainty products of local periodic wavelets. Advances in Computational Mathematics, 13(3):319–333, July 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018962428951.

Yi:2012:GJR

[gYyG12] Yong gang Yi and Ben yu Guo.

Generalized Jacobi rational spectral method on the half line. Advances in Computational Mathematics, 37(1):1–37, July 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9193-4.

Guo:2017:ORL

[GYZ17] Zheng-Chu Guo, Yiming Ying, and Ding-Xuan Zhou. Online regularized learning with pairwise loss functions. Advances in Computational Mathematics, 43(1):127–150, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9479-7.

Guo:2013:CEL

[GZ13] Zheng-Chu Guo and Ding-Xuan Zhou. Concentration estimates for learning with unbounded sampling. Advances in Computational Mathematics, 38(1):207–223, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9238-8.

Hagger:1998:ADD

[Hag98] M. J. Hagger. Automatic domain decomposition on unstructured grids (DOUG). Advances in Computational Mathematics, 9(3–4):281–310, November 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018997725374.

Hakopian:2000:CHI

[Hak00]  Hakop A. Hakopian.  On

a class of Hermite interpolation problems. Advances in Computational Mathematics, 12(4):303–309, March 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018933622590. Mul-

tivariate polynomial interpolation.

Hall:1995:NSS

[Hal95] George Hall. A new stepsize strategy for explicit Runge– Kutta codes. Advances in Computational Mathematics, 3(4): 343–352, May 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/BF03028367.

Han:1998:SOS

[Han98]  Bin Han.  Symmetric or-

thonormal scaling functions and wavelets with dilation factor

4. Advances in Computational Mathematics, 8(4):221–247, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018948314471.

Hansen:2005:FCM

[Han05] Olaf Hansen. A fast collocation method for the radiosity equation, based on the hierarchical algorithm of Hanrahan and Salzman: the 1 d case. Advances in Computational Mathematics, 22(4):353–375, May 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-3962-7.

Han:2006:SSS

[Han06] Bin Han. Solutions in Sobolev spaces of vector refinement equations with a general dilation matrix. Advances in Computational Mathematics, 24 (1–4):375–403, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7615-2.

Han:2010:SOC

[Han10]  Bin Han.  Symmetric or-

thonormal complex wavelets with masks of arbitrarily high linear-phase moments and sum rules. Advances in Computational Mathematics, 32 (2):209–237, February 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9102-7.

Hangelbroek:2012:LRA

[Han12]  T. Hangelbroek.  On local

RBF approximation. Advances in Computational Mathematics, 37(2):285–299, August 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9212-5.

Hollig:2005:IWM

[HAS05] Klaus Höllig, Christian Apprich, and Anja Streit. Introduction to the Web-method and its applications. Advances in Computational Mathematics, 23(1–2):215–237, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1811-y.

Hauber:1997:NTR

[Hau97] R. Hauber. Numerical treatment of retarded differential– algebraic equations by collocation methods. Advances in Computational Mathematics, 7(4):573–592, August 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018919508111.

Hoefkens:2003:CVS

[HBM03] Jens Hoefkens, Martin Berz, and Kyoko Makino. Computing validated solutions of implicit differential equations. Advances in Computational Mathematics, 19(1–3):231–253, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1022858921155. Chal-

lenges in computational mathematics (Pohang, 2001).

Hao:2014:HOA

[HBMY14] S. Hao, A. H. Barnett, P. G.

Martinsson, and P. Young. High-order accurate methods for Nyström discretization of integral equations on smooth curves in the plane. Advances in Computational Mathematics, 40(1):245–272, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9306-3.

Hemker:1995:SGF

[Hem95] P. W. Hemker. Sparse-grid finite-volume multigrid for 3Dproblems. Advances in Computational Mathematics, 4(1– 2):83–110, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02123474. ???

Hemker:1996:CSG

[Hem96]  P. W. Hemker.  Corrigen-

dum: “Sparse-grid finite-volume multigrid for 3D-problems” [Adv. Comput. Math. 4 (1995), no. 1–2, 83–110; MR1338896 (96c:65193)]. Advances in Computational Mathematics, 5 (1):125–126, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124737;   http:
                                        //link.springer.com/content/
                                        pdf/10.1007/BF02124737.pdf.

See [Hem95].

Hesse:2009:ESD

[Hes09] Kerstin Hesse. The s-energy of spherical designs on S^{2}. Advances in Computational Mathematics, 30(1):37–59, January 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9057-0.

Hass:2007:GCS

[HFH^{+}07] Joel Hass, Rida T. Farouki, Chang Yong Han, Xiaowen Song, and Thomas W. Sederberg. Guaranteed consistency of surface intersections and trimmed surfaces using a coupled topology resolution and domain decomposition scheme. Advances in Computational Mathematics, 27(1):1–26, July 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7539-5.

Huang:2015:ETS

[HFH15] Pengzhan Huang, Xinlong Feng, and Yinnian He. An efficient two-step algorithm for the incompressible flow problem. Advances in Computational Mathematics, 41(6): 1495–1513, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9400-1.

Helsing:2015:VEK

[HH15] Johan Helsing and Anders Holst. Variants of an explicit kernel-split panel-based Nyström discretization scheme for Helmholtz boundary value problems. Advances in Computational Mathematics, 41(3): 691–708, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
014-9383-y.

He:2016:DDR

[HH16] Hongjin He and Deren Han. A distributed Douglas–Rachford splitting method for multi-block convex minimization problems. Advances in Computational Mathematics, 42(1):27– 53, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1007/s10444-
                                        015-9408-1.

Hardin:2004:MVR

[HHS04] Douglas P. Hardin, Thomas A. Hogan, and Qiyu Sun. The matrix-valued Riesz lemma and local orthonormal bases in shiftinvariant spaces. Advances in Computational Mathematics, 20(4):367–384, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1027389826705.

Hiptmair:2012:MTB

[HJH12] R. Hiptmair and C. Jerez-Hanckes. Multiple traces boundary integral formulation for Helmholtz transmission problems. Advances in Computational Mathematics, 37(1):39–91, July 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9194-3.

Hu:1997:ROC

[HJJV97] Min Hu, Ken Jackson, Jan Janssen, and Stefan Vandewalle. Remarks on the optimal convolution kernel for CSOR waveform relaxation. Advances in Computational Mathematics, 7(1–2):135–156, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018938617680. Par-

allel methods for ODEs.

Healy:2004:TSE

[HKR04] D. M. Healy, Jr., P. J. Kostelec, and D. Rockmore. Towards safe and effective high-order Legendre transforms with applications to FFTs for the 2-sphere. Advances in Computational Mathematics, 21(1–2):59–105, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000016431.
03652.65.

Haslinger:2014:DDM

[HKRS14] Jaroslav Haslinger, Radek Kučera, Julien Riton, and Taoufik Sassi. A domain decomposition method for twobody contact problems with Tresca friction. Advances in Computational Mathematics, 40(1):65–90, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9299-y.

Hansen:1996:LUC

[HL96] Lars Kai Hansen and Jan Larsen. Linear unlearning for cross-validation. Advances in Computational Mathematics, 5 (1):269–280, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124747.

Han:2005:FPA

[HL05] Weimin Han and Wing Kam Liu. Flexible piecewise approximations based on partition of unity. Advances in Computational Mathematics, 23(1–2):191–199, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1810-z.

He:2006:MLS

[HL06] Yinnian He and Kam-Moon Liu. Multi-level spectral Galerkin method for the Navier–Stokes equations. II. Time discretization. Advances in Computational Mathematics, 25 (4):403–433, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7640-1.

Hesse:2008:CES

[HL08] Kerstin Hesse and Paul Leopardi. The Coulomb energy of spherical designs on S^{2}. Advances in Computational Mathematics, 28(4):331–354, May 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9026-7.

Hamm:2016:CIG

[HL16a] Keaton Hamm and Jeff Ledford. Cardinal interpolation with general multiquadrics. Advances in Computational Mathematics, 42(5):1149–1186, October 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9457-0.

Ho:2016:QQC

[HL16b] Kin Tat Ho and Lok Ming Lui. QCMC: quasi-conformal parameterizations for multiplyconnected domains. Advances in Computational Mathematics, 42(2):279–312, April 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9424-1.

Hintermuller:2012:SOT

[HLN12] M. Hintermüller, A. Laurain, and A. A. Novotny. Secondorder topological expansion for electrical impedance tomography. Advances in Computational Mathematics, 36 (2):235–265, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9205-4.

Han:2011:FTA

[HLT11] Deguang Han, Pengtong Li, and Wai-Shing Tang. Frames and their associated H_{F}^{p} - subspaces. Advances in Computational Mathematics, 34 (2):185–200, February 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9149-0.

Han:2003:MFR

[HM03] Bin Han and Qun Mo. Multiwavelet frames from refinable function vectors. Advances in Computational Mathematics, 18(2–4):211–245, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021360312348. Frames.

Howlett:1996:F

[HN96] Jack Howlett and J. C. Newby. Foreword. Advances in Computational Mathematics, 6(1):iii– vi, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/BF02127692.

Himpe:2015:DDC

[HO15] Christian Himpe and Mario Ohlberger. Data-driven combined state and parameter reduction for inverse problems. Advances in Computational Mathematics, 41 (5):1343–1364, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9420-5.

Hegarty:2017:PUN

[HO17] A. F. Hegarty and E. O’Riordan.

Parameter-uniform numerical method for singularly perturbed convection–diffusion problem on a circular domain. Advances in Computational Mathematics, 43(5):885–909, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9510-z.

Hoffman:2006:GVI

[Hof06] Alan J. Hoffman. Gersgorin variations. II. On themes of Fan and Gudkov. Advances in Computational Mathematics, 25(1–3):1–6, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7624-1.

Holst:2001:ANT

[Hol01] M. Holst. Adaptive numerical treatment of elliptic systems on manifolds. Advances in Computational Mathematics, 15(1–4):139–191, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014246117321. A

posteriori error estimation and adaptive computational methods.

Howlett:1996:FJC

[How96] Jack Howlett. Foreword: John Crank. Advances in Computational Mathematics, 6(1):vii– viii, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). John Crank 80th birthday special issue.

Heuveline:2001:PEC

[HR01] Vincent Heuveline and Rolf Rannacher. A posteriori error control for finite element approximations of elliptic eigenvalue problems. Advances in Computational Mathematics, 15 (1–4):107–138, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014291224961. A

posteriori error estimation and adaptive computational methods.

Hsiao:2002:WCM

[HR02] George C. Hsiao and Andreas Rathsfeld. Wavelet collocation methods for a first kind boundary integral equation in acoustic scattering. Advances in Computational Mathematics, 17(4):281–308, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016249415373.

Hearn:2013:EJR

[HR13] Tristan A. Hearn and Lothar Reichel. Extensions of the Justen–Ramlau blind deconvolution method. Advances in Computational Mathematics, 39 (3–4):465–491, October 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-012-9290-z.

Holhos:2017:APM

[HR17] Adrian Holhoş and Daniela Roşca. Area preserving maps and volume preserving maps between a class of polyhedrons and a sphere. Advances in Computational Mathematics, 43(4):677–697, August 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9502-z.

Heuer:2003:ODD

[HS03] Norbert Heuer and Ernst P.

Stephan. An overlapping domain decomposition preconditioner for high order BEM with anisotropic elements. Advances in Computational Mathematics, 19(1–3):211–230, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1022899205324. Chal-

lenges in computational mathematics (Pohang, 2001).

Hon:2008:SPD

[HS08] Y. C. Hon and Robert Schaback. Solvability of partial differential equations by meshless kernel methods. Advances in Computational Mathematics, 28(3):283–299, April 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-006-9023-2.

Hon:2013:SLE

[HS13] Y. C. Hon and R. Schaback. Solving the 3D Laplace equation by meshless collocation via harmonic kernels. Advances in Computational Mathematics, 38(1):1–19, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9224-1.

Huang:2016:SCM

[HS16] Can Huang and Martin Stynes. A spectral collocation method for a weakly singular Volterra integral equation of the second kind. Advances in Computational Mathematics, 42 (5):1015–1030, October 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9451-6.

Hong:2009:ELC

[HSV09] Byung-Woo Hong, Stefano Soatto, and Luminita A. Vese. Enforcing local context into shape statistics. Advances in Computational Mathematics, 31 (1–3):185–213, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9104-5.

Hu:2016:STD

[HSY16] Jun Hu, Zhongci Shi, and Xueqin Yang. Superconvergence of three dimensional Morley elements on cuboid meshes for biharmonic equations. Advances in Computational Mathematics, 42(6):1453–1471, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9470-3.

Hon:2011:DTR

[HT11] Y. C. Hon and Tomoya Takeuchi. Discretized Tikhonov regularization by reproducing kernel Hilbert space for backward heat conduction problem. Advances in Computational Mathematics, 34 (2):167–183, February 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-010-9148-1.

He:2012:TLL

[HT12] Lianhua He and Lutz Tobiska. The two-level local projection stabilization as an enriched one-level approach. Advances in Computational Mathematics, 36(4):503–523, May 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9188-1.

Hu:2007:RDD

[Hu07] Qiya Hu. A regularized domain decomposition method with Lagrange multiplier. Advances in Computational Mathematics, 26(4):367–401, May 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-4094-4.

Huang:2006:CMB

[Hua06]  Min Huang.  A construc-

tion of multiscale bases for Petrov–Galerkin methods for integral equations. Advances in Computational Mathematics, 25(1–3):7–22, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-003-7607-7.

Hubbert:2012:CFR

[Hub12] Simon Hubbert. Closed form representations for a class of compactly supported radial basis functions. Advances in Computational Mathematics, 36(1):115–136, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9184-5.

Hickernell:2000:IAA

[HW00] F. J. Hickernell and H. Woźniakowski.

Integration and approximation in arbitrary dimensions. Advances in Computational Mathematics, 12(1):25–58, January 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018948631251. High

dimensional integration.

Hesse:2012:NIP

[HW12] Kerstin Hesse and Robert S.

Womersley. Numerical integration with polynomial exactness over a spherical cap. Advances in Computational Mathematics, 36(3):451–483, April 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9187-2.

Huang:2013:AEE

[HWZ13] Jin Huang, Zhu Wang, and Rui Zhu. Asymptotic error expansions for hypersingular integrals. Advances in Computational Mathematics, 38 (2):257–279, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9236-x.

Huang:2002:CFE

[HX02] Yunqing Huang and Weimin Xue. Convergence of finite element approximations and multilevel linearization for Ginzburg– Landau model of d-wave superconductors. Advances in Computational Mathematics, 17 (4):309–330, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016293508648.

Hou:2017:FSM

[HX17] Dianming Hou and Chuanju Xu. A fractional spectral method with applications to some singular problems. Advances in Computational Mathematics, 43(5):911–944, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9511-y.

Hu:2010:NMI

[HXC10] Xiaoyi Hu, Xiang Xu, and Wenbin Chen. Numerical method for the inverse heat transfer problem in composite materials with Stefan–Boltzmann conditions. Advances in Computational Mathematics, 33 (4):471–489, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9131-x.

Hu:2014:NPM

[HY14] Qiya Hu and Long Yuan. A Newton-penalty method for a simplified liquid crystal model. Advances in Computational Mathematics, 40 (1):201–244, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9305-4.

Hu:2018:PWM

[HY18] Qiya Hu and Long Yuan. A plane wave method combined with local spectral elements for nonhomogeneous Helmholtz equation and time-harmonic Maxwell equations. Advances in Computational Mathematics, 44(1):245–275, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9542-z.

Huang:2012:SEA

[HZHL12] Jin Huang, Guang Zeng, Xiaoming He, and Zi-Cai Li. Splitting extrapolation algorithm for first kind boundary integral equations with singularities by mechanical quadrature methods. Advances in Computational Mathematics, 36(1):79–97, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9181-8.

Iacono:2017:NAP

[IB17] Roberto Iacono and John P. Boyd. New approximations to the principal real-valued branch of the Lambert W - function. Advances in Computational Mathematics, 43(6): 1403–1436, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9530-3.

Ivanyshyn:2010:HPI

[IKS10] Olha Ivanyshyn, Rainer Kress, and Pedro Serranho. Huygens’ principle and iterative methods in inverse obstacle scattering. Advances in Computational Mathematics, 33 (4):413–429, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9135-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        009-9135-6.pdf.

Ionescu:2016:MSA

[IL16] Ioan R. Ionescu and Oana Lupaşcu. Modeling shallow avalanche onset over complex basal topography. Advances in Computational Mathematics, 42(1):5–26, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9407-2.

Ivanov:1999:BLA

[IMS99] Tjavdar Ivanov, Vladimir Maz’ya, and Gunther Schmidt. Boundary layer approximate approximations and cubature of potentials in domains. Advances in Computational Mathematics, 10(3–4):311–342, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018990918531.

Ivanov:2015:FME

[IP15] Kamen Ivanov and Pencho Petrushev. Fast memory efficient evaluation of spherical polynomials at scattered points. Advances in Computational Mathematics, 41 (1):191–230, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9354-3.

Ismail:2013:AGD

[IS13] Mourad E. H. Ismail and Plamen Simeonov. Asymptotics of generalized derangements. Advances in Computational Mathematics, 39(1):101–127, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9271-7.

Ito:1996:NCL

[Ito96] Yoshifusa Ito. Nonlinearity creates linear independence. Advances in Computational Mathematics, 5(1):189–203, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124743.

Iapichino:2017:MPC

[IUV17] L. Iapichino, S. Ulbrich, and S. Volkwein. Multiobjective PDE-constrained optimization using the reduced-basis method. Advances in Computational Mathematics, 43 (5):945–972, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9512-x.

Janik:1998:PIV

[Jan98] Tadeusz J. Janik. Parallel implementation of the p-version of the finite element method for elliptic equations on a sharedmemory architecture. Advances in Computational Mathematics, 8(1–2):97–110, January 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018936011745.

Jeon:1998:NBE

[Jeo98] Youngmok Jeon. New boundary element formulas for the biharmonic equation. Advances in Computational Mathematics, 9(1–2):97–115, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018985322648. Nu-

merical treatment of boundary integral equations.

Juttler:2002:LSF

[JF02] Bert Jüttler and Alf Felis.

Least-squares fitting of algebraic spline surfaces. Advances in Computational Mathematics, 17(1–2):135–152, July 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1015200504295. Ad-

vances in geometrical algorithms and representations.

Jiang:2015:ESE

[JGW15] Shidong Jiang, Leslie Greengard, and Shaobo Wang. Efficient sum-of-exponentials approximations for the heat kernel and their applications. Advances in Computational Mathematics, 41(3):529–551, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9372-1.

Jia:1995:SSS

[Jia95] Rong-Qing Jia. Subdivision schemes in L_{p} spaces. Advances in Computational Mathematics, 3(4):309–341, May 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF03028366.

Jiang:2000:PCO

[Jia00] Qingtang Jiang. Parameterization of m-channel orthogonal multifilter banks. Advances in Computational Mathematics, 12(2–3):189–211, February 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018965102118.

Jiang:2003:PMT

[Jia03] Qingtang Jiang. Parameterizations of masks for tight affine frames with two symmetric/ antisymmetric generators. Advances in Computational Mathematics, 18(2–4):247–268, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1021339707805. Frames.

Jia:2009:SWI

[Jia09a] Rong-Qing Jia. Spline wavelets on the interval with homogeneous boundary conditions. Advances in Computational Mathematics, 30(2):177–200, February 2009. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9064-9.

Jiang:2009:HTF

[Jia09b] Qingtang Jiang. Hexagonal tight frame filter banks with idealized high-pass filters. Advances in Computational Mathematics, 31(1–3):215–236, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9085-4.

Jiang:2011:BWF

[Jia11] Qingtang Jiang. Biorthogonal wavelets with 4-fold axial symmetry for quadrilateral surface multiresolution processing. Advances in Computational Mathematics, 34(2):127–165, February 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9144-5.

Jang:2013:CSM

[JJK13] Sumi Jang, Byeongseon Jeong, and Hong Oh Kim. Compactly supported multiwindow dual Gabor frames of rational sampling density. Advances in Computational Mathematics, 38(1):159–186, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9234-z.

Jenkins:2014:PCG

[JJLR14] Eleanor W. Jenkins, Volker John, Alexander Linke, and Leo G. Rebholz. On the parameter choice in grad–div stabilization for the Stokes equations. Advances in Computational Mathematics, 40(2):491–516, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9316-1.

John:2008:FEE

[JK08] Volker John and Songul Kaya. Finite element error analysis of a variational multiscale method for the Navier–Stokes equations. Advances in Computational Mathematics, 28(1):43–61, January 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-005-9010-z.

Jaklic:2012:AGI

[JKK^{+}12] Gašper Jaklič, Jernej Kozak, Marjeta Krajnc, Vito Vitrih, and Emil Žagar. An approach to geometric interpolation by Pythagorean-hodograph curves. Advances in Computational Mathematics, 37(1):123–150, July 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9209-0.

Jeong:2013:EPR

[JKLY13] Byeongseon Jeong, Hong Oh Kim, Yeon Ju Lee, and Jungho Yoon. Exponential polynomial reproducing property of non-stationary symmetric subdivision schemes and normalized exponential B-splines. Advances in Computational Mathematics, 38(3):647–666, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9253-9.

Jiang:1997:SPM

[JL97] Qingtang Jiang and S. L.

Lee. Spectral properties of matrix continuous refinement operators. Advances in Computational Mathematics, 7(3):383–399, May 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018911306293.

Jia:2006:WBH

[JL06] Rong-Qing Jia and Song-Tao Liu. Wavelet bases of Hermite cubic splines on the interval. Advances in Computational Mathematics, 25(1–3):23–39, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-7609-5.

Jia:2008:AER

[JLZ08] Shanghui Jia, Deli Li, and Shuhua Zhang. Asymptotic expansions and Richardson extrapolation of approximate solutions for integro-differential equations by mixed finite element methods. Advances in Computational Mathematics, 29 (4):337–356, November 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9052-5.

Jodar:2000:SOA

[JM00] Lucas Jódar and Marco Marletta. Solving ODEs arising from non-selfadjoint Hamiltonian eigenproblems. Advances in Computational Mathematics, 13(3):231–256, July 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018954227133.

Jeon:2003:NBE

[JM03] Youngmok Jeon and William McLean. A new boundary element method for the biharmonic equation with Dirichlet boundary conditions. Advances in Computational Mathematics, 19(4):339–354, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1024206232212.

Jackiewicz:2000:CRK

[JMO00] Z. Jackiewicz, A. Marthinsen, and B. Owren. Construction of Runge–Kutta methods of Crouch–Grossman type of high order. Advances in Computational Mathematics, 13 (4):405–415, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016645730465.

Jung:2018:FSS

[JN18] Chang-Yeol Jung and Thien Binh Nguyen. Fine structures for the solutions of the two-dimensional Riemann problems by highorder WENO schemes. Advances in Computational Mathematics, 44(1):147–174, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9538-8.

Johnson:2013:RIS

[Joh13] Michael J. Johnson. The role of inner summaries in the fast evaluation of thin-plate splines. Advances in Computational Mathematics, 39(1):1–25, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9265-5.

Jones:2017:CUP

[Jon17] Robert Stephen Jones. Computing ultra-precise eigenvalues of the Laplacian within polygons. Advances in Computational Mathematics, 43(6): 1325–1354, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9527-y.

Junior:2015:OIF

[JRS15] D. S. Almeida Júnior, A. J. A. Ramos, and M. L. Santos. Observability inequality for the finite-difference semidiscretization of the 1-d coupled wave equations. Advances in Computational Mathematics, 41(1):105–130, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9351-6.

Jetter:1995:GBS

[JS95] K. Jetter and J. Stöckler. A generalization of de Boor’s stability result and symmetric preconditioning. Advances in Computational Mathematics, 3(4):353–367, May 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF03028368.

Ji:1999:CSB

[JS99] Hui Ji and Zuowei Shen. Compactly supported (bi)orthogonal wavelets generated by interpolatory refinable functions. Advances in Computational Mathematics, 11(1):81–104, July 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018999220348.

Jeon:2005:ACB

[JS05] Youngmok Jeon and Dongwoo Sheen. Analysis of a cell boundary element method. Advances in Computational Mathematics, 22(3):201–222, April 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-7671-z.

Jeon:1997:DQM

[JSSE97] Y. Jeon, I. H. Sloan, E. P.

Stephan, and J. Elschner. Discrete qualocation methods for logarithmic-kernel integral equations on a piecewise smooth boundary. Advances in Computational Mathematics, 7(4):547–571, August 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018967424040.

Juttler:1998:DBF

[Jüt98] Bert Jüttler. The dual basis functions for the Bernstein polynomials. Advances in Computational Mathematics, 8(4):345–352, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018912801267.

Jones:1995:ZFR

[JW95] John W. Jones and Bruno D.

Welfert. Zero-free regions for a rational function with applications. Advances in Computational Mathematics, 3(3):265–289, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02988626.

Jetter:2004:PRS

[JZ04] K. Jetter and G. Zimmermann. Polynomial reproduction in subdivision. Advances in Computational Mathematics, 20(1–3):67–86, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1025859224071. Ap-

proximation and applications.

Jia:2010:FAT

[JZ10] Rong-Qing Jia and Hanqing Zhao. A fast algorithm for the total variation model of image denoising. Advances in Computational Mathematics, 33(2):231–241, August 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9128-5.

Konstantinov:1997:SAD

[KA97] Mihail Konstantinov and Vera Angelova. Sensitivity analysis of the differential matrix Riccati equation based on the associated linear differential system. Advances in Computational Mathematics, 7(3):295–301, May 1997.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018903104476.

Kosinka:2015:CWC

[KB15] Jiří Kosinka and Michael Bartoň. Convergence of Wachspress coordinates: from polygons to curved domains. Advances in Computational Mathematics, 41(3):489–505, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9370-3.

Kublik:2016:LAT

[KC16] Richard A. Kublik and David L.

Chopp. A locally adaptive time stepping algorithm for the solution to reaction diffusion equations on branched structures. Advances in Computational Mathematics, 42(3):621–649, June 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9437-9.

Kordy:2017:NSC

[KCW17] Michal Kordy, Elena Cherkaev, and Philip Wannamaker. Null space correction and adaptive model order reduction in multi-frequency Maxwell’s problem. Advances in Computational Mathematics, 43 (1):171–193, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9482-z.

Keinert:1995:NSB

[Kei95] Fritz Keinert. Numerical stability of biorthogonal wavelet transforms. Advances in Computational Mathematics, 4 (1–2):1–26, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02123471.

Krivodonova:2003:EED

[KF03] Lilia Krivodonova and Joseph E. Flaherty. Error estimation for discontinuous Galerkin solutions of two-dimensional hyperbolic problems. Advances in Computational Mathematics, 19(1–3):57–71, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1022894504834. Chal-

lenges in computational mathematics (Pohang, 2001).

Keck:2005:PTI

[KH05] R. Keck and D. Hietel. A projection technique for incompressible flow in the meshless finite volume particle method. Advances in Computational Mathematics, 23(1–2):143–169, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1831-7.

Kriva:2016:ACC

[KHM16] Zuzana Krivá, Angela Handlovičová, and Karol Mikula. Adaptive cell-centered finite volume method for diffusion equations on a consistent quadtree grid. Advances in Computational Mathematics, 42(2):249–277, April 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9423-2.

Kim:2006:GDA

[Kim06] M.-Y. Y. Kim. Global dynamics of approximate solutions to an age-structured epidemic model with diffusion. Advances in Computational Mathematics, 25(4):451–474, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7639-7.

Kirkup:1998:FCC

[Kir98]  S. M. Kirkup.  Fortran

codes for computing the discrete Helmholtz integral operators. Advances in Computational Mathematics, 9(3– 4):391–409, November 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018953910353.

Kosinka:2009:HIP

[KJ09] Jiří Kosinka and Bert Jüttler.

C^{1} Hermite interpolation by Pythagorean hodograph quintics in Minkowski space. Advances in Computational Mathematics, 30(2):123–140, February 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9059-y.

Koras:1999:CCP

[KK99] G. D. Koras and P. D. Kaklis. Convexity conditions for parametric tensor-product Bspline surfaces. Advances in Computational Mathematics, 10(3–4):291–309, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018938901692.

Kuchment:2002:DOG

[KK02] P. Kuchment and L. Kunyansky. Differential operators on graphs and photonic crystals. Advances in Computational Mathematics, 16(2–3):263–290, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014481629504. Mod-

eling and computation in optics and electromagnetics.

Karni:2005:LEA

[KK05] Smadar Karni and Alexander Kurganov. Local error analysis for approximate solutions of hyperbolic conservation laws. Advances in Computational Mathematics, 22(1):79–99, January 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7099-8.

Kim:2003:QBF

[KKL03] Hong Oh Kim, Rae Young Kim, and Jae Kun Lim. Quasibiorthogonal frame multiresolution analyses and wavelets. Advances in Computational Mathematics, 18(2–4):269–296, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021371227388. Frames.

Kim:2010:QIR

[KKLY10] Hong Oh Kim, Rae Young Kim, Yeon Ju Lee, and Jungho Yoon. Quasi-interpolatory refinable functions and construction of biorthogonal wavelet systems. Advances in Computational Mathematics, 33 (3):255–283, October 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9129-4.

Kozak:2015:PCP

[KKV15] Jernej Kozak, Marjeta Krajnc, and Vito Vitrih. Parametric curves with Pythagorean binormal. Advances in Computational Mathematics, 41 (4):813–832, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9387-7.

Kim:2000:MMF

[KL00] Yongdeok Kim and Sungyun Lee. Modified Mini finite element for the Stokes problem in R^{2} or R^{3}. Advances in Computational Mathematics, 12(2–3):261–272, February 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018973303935.

Konovalov:2004:FKS

[KL04] V. N. Konovalov and D. Leviatan. Free-knot splines approximation of s-monotone functions. Advances in Computational Mathematics, 20(4):347–366, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1027324000817.

Konovalov:2007:FSA

[KL07] V. N. Konovalov and D. Leviatan. Freeknot splines approximation of Sobolev-type classes of s-monotone functions. Advances in Computational Mathematics, 27(2):211–236, August 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9032-9.

Kruschel:2015:CAR

[KL15] Christian Kruschel and Dirk A.

Lorenz. Computing and analyzing recoverable supports for sparse reconstruction. Advances in Computational Mathematics, 41(6):1555–1580, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9403-6.

Kosmas:2016:AHO

[KL16] Odysseas Kosmas and Sigrid Leyendecker. Analysis of higher order phase fitted variational integrators. Advances in Computational Mathematics, 42(3):605–619, June 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9436-x.

Kopteva:2017:IMN

[KL17] Natalia Kopteva and Torsten Linß. Improved maximumnorm a posteriori error estimates for linear and semilinear parabolic equations. Advances in Computational Mathematics, 43(5):999–1022, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9514-3.

Koch:1995:CCK

[KLNS95] Per Erik Koch, Tom Lyche, Marian Neamtu, and Larry L.

Schumaker. Control curves and knot insertion for trigonometric splines. Advances in Computational Mathematics, 3(4):405–424, May 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF03028369.

Kobarg:2014:NEM

[KMO^{+}14] Jan Hendrik Kobarg, Peter Maass, Janina Oetjen, Oren Tropp, Eyal Hirsch, Chen Sagiv, Mohammad Golbabaee, and Pierre Vandergheynst. Numerical experiments with MALDI Imaging data. Advances in Computational Mathematics, 40(3):667–682, June 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9325-0.

Kropinski:2014:FIE

[KN14] Mary Catherine A. Kropinski and Nilima Nigam. Fast integral equation methods for the Laplace–Beltrami equation on the sphere. Advances in Computational Mathematics, 40(2):577–596, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9319-y.

Knobloch:2009:CSP

[Kno09] Petr Knobloch. On the choice of the SUPG parameter at outflow boundary layers. Advances in Computational Mathematics, 31(4):369–389, November 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9075-6.

Kaneko:1998:SPG

[KNP98] Hideaki Kaneko, Richard D.

Noren, and Peter A. Padilla. Singularity preserving Galerkin method for Hammerstein equations with logarithmic kernel. Advances in Computational Mathematics, 9(3– 4):363–376, November 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018910128100.

Kropinski:2016:IEM

[KNQ16] M. C. Kropinski, N. Nigam, and B. Quaife. Integral equation methods for the Yukawa–Beltrami equation on the sphere. Advances in Computational Mathematics, 42(2):469–488, April 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9431-2.

Krizek:2001:PEE

[KNV01] Michal Křížek, Jan Němec, and Tomáš Vejchodský. A posteriori error estimates for axisymmetric and nonlinear problems. Advances in Computational Mathematics, 15(1– 4):219–236, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1014234911830. A

posteriori error estimation and adaptive computational methods.

Kazeev:2017:QFE

[KORS17] Vladimir Kazeev, Ivan Oseledets, Maxim Rakhuba, and Christoph Schwab. QTT-finite-element approximation for multiscale problems I: model problems in one dimension. Advances in Computational Mathematics, 43(2):411–442, April 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9491-y.

Kostic:2015:GPE

[Kos15]  Vladimir Kostić.  On gen-

eral principles of eigenvalue localizations via diagonal dominance. Advances in Computational Mathematics, 41(1):55– 75, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1007/s10444-
                                        014-9349-0.

Kotnik:2008:PCF

[Kot08] Tadej Kotnik. The primecounting function and its analytic approximations: π(x) and its approximations. Advances in Computational Mathematics, 29(1):55–70, July 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9039-2.

Khoromskij:1995:MPR

[KP95] Boris N. Khoromskij and Siegfried Prössdorf. Multilevel preconditioning on the refined interface and optimal boundary solvers for the Laplace equation. Advances in Computational Mathematics, 4(1– 2):331–355, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02123480.

Khoromskij:1998:FCH

[KP98] Boris N. Khoromskij and Siegfried Prössdorf. Fast computations with the harmonic Poincaré–Steklov operators on nested refined meshes. Advances in Computational Mathematics, 8(1–2):111–135, January 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018988028583.

Kouibia:2004:ASF

[KP04] A. Kouibia and M. Pasadas.

Approximation of surfaces by fairness bicubic splines. Advances in Computational Mathematics, 20(1–3):87–103, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1025805701726. Ap-

proximation and applications.

Krishnan:2009:PDA

[KPY09] D. Krishnan, Quang Vinh Pham, and Andy M. Yip. A primal-dual active-set algorithm for bilaterally constrained total variation deblurring and piecewise constant Mumford–Shah segmentation problems. Advances in Computational Mathematics, 31(1–3):237–266, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9101-8.

Krasauskas:2002:TSP

[Kra02] Rimvydas Krasauskas. Toric surface patches. Advances in Computational Mathematics, 17(1–2):89–113, July 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1015289823859. Ad-

vances in geometrical algorithms and representations.

Kopteva:2017:ANS

[KS17] Natalia Kopteva and Martin Stynes. Analysis and numerical solution of a Riemann–Liouville fractional derivative two-point boundary value problem. Advances in Computational Mathematics, 43(1):77–99, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9476-x.

Kunoth:1995:MPA

[Kun95]  Angela Kunoth.  Multilevel

preconditioning — appending boundary conditions by Lagrange multipliers. Advances in Computational Mathematics, 4(1–2):145–170, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02123477.

Kunert:2001:RPE

[Kun01] Gerd Kunert. Robust a posteriori error estimation for a singularly perturbed reaction– diffusion equation on anisotropic tetrahedral meshes. Advances in Computational Mathematics, 15(1–4):237–259, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014248711347. A

posteriori error estimation and adaptive computational methods.

Kunis:2009:NSR

[Kun09]  Stefan Kunis.  A note on

stability results for scattered data interpolation on Euclidean spheres. Advances in Computational Mathematics, 30(4):303–314, May 2009.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9069-4.

Kvasov:2014:MCI

[Kva14] Boris I. Kvasov. Monotone and convex interpolation by weighted quadratic splines. Advances in Computational Mathematics, 40(1):91–116, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9300-9.

Kuijt:2001:LAS

[KvD01] Frans Kuijt and Ruud van Damme. A linear approach to shape preserving spline approximation. Advances in Computational Mathematics, 14(1):25–48, January 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1016660513725.

Koren:1995:APM

[KvL95] B. Koren and B. van Leer. Analysis of preconditioning and multigrid for Euler flows with low-subsonic regions. Advances in Computational Mathematics, 4(1–2):127–144, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02123476.

Kim:2004:UCM

[KXZ04] Hwanho Kim, Jinchao Xu, and Ludmil Zikatanov. Uniformly convergent multigrid methods for convection–diffusion problems without any constraint on coarse grids. Advances in Computational Mathematics, 20(4):385–399, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1027378015262.

Kang:2001:SPE

[KY01] Tong Kang and Dehao Yu.

Some a posteriori error estimates of the finite-difference streamline–diffusion method for convection-dominated diffusion equations. Advances in Computational Mathematics, 15(1– 4):193–218, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014294002686. A

posteriori error estimation and adaptive computational methods.

Lai:2006:CMC

[Lai06] Ming-Jun Lai. Construction of multivariate compactly supported orthonormal wavelets. Advances in Computational Mathematics, 25(1–3):41–56, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7643-y.

Lasserre:2016:LDA

[Las16] Jean B. Lasserre. Lebesgue decomposition in action via semidefinite relaxations. Advances in Computational Mathematics, 42(5):1129–1148, October 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9456-1.

Liu:1993:EBF

[LB93] W. B. Liu and John W. Barrett. Error bounds for the finite element approximation of a degenerate quasilinear parabolic variational inequality. Advances in Computational Mathematics, 1(2):223–239, June 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02071387.

Li:2009:NSF

[LC09] Jian Li and Zhangxin Chen. A new stabilized finite volume method for the stationary Stokes equations. Advances in Computational Mathematics, 30(2):141–152, February 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9060-5.

Li:2013:SDS

[LC13] Jian Li and Zhangxin Chen.

On the semi-discrete stabilized finite volume method for the transient Navier–Stokes equations. Advances in Computational Mathematics, 38 (2):281–320, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9237-9.

Lipovetsky:2007:EAC

[LD07] Evgeny Lipovetsky and Nira Dyn. An efficient algorithm for the computation of the metric average of two intersecting convex polygons, with application to morphing. Advances in Computational Mathematics, 26(1–3):269–282, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7473-6.

Li:2016:HOS

[LD16] Can Li and Weihua Deng. High order schemes for the tempered fractional diffusion equations. Advances in Computational Mathematics, 42(3):543–572, June 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9434-z.

LeMehaute:2000:SAM

[Le 00] A. Le Méhauté. On some aspects of multivariate polynomial interpolation. Advances in Computational Mathematics, 12(4):311–333, March 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018985606661. Mul-

tivariate polynomial interpolation.

LeGia:2005:APP

[Le 05] Q. T. Le Gia. Approximation of parabolic PDEs on spheres using spherical basis functions. Advances in Computational Mathematics, 22(4):377–397, May 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-3960-9.

Ledford:2015:CRF

[Led15] Jeff Ledford. On the convergence of regular families of cardinal interpolators. Advances in Computational Mathematics, 41(2):357–371, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9361-4.

Lee:2003:ACM

[Lee03]  Hyung-Chun Lee.  Analy-

sis and computational methods of Dirichlet boundary optimal control problems for 2D Boussinesq equations. Advances in Computational Mathematics, 19(1–3):255–275, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1022872602498. Chal-

lenges in computational mathematics (Pohang, 2001).

Lee:2016:TUA

[Lee16] Jeonghun J. Lee. Towards a unified analysis of mixed methods for elasticity with weakly symmetric stress. Advances in Computational Mathematics, 42(2):361–376, April 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9427-y.

Lemvig:2009:CPD

[Lem09]  Jakob Lemvig.  Construct-

ing pairs of dual bandlimited framelets with desired time localization. Advances in Computational Mathematics, 30(3):231–247, April 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9066-7.

Lenze:1994:LGS

[Len94] Burkhard Lenze. On local and global sigma–pi neural networks. A common link. Advances in Computational Mathematics, 2(4):479–491, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02521610.

Lenze:1996:HSP

[Len96] Burkhard Lenze. Hyperbolic sigma–pi neural network operators for compactly supported continuous functions. Advances in Computational Mathematics, 5(1):163–172, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124741.

Leopoldseder:2002:CSS

[Leo02]  Stefan Leopoldseder.  Cone

spline surfaces and spatial arc splines — a sphere geometric approach. Advances in Computational Mathematics, 17(1–2):49–66, July 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1015215802130. Ad-

vances in geometrical algorithms and representations.

Leopardi:2013:DSR

[Leo13]  Paul Leopardi.  Discrep-

ancy, separation and Riesz energy of finite point sets on the unit sphere. Advances in Computational Mathematics, 39(1):27–43, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9266-4.

Levin:1999:ULP

[Lev99] David Levin. Using Laurent polynomial representation for the analysis of non-uniform binary subdivision schemes. Advances in Computational Mathematics, 11(1):41–54, July 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018907522165.

Li:2013:CCN

[LG13] J. Li and O. Gonzalez. Convergence and conditioning of a Nyström method for Stokes flow in exterior three-dimensional domains. Advances in Computational Mathematics, 39(1): 143–174, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
012-9272-1.

Leng:2013:OPD

[LH13] Jinsong Leng and Deguang Han.

Orthogonal projection decomposition of matrices and construction of fusion frames. Advances in Computational Mathematics, 38(2):369–381, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9241-0.

Liu:2008:NOD

[LHY08] Yang Liu, Qiya Hu, and Dehao Yu. A non-overlapping domain decomposition for low-frequency time-harmonic Maxwell’s equations in unbounded domains. Advances in Computational Mathematics, 28(4):355–382, May 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9027-6.

Li:1994:HOB

[Li94]  Xin Li.  Hankel operators

and best Hankel approximation on the half-plane. Advances in Computational Mathematics, 2(3):343–355, April 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02521115.

Li:2003:IMC

[Li03]  Yuesheng Li.  The inver-

sion of multiscale convolution approximation and average of distributions. Advances in Computational Mathematics, 19(1–3):293–306, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1022811005225. Chal-

lenges in computational mathematics (Pohang, 2001).

Li:2004:CSM

[Li04]  Song Li.  Characterization

of smoothness of multivariate refinable functions and convergence of cascade algorithms of nonhomogeneous refinement equations. Advances in Computational Mathematics, 20(4):311–331, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1027373528712.

Li:2005:MMF

[Li05a] Jichun Li. Mixed methods for fourth-order elliptic and parabolic problems using radial basis functions. Advances in Computational Mathematics, 23(1–2):21–30, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1807-7.

Li:2005:CMF

[Li05b]  Xin Li.  On convergence

of the method of fundamental solutions for solving the Dirichlet problem of Poisson’s equation. Advances in Computational Mathematics, 23 (3):265–277, October 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1782-z.

Li:2008:CMF

[Li08a] Xin Li. Convergence of the method of fundamental solutions for Poisson’s equation on the unit sphere. Advances in Computational Mathematics, 28(3):269–282, April 2008.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-006-9022-3.

Li:2008:RCM

[Li08b] Xin Li. Rate of convergence of the method of fundamental solutions and hyperinterpolation for modified Helmholtz equations on the unit ball. Advances in Computational Mathematics, 29(4):393–413, November 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9056-1.

Li:2009:API

[Li09] Xin Li. Approximation of potential integral by radial bases for solutions of Helmholtz equation. Advances in Computational Mathematics, 30(3):201–230, April 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9065-8.

Li:2010:RBAa

[Li10a] Xin Li. Radial basis approximation and its application to biharmonic equation. Advances in Computational Mathematics, 32(3):275–302, April 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9107-2.

Li:2010:RBAb

[Li10b] Xin Li. Radial basis approximation for Newtonian potentials. Advances in Computational Mathematics, 33(1):1–24, July 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9117-8.

Liu:1999:RKC

[Liu99] Yunkang Liu. Runge–Kuttacollocation methods for systems of functional-differential and functional equations. Advances in Computational Mathematics, 11(4):315–329, December 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018936328525.

Ling:2005:LSP

[LK05] Leevan Ling and Edward J.

Kansa. A least-squares preconditioner for radial basis functions collocation methods. Advances in Computational Mathematics, 23(1–2):31–54, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1809-5.

Lewanowicz:2017:BFD

[LKW17] Stanisław Lewanowicz, Paweł Keller, and Paweł Woźny. Bézier form of dual bivariate Bernstein polynomials. Advances in Computational Mathematics, 43(4):777–793, August 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9506-8.

Laghchim-Lahlou:1998:FSC

[LL98] Mohammed Laghchim-Lahlou. The C^{r}-fundamental splines of Clough–Tocher and Powell– Sabin types for Lagrange interpolation on a three direction mesh. Advances in Computational Mathematics, 8(4):353–366, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018904532218.

Levesley:1999:DFS

[LL99] Jeremy Levesley and Will Light. Direct form seminorms arising in the theory of interpolation by translates of a basis function. Advances in Computational Mathematics, 11(2– 3):161–182, November 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018971808961. Ra-

dial basis functions and their applications.

Lai:2004:NKT

[LL04] Ming-Jun Lai and A. LeMéhauté. A new kind of trivariate C^{1} macro-element. Advances in Computational Mathematics, 21 (3–4):273–292, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000032047.
05052.27.

Lin:2007:CAE

[LL07a] Qun Lin and Jiaquan Liu.

Counterexamples to the asymptotic expansion of interpolation in finite elements. Advances in Computational Mathematics, 27(2):167–177, August 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9030-y.

Lu:2007:FSA

[LL07b] Shujuan Lü and Qishao Lu.

Fourier spectral approximation to long-time behavior of three dimensional Ginzburg– Landau type equation. Advances in Computational Mathematics, 27(3):293–318, October 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-9004-x.

Lv:2010:EEF

[LL10] Junliang Lv and Yonghai Li.

L^{2} error estimate of the finite volume element methods on quadrilateral meshes. Advances in Computational Mathematics, 33(2):129–148, August 2010. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9121-z.

Lian:2011:GFS

[LL11] Qiao-Fang Lian and Yun-Zhang Li. Gabor frame sets for subspaces. Advances in Computational Mathematics, 34(4):391–411, May 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-010-9161-4.

Lin:2012:AHO

[LL12a] Fu-Rong Lin and Fen Liang. Application of high order numerical quadratures to numerical inversion of the Laplace transform. Advances in Computational Mathematics, 36 (2):267–278, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9202-7.

Lv:2012:EES

[LL12b] Junliang Lv and Yonghai Li. L^{2} error estimates and superconvergence of the finite volume element methods on quadrilateral meshes. Advances in Computational Mathematics, 37(3):393–416, October 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9215-2.

Li:2013:SGF

[LL13] Yun-Zhang Li and Qiao-Fang Lian. Super Gabor frames on discrete periodic sets. Advances in Computational Mathematics, 38(4):763–799, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9259-3.

Liu:2016:WGF

[LLC16] Xin Liu, Jian Li, and Zhangxin Chen. A weak Galerkin finite element method for the Oseen equations. Advances in Computational Mathematics, 42(6):1473–1490, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9471-2.

Laghchim-Lahlou:1994:TFE

[LLS94] M. Laghchim-Lahlou and P. Sablonnière.

Triangular finite elements of HCT type and class C^{ρ}. Advances in Computational Mathematics, 2(1):101–122, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02519038.

Lawton:1995:CCS

[LLS95] Wayne Lawton, S. L. Lee, and Zuowei Shen. Characterization of compactly supported refinable splines. Advances in Computational Mathematics, 3(1–2):137–145, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF03028364.

Laghchim-Lahlou:1996:FEP

[LLS96] M. Laghchim-Lahlou and P. Sablonnière. C^{r}-finite elements of Powell– Sabin type on the three direction mesh. Advances in Computational Mathematics, 6 (1):191–206, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127703.

Lee:2003:LFN

[LLS03] Chang-Ock Lee, Jongwoo Lee, and Dongwoo Sheen. A locking-free nonconforming finite element method for planar linear elasticity. Advances in Computational Mathematics, 19(1–3):277–291, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1022838628615. Chal-

lenges in computational mathematics (Pohang, 2001).

Lee:2009:CNA

[LLS09] Cheng-Feng Lee, Leevan Ling, and Robert Schaback. On convergent numerical algorithms for unsymmetric collocation. Advances in Computational Mathematics, 30(4):339–354, May 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9071-x.

Lee:2006:SBS

[LLY06] Byung-Gook Lee, Yeon Ju Lee, and Jungho Yoon. Stationary binary subdivision schemes using radial basis function interpolation. Advances in Computational Mathematics, 25(1–3):57–72, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7642-z.

Luo:2012:PPC

[LLY12] Xingjun Luo, Fanchun Li, and Suhua Yang. A posteriori parameter choice strategy for fast multiscale methods solving illposed integral equations. Advances in Computational Mathematics, 36(2):299–314, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9229-9.

Li:2010:ESS

[LLZ10] Mingxia Li, Qun Lin, and Shuhua Zhang. Extrapolation and superconvergence of the Steklov eigenvalue problem. Advances in Computational Mathematics, 33(1):25–44, July 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9118-7.

Lamberti:2004:TQI

[LM04] P. Lamberti and C. Manni. Tensioned quasi-interpolation via geometric continuity. Advances in Computational Mathematics, 20(1–3):105–127, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1025823221346. Ap-

proximation and applications.

Lyche:2006:TPE

[LM06] Tom Lyche and Marie-Laurence Mazure. Total positivity and the existence of piecewise exponential B-splines. Advances in Computational Mathematics, 25(1–3):105–133, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7633-0.

LeGia:2014:SHE

[LM14] Quoc Thong Le Gia and William McLean. Solving the heat equation on the unit sphere via Laplace transforms and radial basis functions. Advances in Computational Mathematics, 40(2):353–375, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9311-6.

Lemvig:2014:PTF

[LMO14] Jakob Lemvig, Christopher Miller, and Kasso A. Okoudjou. Prime tight frames. Advances in Computational Mathematics, 40(2):315–334, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9309-0.

Liu:2001:MEE

[LMT01] Wenbin Liu, Heping Ma, and Tao Tang. On mixed error estimates for elliptic obstacle problems. Advances in Computational Mathematics, 15(1–4):261–283, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014261013164. A

posteriori error estimation and adaptive computational methods.

Linh:2011:MEA

[LMV11] Vu Hoang Linh, Volker Mehrmann, and Erik S. Van Vleck. QR methods and error analysis for computing Lyapunov and Sacker–Sell spectral intervals for linear differential-algebraic equations. Advances in Computational Mathematics, 35(2– 4):281–322, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9156-1.

Liu:2013:FRK

[LMWX13] Jianqiang Liu, Charles A. Micchelli, Rui Wang, and Yuesheng Xu. Finite rank kernels for multi-task learning. Advances in Computational Mathematics, 38(2):427–439, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9244-x.

Lechleiter:2014:TGM

[LN14] Armin Lechleiter and Dinh-Liem Nguyen. A trigonometric Galerkin method for volume integral equations arising in TM grating scattering. Advances in Computational Mathematics, 40(1):1–25, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9295-2.

Lyche:2002:PIM

[LNW02] Tom Lyche, Trygve K. Nilssen, and Ragnar Winther. Preconditioned iterative methods for scattered data interpolation. Advances in Computational Mathematics, 17 (3):237–256, October 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1016050415412.

Li:2008:ONS

[LO08] Shidong Li and Hidemitsu Ogawa. Optimal noise suppression: a geometric nature of pseudoframes for subspaces. Advances in Computational Mathematics, 28(2):141–155, February 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-006-9014-3.

Lopez:2018:CEB

[Lóp18] José L. López. Convergent expansions of the Bessel functions in terms of elementary functions. Advances in Computational Mathematics, 44 (1):277–294, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9543-y.

Lowitzsch:2005:MVR

[Low05]  Svenja Lowitzsch.  Matrix-

valued radial basis functions: stability estimates and applications. Advances in Computational Mathematics, 23 (3):299–315, October 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1786-8.

Lyche:2004:OSM

[LP04] T. Lyche and J. M. Peña.

Optimally stable multivariate bases. Advances in Computational Mathematics, 20 (1–3):149–159, January 2004.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1025863309959. Ap-

proximation and applications.

Li:2010:SSP

[LP10] Song Li and Yali Pan. Subdivision schemes with polynomially decaying masks. Advances in Computational Mathematics, 32(4):487–507, May 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9116-9.

Lyon:2014:FAS

[LP14] M. Lyon and J. Picard. The Fourier approximation of smooth but non-periodic functions from unevenly spaced data. Advances in Computational Mathematics, 40(5– 6):1073–1092, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9342-7.

Lorente-Pardo:2000:CSA

[LPSSP00] J. Lorente-Pardo, P. Sablonnière, and M. C. Serrano-Pérez. On the convexity of C^{1} surfaces associated with some quadrilateral finite elements. Advances in Computational Mathematics, 13(3):271–292, July 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018958328042.

Lammers:2010:ADF

[LPY10] Mark Lammers, Alexander M.

Powell, and Özgür Yılmaz. Alternative dual frames for digitalto-analog conversion in sigma– delta quantization. Advances in Computational Mathematics, 32(1):73–102, January 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9088-1.

Li:2014:HWS

[LQ14] Gang Li and Jianxian Qiu.

Hybrid WENO schemes with different indicators on curvilinear grids. Advances in Computational Mathematics, 40(4):747–772, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9322-3.

Levesley:2007:RBI

[LR07] J. Levesley and D. L. Ragozin.

Radial basis interpolation on homogeneous manifolds: convergence rates. Advances in Computational Mathematics, 27(2):237–246, August 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-9000-1.

Linss:2012:NMS

[LRS12] Torsten Linß, Hans-Görg Roos, and Martin Schopf. Nitschemortaring for singularly perturbed convection–diffusion problems. Advances in Computational Mathematics, 36(4): 581–603, May 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1007/s10444-
                                        011-9195-2.

Lai:1998:APB

[LS98] Ming-Jun Lai and Larry L. Schumaker. On the approximation power of bivariate splines. Advances in Computational Mathematics, 9(3– 4):251–279, November 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018958011262.

Lu:2005:PDD

[LS05] Fangyan Lu and Hongwei Sun. Positive definite dot product kernels in learning theory. Advances in Computational Mathematics, 22(2):181–198, February 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-3140-6.

Levie:2014:ATA

[LSLS14] R. Levie, H.-G. Stark, F. Lieb, and N. Sochen. Adjoint translation, adjoint observable and uncertainty principles. Advances in Computational Mathematics, 40(3):609–627, June 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9336-x.

LeGia:2011:SNP

[LST11] Q. T. Le Gia, E. P. Stephan, and T. Tran. Solution to the Neumann problem exterior to a prolate spheroid by radial basis functions. Advances in Computational Mathematics, 34(1):83–103, January 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9145-4.

LeGia:2017:ZGL

[LSW17] Q. T. Le Gia, I. H. Sloan, and H. Wendland. Zooming from global to local: a multiscale RBF approach. Advances in Computational Mathematics, 43(3):581–606, June 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9498-4.

Li:2015:MSF

[LSXZ15] Qia Li, Lixin Shen, Yuesheng Xu, and Na Zhang. Multistep fixed-point proximity algorithms for solving a class of optimization problems arising from image processing. Advances in Computational Mathematics, 41(2):387–422, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9363-2.

Liandrat:1998:FAS

[LT98] J. Liandrat and Ph. Tchamitchian. On the fast approximation of some nonlinear operators in nonregular wavelet spaces. Advances in Computational Mathematics, 8(4):179–192, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018992129492.

Leviatan:2006:SAG

[LT06] D. Leviatan and V. N. Temlyakov. Simultaneous approximation by greedy algorithms. Advances in Computational Mathematics, 25(1–3):73–90, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7613-4.

Liu:2012:SGT

[LT12] Entao Liu and Vladimir N. Temlyakov. Super greedy type algorithms. Advances in Computational Mathematics, 37(4):493–504, November 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9220-5.

Lopez:2013:NSE

[LT13] José L. López and Nico M.

Temme. New series expansions of the Gauss hypergeometric function. Advances in Computational Mathematics, 39(2):349–365, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-012-9283-y.

Luo:1999:SPD

[Luo99]  Zuhua Luo.  Strictly posi-

tive definiteness of Hermite interpolation on spheres. Advances in Computational Mathematics, 10(3–4):261–270, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018982716713.

Luther:2004:ASN

[Lut04]  U. Luther.  Approximation

spaces in the numerical analysis of operator equations. Advances in Computational Mathematics, 20(1–3):129–147, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1025806612592. Ap-

proximation and applications.

Liu:2003:UNU

[LW03] Youming Liu and Yang Wang.

The uniformity of non-uniform Gabor bases. Advances in Computational Mathematics, 18 (2–4):345–355, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021350103925. Frames.

Lai:2004:SMS

[LW04] Ming-Jun Lai and Paul Wenston. L_{1} spline methods for scattered data interpolation and approximation. Advances in Computational Mathematics, 21 (3–4):293–315, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000032042.
35918.c2.

LeGia:2014:DCS

[LW14] Q. T. Le Gia and H. Wendland. Data compression on the sphere using multiscale radial basis function approximation. Advances in Computational Mathematics, 40(4):923–943, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9334-z.

Liu:2017:FDS

[LWL17] Fu-Jun Liu, Zhong-Qing Wang, and Hui-Yuan Li. A fully diagonalized spectral method using generalized Laguerre functions on the half line. Advances in Computational Mathematics, 43(6):1227–1259, December 2017. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9522-3.

Lian:2006:CLT

[LWY06] Qiaofang Lian, Yongge Wang, and Dunyan Yan. A characterization of L^{p}(R) by local trigonometric bases with 1 < p < ∞. Advances in Computational Mathematics, 25(1–3):91–104, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7625-0.

Wang:2006:SLP

[lWyG06] Li lian Wang and Ben yu Guo.

Stair Laguerre pseudospectral method for differential equations on the half line. Advances in Computational Mathematics, 25(1–3):305–322, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-7608-6.

Xu:2002:MLL

[lXyG02] Cheng long Xu and Ben yu Guo.

Mixed Laguerre–Legendre spectral method for incompressible flow in an infinite strip. Advances in Computational Mathematics, 16(1):77–96, January 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014249613222.

Liu:2001:PEE

[LY01] Wenbin Liu and Ningning Yan. A posteriori error estimates for distributed convex optimal control problems. Advances in Computational Mathematics, 15(1–4):285–309, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1014239012739. A

posteriori error estimation and adaptive computational methods.

Li:2007:CMK

[LY07] Haitao Li and Dunyan Yan. Characterizations of multi-knot piecewise linear spectral sequences. Advances in Computational Mathematics, 27 (4):401–422, November 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-005-9006-8.

Liu:2008:SAN

[LY08] Huipo Liu and Ningning Yan. Superconvergence analysis of the nonconforming quadrilateral linear-constant scheme for Stokes equations. Advances in Computational Mathematics, 29(4):375–392, November 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9054-3.

Li:2013:BMS

[LYY13] Youfa Li, Shouzhi Yang, and Dehui Yuan. Bessel multiwavelet sequences and dual multiframelets in Sobolev spaces. Advances in Computational Mathematics, 38(3):491–529, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9246-8.

Zhu:2003:ASS

[lZmCRX03] You lan Zhu, Bin mu Chen, Hongliang Ren, and Hanping Xu. Application of the singularity-separating method to American exotic option pricing. Advances in Computational Mathematics, 19(1–3):147–158, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1022835722199. Chal-

lenges in computational mathematics (Pohang, 2001).

Lin:1998:AMI

[LZY98] Qun Lin, Shuhua Zhang, and Ningning Yan. An acceleration method for integral equations by using interpolation post-processing. Advances in Computational Mathematics, 9 (1–2):117–129, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018925103993. Nu-

merical treatment of boundary integral equations.

Lin:2009:SSP

[LZZ09] Yanping Lin, Kai Zhang, and Jun Zou. Studies on some perfectly matched layers for one-dimensional timedependent systems. Advances in Computational Mathematics, 30(1):1–35, January 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9055-2.

MacLeod:1994:TBT

[Mac94]  Allan J. MacLeod.  Table-

based tests for Bessel function software. Advances in Computational Mathematics, 2(2):251–260, March 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02521111.

Maiorov:2005:LBR

[Mai05] V. Maiorov. On lower bounds in radial basis approximation. Advances in Computational Mathematics, 22(2):103–113, February 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1090-7.

Maiorov:2006:PDE

[Mai06] V. Maiorov. Pseudo-dimension and entropy of manifolds formed by affine-invariant dictionary. Advances in Computational Mathematics, 25 (4):435–450, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7645-9.

Maiorov:2010:GPR

[Mai10]  Vitaly Maiorov.  Geometric

properties of the ridge function manifold. Advances in Computational Mathematics, 32 (2):239–253, February 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9106-3;
 http://link.springer.com/
content/pdf/10.1007/s10444-
008-9106-3.pdf.

Marletta:1994:NSE

[Mar94] Marco Marletta. Numerical solution of eigenvalue problems for Hamiltonian systems. Advances in Computational Mathematics, 2(2):155–184, March 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02521106.

Maroni:1995:ISD

[Mar95] P. Maroni. An introduction to second degree forms. Advances in Computational Mathematics, 3(1–2):59–88, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02431996.

Matthies:2009:LPS

[Mat09] Gunar Matthies. Local projection stabilisation for higher order discretisations of convection– diffusion problems on Shishkin meshes. Advances in Computational Mathematics, 30(4): 315–337, May 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1007/s10444-
                                        008-9070-y.

Mazure:1998:VTD

[Maz98] Marie-Laurence Mazure. Vandermonde type determinants and blossoming. Advances in Computational Mathematics, 8(4):291–315, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018956616288.

Mazure:1999:CSP

[Maz99] Marie-Laurence Mazure. Chebyshev spaces with polynomial blossoms. Advances in Computational Mathematics, 10(3–4):219–238, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018995019439.

Mazure:2001:CSB

[Maz01] Marie-Laurence Mazure. Chebyshev splines beyond total positivity. Advances in Computational Mathematics, 14 (2):129–156, February 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016616731472.

Mazure:2004:BOB

[Maz04] Marie-Laurence Mazure. Blossoms and optimal bases. Advances in Computational Mathematics, 20(1–3):177–203, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1025855123163. Ap-

proximation and applications.

Mazure:2007:PBA

[Maz07] Marie-Laurence Mazure. Pseudoaffinity, de Boor algorithm, and blossoms. Advances in Computational Mathematics, 26 (1–3):305–322, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7450-0.

Morton:1996:SBC

[MB96] K. W. Morton and N. A.

Burgess. The stability of boundary conditions for an angledderivative difference scheme. Advances in Computational Mathematics, 6(1):263–279, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/BF02127707.

John Crank 80th birthday special issue.

Mao:2008:QNF

[MC08] Shipeng Mao and Shaochun Chen. A quadrilateral nonconforming finite element for linear elasticity problem. Advances in Computational Mathematics, 28(1):81–100, January 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-006-9017-0.

Milovanovic:2010:NGQ

[MC10] Gradimir V. Milovanović and Aleksandar S. Cvetković. Nonstandard Gaussian quadrature formulae based on operator values. Advances in Computational Mathematics, 32(4):431–486, May 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9114-y.

Massar:2011:FCS

[MFB^{+}11] Melody L. Massar, Matthew Fickus, Erik Bryan, Douglas T. Petkie, and Andrew J. Terzuoli, Jr. Fast computation of spectral centroids. Advances in Computational Mathematics, 35(1):83–97, July 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9167-y.

Mustafa:2017:FUB

[MH17] Ghulam Mustafa and Rabia Hameed. Families of univariate and bivariate subdivision schemes originated from quartic B-spline. Advances in Computational Mathematics, 43(5):1131–1161, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9519-y.

Mhaskar:1993:APM

[Mha93]  H. N. Mhaskar.  Approxi-

mation properties of a multilayered feedforward artificial neural network. Advances in Computational Mathematics, 1(1):61–80, February 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02070821.

Macon:2002:RSC

[MKS02] Charles A. Macon, Leo C. Kempel, and Stephen W. Schneider. Radiation and scattering by complex conformal antennas on a circular cylinder. Advances in Computational Mathematics, 16(2–3):191–209, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/

                                        10.1023/A:1014425527687. Mod-

eling and computation in optics

and electromagnetics.

Mason:1993:E

[MM93] John C. Mason and Charles A. Micchelli. Editorial. Advances in Computational Mathematics, 1(1):i, February 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02070817.

Mason:1999:CE

[MM99] John C. Mason and Charles A. Micchelli. The current Editor. Advances in Computational Mathematics, 11(2–3):i, November 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018984211687.

Maiorov:2000:NOS

[MM00] V. E. Maiorov and R. Meir. On the near optimality of the stochastic approximation of smooth functions by neural networks. Advances in Computational Mathematics, 13(1): 79–103, April 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1023/A:1018993908478.

Micchelli:2013:RSS

[MMP13] Charles A. Micchelli, Jean M. Morales, and Massimiliano Pontil. Regularizers for structured sparsity. Advances in Computational Mathematics, 38(3):455–489, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9245-9.

Morigi:2000:SRC

[MN00] Serena Morigi and Marian Neamtu. Some results for a class of generalized polynomials. Advances in Computational Mathematics, 12(2–3):133–149, February 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018908917139.

Manzoni:2015:HSA

[MN15] Andrea Manzoni and Federico Negri. Heuristic strategies for the approximation of stability factors in quadratically nonlinear parametrized PDEs. Advances in Computational Mathematics, 41(5):1255–1288, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9413-4.

Mallik:2016:CFE

[MN16] Gouranga Mallik and Neela Nataraj. Conforming finite element methods for the von Kármán equations. Advances in Computational Mathematics, 42(5):1031–1054, October 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9452-5.

Mustapha:2016:CSA

[MNC16] Kassem Mustapha, Maher Nour, and Bernardo Cockburn. Convergence and superconvergence analyses of HDG methods for time fractional diffusion problems. Advances in Computational Mathematics, 42(2):377–393, April 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9428-x.

Mukherjee:2006:LTS

[MNPR06] Sayan Mukherjee, Partha Niyogi, Tomaso Poggio, and Ryan Rifkin. Learning theory: stability is sufficient for generalization and necessary and sufficient for consistency of empirical risk minimization. Advances in Computational Mathematics, 25(1–3):161–193, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7634-z.

Mhaskar:2000:PFS

[MNPW00] H. N. Mhaskar, F. J. Narcowich, J. Prestin, and J. D. Ward. Polynomial frames on the sphere. Advances in Computational Mathematics, 13 (4):387–403, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016639802349.

Meinardus:1996:BSA

[MNW96] G. Meinardus, G. Nürnberger, and G. Walz. Bivariate segment approximation and splines. Advances in Computational Mathematics, 6(1):25–45, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127694.

Mhaskar:1999:APZ

[MNW99] H. N. Mhaskar, F. J. Narcowich, and J. D. Ward. Approximation properties of zonal function networks using scattered data on the sphere. Advances in Computational Mathematics, 11(2–3):121–137, November 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018967708053. Ra-

dial basis functions and their applications.

Mhaskar:2004:RBD

[MNW04] H. N. Mhaskar, F. J. Narcowich, and J. D. Ward. On the representation of band-dominant functions on the sphere using finitely many bits. Advances in Computational Mathematics, 21(1–2):127–146, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/B:ACOM.0000016434.
                                        97920.1f.

Marthinsen:2016:GIN

[MO16] Håkon Marthinsen and Brynjulf Owren. Geometric integration of non-autonomous linear Hamiltonian problems. Advances in Computational Mathematics, 42(2):313–332, April 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-015-9425-0.

Monterde:2009:CHP

[Mon09] J. Monterde. A characterization of helical polynomial curves of any degree. Advances in Computational Mathematics, 30(1):61–78, January 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9063-x.

Moosmuller:2017:HSM

[Moo17]  Caroline Moosmüller.  Her-

mite subdivision on manifolds via parallel transport. Advances in Computational Mathematics, 43(5):1059–1074, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9516-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        017-9516-1.pdf.

Mossner:2010:JSR

[Möß10] Bernhard Mößner. On the joint spectral radius of matrices of order 2 with equal spectral radius. Advances in Computational Mathematics, 33(2):243–254, August 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9130-y.

Mortreux:1996:APA

[MP96] P. Mortreux and M. Prévost.

An acceleration property of the E-algorithm for alternate sequences. Advances in Computational Mathematics, 5 (1):443–482, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124755.

Mhaskar:2000:DSP

[MP00] H. N. Mhaskar and J. Prestin.

On the detection of singularities of a periodic function. Advances in Computational Mathematics, 12(2–3):95–131, February 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018921319865.

Mainar:2004:QCC

[MP04] E. Mainar and J. M. Peña.

Quadratic-cycloidal curves. Advances in Computational Mathematics, 20(1–3):161–175, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1025813919473. Ap-

proximation and applications.

Marcellan:1996:OPL

[MPS96] Francisco Marcellán, Franz Peherstorfer, and Robert Steinbauer. Orthogonality properties of linear combinations of orthogonal polynomials. Advances in Computational Mathematics, 5(1):281–295, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124748.

Marcellan:1997:OPL

[MPS97] Francisco Marcellán, Franz Peherstorfer, and Robert Steinbauer. Orthogonality properties of linear combinations of orthogonal polynomials. II. Advances in Computational Mathematics, 7(3):401–428, May 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018963323132.

Marusic:1996:CMS

[MR96] Miljenko Marušić and Mladen Rogina. A collocation method for singularly perturbed twopoint boundary value problems with splines in tension. Advances in Computational Mathematics, 6(1):65–76, December 1996. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127696.

Massey:2010:MCF

[MR10] Pedro Massey and Mariano Ruiz. Minimization of convex functionals over frame operators. Advances in Computational Mathematics, 32 (2):131–153, February 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9092-5.

Madden:2015:MSG

[MR15] Niall Madden and Stephen Russell. A multiscale sparse grid finite element method for a two-dimensional singularly perturbed reaction–diffusion problem. Advances in Computational Mathematics, 41(6): 1423–1450, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9395-7.

Martini:2015:RBA

[MRH15] Immanuel Martini, Gianluigi Rozza, and Bernard Haasdonk. Reduced basis approximation and a-posteriori error estimation for the coupled Stokes– Darcy system. Advances in Computational Mathematics, 41 (5):1131–1157, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9396-6.

Matache:2003:HOC

[MRS03] Ana-Maria Matache, Tomáš Roubíček, and Christoph Schwab. Higher-order convex approximations of Young measures in optimal control. Advances in Computational Mathematics, 19(1–3):73–97, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1022883930432. Chal-

lenges in computational mathematics (Pohang, 2001).

Massey:2014:OFC

[MRS14] Pedro G. Massey, Mariano A. Ruiz, and Demetrio Stojanoff. Optimal frame completions. Advances in Computational Mathematics, 40(5– 6):1011–1042, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9339-7.

Massey:2018:FCP

[MRS18] Pedro G. Massey, Noelia B. Rios, and Demetrio Stojanoff. Frame completions with prescribed norms: local minimizers and applications. Advances in Computational Mathematics, 44(1):51–86, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9535-y.

Micchelli:1997:RM

[MS97] C. A. Micchelli and Thomas Sauer. Regularity of multiwavelets. Advances in Computational Mathematics, 7(4):455–545, August 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018971524949.

Mund:1998:PGM

[MS98] Patrick Mund and Ernst P.

Stephan. The preconditioned GMRES method for systems of coupled FEM–BEM equations. Advances in Computational Mathematics, 9(1– 2):131–144, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018921003084. Nu-

merical treatment of boundary integral equations.

McLean:1999:BEP

[MS99] W. McLean and O. Steinbach.

Boundary element preconditioners for a hypersingular integral equation on an interval. Advances in Computational Mathematics, 11(4):271–286, December 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018944530343.

Moller:2000:BPI

[MS00] H. Michael Möller and Thomas Sauer. H-bases for polynomial interpolation and system solving. Advances in Computational Mathematics, 12(4):335–362, March 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018937723499. Mul-

tivariate polynomial interpolation.

Moller:2004:MRF

[MS04] H. Michael Möller and Thomas Sauer. Multivariate refinable functions of high approximation order via quotient ideals of Laurent polynomials. Advances in Computational Mathematics, 20(1–3):205–228, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1025889132677. Ap-

proximation and applications.

Manni:2007:QSQ

[MS07] Carla Manni and Paul Sablonnière. Quadratic spline quasi-interpolants on Powell–Sabin partitions. Advances in Computational Mathematics, 26(1–3):283–304, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-006-9025-0.

Merrien:2012:HSS

[MS12] Jean-Louis Merrien and Tomas Sauer. From Hermite to stationary subdivision schemes in one and several variables. Advances in Computational Mathematics, 36(4):547–579, May 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9190-7.

Munch:2016:MFD

[MS16] Arnaud Münch and Diego A.

Souza. A mixed formulation for the direct approximation of L^{2}-weighted controls for the linear heat equation. Advances in Computational Mathematics, 42(1):85–125, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9412-5.

Mazroui:2002:ECS

[MSS02] A. Mazroui, D. Sbibih, and P. Sablonnière. Existence and construction of H_{1}-splines of class C^{k} on a three directional mesh. Advances in Computational Mathematics, 17(3):167–198, October 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016066825676.

Maass:2014:SRU

[MSST14] Peter Maass, Chen Sagiv, Hans-Georg Stark, and Bruno Torresani. Signal representation, uncertainty principles and localization measures. Advances in Computational Mathematics, 40(3):597–607, June 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9341-8.

Mazroui:2005:SMS

[MST05] A. Mazroui, D. Sbibih, and A. Tijini. A simple method for smoothing functions and compressing Hermite data. Advances in Computational Mathematics, 23(3):279–297, October 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1783-y.

Micchelli:2013:PAL

[MSXZ13] Charles A. Micchelli, Lixin Shen, Yuesheng Xu, and Xueying Zeng. Proximity algorithms for the L1/TV image denoising model. Advances in Computational Mathematics, 38(2):401–426, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9243-y.

Mao:2006:AFE

[MSZ06] Dong Mao, Lihua Shen, and Aihui Zhou. Adaptive finite element algorithms for eigenvalue problems based on local averaging type a posteriori error estimates. Advances in Computational Mathematics, 25(1–3):135–160, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7617-0.

Mehlum:1998:ISM

[MT98] Even Mehlum and Christian Tarrou. Invariant smoothness measures for surfaces. Advances in Computational Mathematics, 8(1–2):49–63, January 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018931910836.

Mikula:2014:NTS

[MU14] Karol Mikula and Jozef Urbán.

A new tangentially stabilized 3D curve evolution algorithm and its application in virtual colonoscopy. Advances in Computational Mathematics, 40(4):819–837, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9328-x.

Muir:1999:ODC

[Mui99] P. H. Muir. Optimal discrete and continuous monoimplicit Runge–Kutta schemes for BVODEs. Advances in Computational Mathematics, 10 (2):135–167, February 1999.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018926631734.

Muradova:2008:SMN

[Mur08]  Aliki D. Muradova.  The

spectral method and numerical continuation algorithm for the von Kármán problem with postbuckling behaviour of solutions. Advances in Computational Mathematics, 29(2):179–206, August 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9050-7.

Mason:1999:CPM

[MV99] J. C. Mason and E. Venturino. A Chebyshev polynomial method for line integrals with singularities. Advances in Computational Mathematics, 10(2):187–208, February 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018978615805.

Meister:2001:EPL

[MV01] Andreas Meister and Christof Vömel. Efficient preconditioning of linear systems arising from the discretization of hyperbolic conservation laws. Advances in Computational Mathematics, 14(1):49–73, January 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016645505973.

Melenk:2001:RBP

[MW01] J. M. Melenk and B. I.

Wohlmuth. On residual-based a posteriori error estimation in hp–FEM. Advances in Computational Mathematics, 15(1– 4):311–331, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014268310921. A

posteriori error estimation and adaptive computational methods.

Micchelli:2013:AAP

[MWW13] Charles A. Micchelli, Jianzhong Wang, and Yi Wang. On an asymptotic analysis of polynomial approximation to halfband filters. Advances in Computational Mathematics, 38(3):601–622, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9251-y.

Melenk:2013:ARS

[MXO13] J. M. Melenk, C. Xenophontos, and L. Oberbroeckling. Analytic regularity for a singularly perturbed system of reaction– diffusion equations with multiple scales. Advances in Computational Mathematics, 39(2):367–394, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-012-9284-x.

Micchelli:2013:CHS

[MXY13] Charles A. Micchelli, Yuesheng Xu, and Bo Yu. On computing with the Hilbert spline transform. Advances in Computational Mathematics, 38(3):623–646, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9252-x.

Micchelli:2008:TIO

[MXZ08] Charles A. Micchelli, Yuesheng Xu, and Haizhang Zhang. On translation invariant operators which preserve the Bspline recurrence. Advances in Computational Mathematics, 28(2):157–169, February 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-006-9018-z.

Mu:2017:SFE

[MY17] Lin Mu and Xiu Ye. A simple finite element method for the Stokes equations. Advances in Computational Mathematics, 43(6):1305–1324, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9526-z.

Micchelli:2013:DR

[MZ13] Charles A. Micchelli and Tao Zhao. On discrete ℓ^{1}-regularization. Advances in Computational Mathematics, 38(2):441–454, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
011-9247-7.

Nair:2012:QBC

[Nai12] M. Thamban Nair. Quadrature based collocation methods for integral equations of the first kind. Advances in Computational Mathematics, 36 (2):315–329, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9196-1.

Neytcheva:2011:FEB

[NBL11] Maya Neytcheva, Erik Bängtsson, and Elisabeth Linnér. Finiteelement based sparse approximate inverses for blockfactorized preconditioners. Advances in Computational Mathematics, 35(2–4):323–355, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9176-5.

Vu:2013:SAD

[nCV13] B. ng Công Vũ. A splitting algorithm for dual monotone inclusions involving cocoercive operators. Advances in Computational Mathematics, 38(3):667–681, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9254-8.

Nees:1996:CAD

[Nee96] Manuela Nees. Chebyshev approximation by discrete superposition. application to neural networks. Advances in Computational Mathematics, 5 (1):137–151, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124739.

Nestler:2016:APT

[Nes16]  Franziska Nestler.  Auto-

mated parameter tuning based on RMS errors for nonequispaced FFTs. Advances in Computational Mathematics, 42(4):889–919, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9446-8.

Newby:1996:JCH

[New96] J. C. Newby. John Crank: his association with Brunel University. Advances in Computational Mathematics, 6(1):ix–x, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). John Crank 80th birthday special issue.

Niyogi:1999:GBF

[NG99] Partha Niyogi and Federico Girosi. Generalization bounds for function approximation from scattered noisy data. Advances in Computational Mathematics, 10(1):51–80, January 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018966213079.

Niessen:1998:ENR

[Nie98] Gero Nießen. An explicit norm representation for the analysis of multilevel methods. Advances in Computational Mathematics, 9(3–4):311–335, November 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018949809444.

Nielsen:2007:PSS

[Nie07]  Morten Nielsen.  On poly-

nomial symbols for subdivision schemes. Advances in Computational Mathematics, 27(2):195–209, August 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7476-3.

Noakes:1996:NCB

[Noa96]  J. Lyle Noakes.  Networks

and closed balls. Advances in Computational Mathematics, 5(1):153–161, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124740.

Noakes:1998:NCC

[Noa98]  Lyle Noakes.  Nonlinear

corner-cutting. Advances in Computational Mathematics, 8(4):165–177, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018940112654.

Noakes:2002:QIS

[Noa02] Lyle Noakes. Quadratic interpolation on spheres. Advances in Computational Mathematics, 17(4):385–395, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1016277023669.

Noakes:2006:DRC

[Noa06]  Lyle Noakes.  Duality and

Riemannian cubics. Advances in Computational Mathematics, 25(1–3):195–209, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7621-4.

Noakes:2014:MAR

[Noa14]  Lyle Noakes.  Minimum

L^{∞} accelerations in Riemannian manifolds. Advances in Computational Mathematics, 40(4):839–863, August 2014.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9329-9.

Noelle:1995:CHO

[Noe95] Sebastian Noelle. Convergence of higher order finite volume schemes on irregular grids. Advances in Computational Mathematics, 3(3):197–218, April 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02431999.

Novak:2000:HDI

[Nov00] Erich Novak. High dimensional integration. Advances in Computational Mathematics, 12(1):1–2, January 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018996514413.

Navas-Palencia:2018:HPC

[NP18] Guillermo Navas-Palencia. Highprecision computation of the confluent hypergeometric functions via Franklin–Friedman expansion. Advances in Computational Mathematics, 44(??):1– 19, 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

Neamtu:2004:AOS

[NS04] Marian Neamtu and Larry L.

Schumaker. On the approximation order of splines on spherical triangulations. Advances in Computational Mathematics, 21(1–2):3–20, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000016430.
93487.ec.

Nouisser:2004:PBS

[NSS04] O. Nouisser, D. Sbibih, and P. Sablonnière. Pairs of Bsplines with small support on the four-directional mesh generating a partition of unity. Advances in Computational Mathematics, 21(3–4):317–355, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000032048.
57043.dc.

Narcowich:1998:SRS

[NSW98] F. J. Narcowich, N. Sivakumar, and J. D. Ward. Stability results for scattered-data interpolation on Euclidean spheres. Advances in Computational Mathematics, 8(4):137–163, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018996230401.

Narcowich:2007:APR

[NSW07] Francis J. Narcowich, Xinping Sun, and Joseph D. Ward. Approximation power of RBFs and their associated SBFs: a connection. Advances in Computational Mathematics, 27(1):107–124, July 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7506-1.

Nuyens:2016:RLR

[NSW16] Dirk Nuyens, Gowri Suryanarayana, and Markus Weimar. Rank-1 lattice rules for multivariate integration in spaces of permutation-invariant functions: error bounds and tractability. Advances in Computational Mathematics, 42(1):55–84, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9411-6.

Nurnberger:2004:LIC

[NSZ04] Günther Nürnberger, Larry L.

Schumaker, and Frank Zeilfelder. Lagrange interpolation by C^{1} cubic splines on triangulated quadrangulations. Advances in Computational Mathematics, 21(3–4):357–380, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000032044.
49282.8a.

Novak:2017:COI

[NUWZ17] Erich Novak, Mario Ullrich, Henryk Woźniakowski, and Shun Zhang. Complexity of oscillatory integrals on the real line. Advances in Computational Mathematics, 43(3):537–553, June 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9496-6.

Nicaise:2010:QED

[NV10] Serge Nicaise and Julie Valein. Quasi exponential decay of a finite difference space discretization of the 1-d wave equation by pointwise interior stabilization. Advances in Computational Mathematics, 32(3):303–334, April 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9108-1.

Narcowich:2004:PCA

[NW04] Francis J. Narcowich and Joseph D. Ward. Preface [Computational and approximation on the sphere]. Advances in Computational Mathematics, 21(1–2):1–2, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000016666.
68274.5d.

Nurnberger:2004:LIB

[NZ04] Günther Nürnberger and Frank Zeilfelder. Lagrange interpolation by bivariate C^{1}-splines with optimal approximation order. Advances in Computational Mathematics, 21 (3–4):381–419, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000032043.
07621.62.

Ober-Blobaum:2015:CAH

[OBS15] Sina Ober-Blöbaum and Nils Saake. Construction and analysis of higher order Galerkin variational integrators. Advances in Computational Mathematics, 41(6):1391–1422, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9394-8.

Ockendon:1996:RCG

[Ock96] J. R. Ockendon. The role of the Crank–Gupta model in the theory of free and moving boundary problems. Advances in Computational Mathematics, 6(1):281–293, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127708.  John

Crank 80th birthday special issue.

Onchis:2014:OFM

[Onc14]  Darian M. Onchis.  Op-

timized frames and multidimensional challenges in timefrequency analysis. Advances in Computational Mathematics, 40(3):703–709, June 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9332-1.

ORiordan:2011:PUN

[OP11] E. O’Riordan and M. L. Pickett. A parameter-uniform numerical method for a singularly perturbed two parameter elliptic problem. Advances in Computational Mathematics, 35(1):57–82, July 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9164-1.

Opfer:2006:MK

[Opf06]  Roland Opfer.  Multiscale

kernels. Advances in Computational Mathematics, 25 (4):357–380, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7622-3.

ORiordan:2009:NAS

[OS09] Eugene O’Riordan and Martin Stynes. Numerical analysis of a strongly coupled system of two singularly perturbed convection–diffusion problems. Advances in Computational Mathematics, 30(2):101–121, February 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
007-9058-z.

Oswald:2004:SNM

[Osw04]  Peter Oswald.  Smooth-

ness of nonlinear medianinterpolation subdivision. Advances in Computational Mathematics, 20(4):401–423, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1027315032100.

Pan:1995:AAA

[Pan95] Victor Y. Pan. An algebraic approach to approximate evaluation of a polynomial on a set of real points. Advances in Computational Mathematics, 3(1–2):41–58, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02431995.

Paulsen:2017:SCP

[PC17] William Paulsen and Samuel Cowgill. Solving F (z + 1) = b^{F (z)} in the complex plane. Advances in Computational Mathematics, 43(6): 1261–1282, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9524-1.

Pericleous:1996:FSN

[PCM^{+}96] K. A. Pericleous, M. Cross, G. Moran, P. Chow, and K. S. Chan. Free surface Navier– Stokes flows with simultaneous heat transfer and solidification/melting. Advances in Computational Mathematics, 6 (1):295–308, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02127709.  John

Crank 80th birthday special issue.

Peixin:2005:CCI

[Pei05] Ye Peixin. Computational complexity of the integration problem for anisotropic classes. Advances in Computational Mathematics, 23(4):375–392, November 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1830-8.

Penzl:1998:NSG

[Pen98] Thilo Penzl. Numerical solution of generalized Lyapunov equations. Advances in Computational Mathematics, 8(1–2):33–48, January 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018979826766.

Pena:2011:DDS

[Peñ11] Juan Manuel Peña. Diagonal dominance, Schur complements and some classes of H-matrices and P -matrices. Advances in Computational Mathematics, 35(2–4):357–373, November 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9160-5.

Peters:1996:IRC

[Pet96] Jörg Peters. Interpolation regions for convex cubic curve segments. Advances in Computational Mathematics, 6 (1):87–96, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127698.

Petras:2000:SCE

[Pet00] Knut Petras. On the Smolyak cubature error for analytic functions. Advances in Computational Mathematics, 12(1):71– 93, January 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1023/A:1018904816230.

High dimensional integration.

Petzoldt:2002:PEE

[Pet02] Martin Petzoldt. A posteriori error estimators for elliptic equations with discontinuous coefficients. Advances in Computational Mathematics, 16(1):47–75, January 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1014221125034.

Pelosi:2005:GHI

[PFMS05] Francesca Pelosi, Rida T. Farouki, Carla Manni, and Alessandra Sestini. Geometric Hermite interpolation by spatial Pythagorean-hodograph cubics. Advances in Computational Mathematics, 22(4):325–352, May 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-003-2599-x.

Pottmann:2010:EOM

[PGB10] Helmut Pottmann, Philipp Grohs, and Bernhard Blaschitz. Edge offset meshes in Laguerre geometry. Advances in Computational Mathematics, 33(1):45–73, July 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9119-6.

Peherstorfer:2015:RMS

[PGB15] Benjamin Peherstorfer, Pablo Gómez, and Hans-Joachim Bungartz. Reduced models for sparse grid discretizations of the multi-asset Black–Scholes equation. Advances in Computational Mathematics, 41 (5):1365–1389, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9421-4.

Pottmann:2009:LMS

[PGM09] Helmut Pottmann, Philipp Grohs, and Niloy J. Mitra. Laguerre minimal surfaces, isotropic geometry and linear elasticity. Advances in Computational Mathematics, 31 (4):391–419, November 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9076-5.

Pinkus:2004:SPD

[Pin04] Allan Pinkus. Strictly positive definite functions on a real inner product space. Advances in Computational Mathematics, 20(4):263–271, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1027362918283.

Pitolli:2016:BSN

[Pit16] Francesca Pitolli. Bell-shaped nonstationary refinable ripplets. Advances in Computational Mathematics, 42(6): 1427–1451, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9469-9.

PazCalvo:2001:SEM

[PIZ01] Mari Paz Calvo, Arieh Iserles, and Antonella Zanna. Semiexplicit methods for isospectral flows. Advances in Computational Mathematics, 14(1): 1–24, January 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1023/A:1016635812817.

Prautzsch:1994:CSD

[PK94] Hartmut Prautzsch and Leif Kobbelt. Convergence of subdivision and degree elevation. Advances in Computational Mathematics, 2(1):143–154, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02519040.

Piquemal:2005:NWP

[PL05] Anne-Sophie Piquemal and Jacques Liandrat. A new wavelet preconditioner for finite difference operators. Advances in Computational Mathematics, 22(2):125–163, February 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1111-6.

Plantevin:1995:WIM

[Pla95] Frédérique Plantevin. Wavelets on irregular meshes. Advances in Computational Mathematics, 4(1–2):293–329, December 1995. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02123479.

Plaskota:2000:EDS

[Pla00] Leszek Plaskota. The exponent of discrepancy of sparse grids is at least 2.1933. Advances in Computational Mathematics, 12(1):3–24, January 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018900715321. High

dimensional integration.

Plato:2012:RPC

[Pla12] Robert Plato. The regularizing properties of the composite trapezoidal method for weakly singular Volterra integral equations of the first kind. Advances in Computational Mathematics, 36(2):331–351, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9217-0.

Plonka:1995:TSS

[Plo95] Gerlind Plonka. Two-scale symbol and autocorrelation symbol for B-splines with multiple knots. Advances in Computational Mathematics, 3(1–2):1–22, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02431993.

Pandey:2017:HAS

[PM17] Rishi Kumar Pandey and Hradyesh Kumar Mishra. Homotopy analysis Sumudu transform method for time-fractional third order dispersive partial differential equation. Advances in Computational Mathematics, 43(2):365–383, April 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9489-5.

Pottmann:1995:CDR

[Pot95] Helmut Pottmann. Curve design with rational Pythagoreanhodograph curves. Advances in Computational Mathematics, 3(1–2):147–170, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF03028365.

Potts:2004:ASD

[Pot04] Daniel Potts. Approximation of scattered data by trigonometric polynomials on the torus and the 2-sphere. Advances in Computational Mathematics, 21(1–2):21–36, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/B:ACOM.0000016432.
                                        97333.e8.

Pouly:1997:AFE

[PP97] L. Pouly and J. Pousin.

Adaptive finite element for semi-linear convection–diffusion problems. Advances in Computational Mathematics, 7(3): 235–259, May 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1023/A:1018946919497.

Prautzsch:1999:DEP

[PR99] Hartmut Prautzsch and Ulrich Reif. Degree estimates for C^{k}-piecewise polynomial subdivision surfaces. Advances in Computational Mathematics, 10(2):209–217, February 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018922530826.

Pilarczyk:2015:CCH

[PR15] Paweł Pilarczyk and Pedro Real. Computation of cubical homology, cohomology, and (co)homological operations via chain contraction. Advances in Computational Mathematics, 41(1):253–275, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9356-1.

Prautzsch:1998:SSS

[Pra98] Hartmut Prautzsch. Smoothness of subdivision surfaces at extraordinary points. Advances in Computational Mathematics, 9(3–4):377–389, November 1998. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018945708536.

Prevost:1994:APA

[Pré94] Marc Prévost. Acceleration property for the E-algorithm and an application to the summation of series. Advances in Computational Mathematics, 2(3):319–341, April 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02521114.

Prossdorf:1998:MAM

[PS98] Siegfried Prössdorf and Jörg Schult. Multiwavelet approximation methods for pseudodifferential equations on curves. Stability and convergence analysis. Advances in Computational Mathematics, 9(1– 2):145–171, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018977120831. Nu-

merical treatment of boundary integral equations.

Pousin:2005:PEE

[PS05] J. Pousin and T. Sassi. A posteriori error estimates and domain decomposition with nonmatching grids. Advances in Computational Mathematics, 23(3):241–263, October 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1779-7.

Pena:2007:EPR

[PS07] Juan Manuel Peña and Tomas Sauer. Efficient polynomial reduction. Advances in Computational Mathematics, 26 (1–3):323–336, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7208-0.

Pradhan:2011:RTN

[PSNP11] Debasish Pradhan, Baskar Shalini, Neela Nataraj, and Amiya K. Pani. A Robin-type non-overlapping domain decomposition procedure for second order elliptic problems. Advances in Computational Mathematics, 34(4):339–368, May 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9157-0.

Plonka:1995:CWB

[PST95] Gerlind Plonka, Kathi Selig, and Manfred Tasche. On the construction of wavelets on a bounded interval. Advances in Computational Mathematics, 4(1–2):357–388, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02123481.

Plonka:2008:NSB

[PST08] Gerlind Plonka, Hagen Schumacher, and Manfred Tasche. Numerical stability of biorthogonal wavelet transforms. Advances in Computational Mathematics, 29(1):1–25, July 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9034-7.

Paluszynski:2003:TFW

[PŠWX03] Maciej Paluszyński, Hrvoje Šikić, Guido Weiss, and Shaoliang Xiao. Tight frame wavelets, their dimension functions, MRA tight frame wavelets and connectivity properties. Advances in Computational Mathematics, 18(2–4):297–327, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1021312110549. Frames.

Pham:2012:OAS

[PTC12] Duong Pham, Thanh Tran, and Simon Crothers. An overlapping additive Schwarz preconditioner for the Laplace–Beltrami equation using spherical splines. Advances in Computational Mathematics, 37(1):93–121, July 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9200-9.

Pottmann:1994:HSE

[PW94] Helmut Pottmann and Michael G.

Wagner. Helix splines as an example of affine Tchebycheffian splines. Advances in Computational Mathematics, 2(1):123–142, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02519039.

Pottmann:2008:FGC

[PW08] Helmut Pottmann and Johannes Wallner. The focal geometry of circular and conical meshes. Advances in Computational Mathematics, 29(3):249–268, October 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9045-4.

Plonka:2016:RBT

[PZ16] Gerlind Plonka and Yi Zheng.

Relation between total variation and persistence distance and its application in signal processing. Advances in Computational Mathematics, 42(3):651–674, June 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9438-8.

Qin:2012:ISP

[QC12] Hai-Hua Qin and David Colton.

The inverse scattering problem for cavities with impedance boundary condition. Advances in Computational Mathematics, 36(2):157–174, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9179-2.

Qian:2011:AFS

[QW11] Tao Qian and Yan-Bo Wang. Adaptive Fourier series — a variation of greedy algorithm. Advances in Computational Mathematics, 34(3):279–293, April 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9153-4.

Qian:2010:OBN

[QWXZ10] Tao Qian, Rui Wang, Yuesheng Xu, and Haizhang Zhang. Orthonormal bases with nonlinear phases. Advances in Computational Mathematics, 33(1):75–95, July 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9120-0.

Romero:2009:GAP

[RAB^{+}09] Juan R. Romero, Simon K. Alexander, Shikha Baid, Saurabh Jain, and Manos Papadakis. The geometry and the analytic properties of isotropic multiresolution analysis. Advances in Computational Mathematics, 31(1–3):283–328, October 2009. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9111-6.

Ramaswami:1995:PCS

[Ram95]  G. Ramaswami.  Perturbed

collocation and symplectic RKN methods. Advances in Computational Mathematics, 3(1–2):23–40, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02431994.

Rauhut:2005:BTL

[Rau05] Holger Rauhut. Best time localized trigonometric polynomials and wavelets. Advances in Computational Mathematics, 22(1):1–20, January 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7411-7.

Reichel:1993:CPO

[Rei93] Lothar Reichel. Construction of polynomials that are orthogonal with respect to a discrete bilinear form. Advances in Computational Mathematics, 1(2):241–258, June 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02071388.

Remogna:2012:BCS

[Rem12] Sara Remogna. Bivariate C^{2} cubic spline quasi-interpolants on uniform Powell–Sabin triangulations of a rectangular domain. Advances in Computational Mathematics, 36(1):39–65, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9178-3.

Ren:2009:CWT

[Ren09] Jian Ren. A cryptographic watermarking technique for multimedia signals. Advances in Computational Mathematics, 31 (1–3):267–281, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9096-1.

Rand:2013:IEE

[RGB13] Alexander Rand, Andrew Gillette, and Chandrajit Bajaj. Interpolation error estimates for mean value coordinates over convex polygons. Advances in Computational Mathematics, 39(2):327–347, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-012-9282-z.

Raus:2012:QOR

[RH12] Toomas Raus and Uno Hämarik. On the quasi-optimal rules for the choice of the regularization parameter in case of a noisy operator. Advances in Computational Mathematics, 36 (2):221–233, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9203-6.

Redeker:2015:PER

[RH15] Magnus Redeker and Bernard Haasdonk. A POD–EIM reduced two-scale model for crystal growth. Advances in Computational Mathematics, 41 (5):987–1013, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9367-y.

Rong:2017:NAS

[RHZ17] Yao Rong, Yanren Hou, and Yuhong Zhang. Numerical analysis of a second order algorithm for simplified magnetohydrodynamic flows. Advances in Computational Mathematics, 43(4):823–848, August 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9508-6.

Rippa:1999:ASG

[Rip99] Shmuel Rippa. An algorithm for selecting a good value for the parameter c in radial basis function interpolation. Advances in Computational Mathematics, 11(2–3):193–210, November 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018975909870. Ra-

dial basis functions and their applications.

Reddy:2000:ADO

[RJ00] Muni V. Reddy and Stephen Joe. An average discrepancy for optimal vertex-modified number-theoretic rules. Advances in Computational Mathematics, 12(1):59–69, January 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018952732160. High

dimensional integration.

Rolania:2007:ABS

[RL07] D. Barrios Rolanía and G. López Lagomasino. Asymptotic behavior of solutions of general three term recurrence relations. Advances in Computational Mathematics, 26(1–3):9–37, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-4137-x.

Roberts:1998:VA

[Rob98] D. E. Roberts. On a vector q–d algorithm. Advances in Computational Mathematics, 8(4):193–219, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018944213562.

Riemenschneider:1999:GRB

[RS99] S. D. Riemenschneider and N. Sivakumar. Gaussian radialbasis functions: cardinal interpolation of l^{p} and powergrowth data. Advances in Computational Mathematics, 11 (2–3):229–251, November 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018980110778. Ra-

dial basis functions and their applications.

Reif:2001:CIS

[RS01] Ulrich Reif and Peter Schröder.

Curvature integrability of subdivision surfaces. Advances in Computational Mathematics, 14(2):157–174, February 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016685104156.

Romani:2014:ATP

[RSA14] Lucia Romani, Laura Saini, and Gudrun Albrecht. Algebraic– trigonometric Pythagorean-hodograph curves and their use for Hermite interpolation. Advances in Computational Mathematics, 40(5–6):977–1010, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9338-8.

Ricaud:2014:OCG

[RST^{+}14] Benjamin Ricaud, Guillaume Stempfel, Bruno Torrésani, Christoph Wiesmeyr, Hélène Lachambre, and Darian Onchis. An optimally concentrated Gabor transform for localized timefrequency components. Advances in Computational Mathematics, 40(3):683–702, June 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9337-9.

Rohwedder:2011:PPI

[RSZ11] Thorsten Rohwedder, Reinhold Schneider, and Andreas Zeiser. Perturbed preconditioned inverse iteration for operator eigenvalue problems with applications to adaptive wavelet discretization. Advances in Computational Mathematics, 34(1):43–66, January 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9141-8.

Ricaud:2014:SUP

[RT14] Benjamin Ricaud and Bruno Torrésani. A survey of uncertainty principles and some signal processing applications. Advances in Computational Mathematics, 40(3):629–650, June 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9323-2.

Rong:2013:AWC

[RWT13] Zhijian Rong, Li-Lian Wang, and Xue-Cheng Tai. Adaptive wavelet collocation methods for image segmentation using TV– Allen–Cahn type models. Advances in Computational Mathematics, 38(1):101–131, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9227-y.

Rim:2013:GPR

[RY13] Kyung Soo Rim and Beong In Yun. Gibbs phenomenon removal by adding Heaviside functions. Advances in Computational Mathematics, 38(4):683–699, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9255-7.

Richardson:2016:MIR

[RY16] Casey L. Richardson and Laurent Younes. Metamorphosis of images in reproducing kernel Hilbert spaces. Advances in Computational Mathematics, 42(3):573–603, June 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9435-y.

Rynne:2000:CGM

[Ryn00]  Bryan P. Rynne.  Conver-

gence of Galerkin method solutions of the integral equation for thin wire antennas. Advances in Computational Mathematics, 12(2–3):251–259, February 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018956800300.

Rude:1998:MPE

[RZ98] Ulrich Rüde and Aihui Zhou. Multi-parameter extrapolation methods for boundary integral equations. Advances in Computational Mathematics, 9(1– 2):173–190, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018989523557. Nu-

merical treatment of boundary integral equations.

Rieger:2010:SII

[RZ10] Christian Rieger and Barbara Zwicknagl. Sampling inequalities for infinitely smooth functions, with applications to interpolation and machine learning. Advances in Computational Mathematics, 32 (1):103–129, January 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9089-0;
                                         http://link.springer.com/

content/pdf/10.1007/s10444-
008-9089-0.pdf.

Sablonniere:2004:BTB

[Sab04] Paul Sablonnière. Bernsteintype bases and corner cutting algorithms for C^{1} Merrien’s curves. Advances in Computational Mathematics, 20 (1–3):229–246, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1025870005361. Ap-

proximation and applications.

Salam:1998:FVO

[Sal98] A. Salam. Formal vector orthogonal polynomials. Advances in Computational Mathematics, 8(4):267–289, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018952415379.

Sangawi:2015:SSM

[San15] Ali W. K. Sangawi. Straight slits map and its inverse of bounded multiply connected regions. Advances in Computational Mathematics, 41(2):439–455, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9368-x.

Sauer:1995:CAM

[Sau95] Thomas Sauer. Computational aspects of multivariate polynomial interpolation. Advances in Computational Mathematics, 3(3):219–237, April 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02432000.

Sauer:2006:DMR

[Sau06] Thomas Sauer. Differentiability of multivariate refinable functions and factorization. Advances in Computational Mathematics, 25(1–3):211–235, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7635-y.

Shih:2008:EPG

[SC08] Shagi-Di Shih and Shue-Sum Chow. Equivalence of n-point Gauss–Chebyshev rule and 4npoint midpoint rule in computing the period of a Lotka– Volterra system. Advances in Computational Mathematics, 28(1):63–79, January 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-006-9013-4.

Schaback:1995:EEC

[Sch95] Robert Schaback. Error estimates and condition numbers for radial basis function interpolation. Advances in Computational Mathematics, 3(3):251–264, April 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02432002.

Schumaker:2004:PMS

[Sch04] Larry L. Schumaker. Preface [Multivariate splines]. Advances in Computational Mathematics, 21(3–4):221–222, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000032130.
41067.34.

Schaback:2009:SLE

[Sch09]  Robert Schaback.  Solving

the Laplace equation by meshless collocation using harmonic kernels. Advances in Computational Mathematics, 31 (4):457–470, November 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9078-3;
 http://link.springer.com/
content/pdf/10.1007/s10444-
008-9078-3.pdf.

Schaback:2011:MWF

[Sch11a] Robert Schaback. The missing Wendland functions. Advances in Computational Mathematics, 34(1):67–81, January 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9142-7;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        009-9142-7.pdf.

Scheuerer:2011:APS

[Sch11b] Michael Scheuerer. An alternative procedure for selecting a good value for the parameter c in RBF-interpolation. Advances in Computational Mathematics, 34(1):105–126, January 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-010-9146-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        010-9146-3.pdf.

Schaback:2015:CTC

[Sch15] Robert Schaback. A computational tool for comparing all linear PDE solvers: error-optimal methods are meshless. Advances in Computational Mathematics, 41(2):333–355, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9360-5.

Chang:2013:BHM

[sCLC13] Shih sen Chang, H. W. Joseph Lee, and Chi Kin Chan. A block hybrid method for solving generalized equilibrium problems and convex feasibility problem. Advances in Computational Mathematics, 38(3):563–580, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9249-5.

Song:1996:MMB

[SEE96] J. H. Song, M. J. Edirisinghe, and J. R. G. Evans. Mathematical modelling of binder removal from plastically-formed engineering ceramics. Advances in Computational Mathematics, 6(1):325–331, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

Smith:2014:DAF

[SF14] Lindsay N. Smith and Matthew Fickus. Determining angular frequency from images of rotating objects via a generalized Fast Fourier Transform. Advances in Computational Mathematics, 40(1):27–47, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9296-1.

Sommer:2015:FCP

[SFDE15] Alexander Sommer, Ortwin Farle, and Romanus Dyczij-Edlinger. A fast certified parametric near-field-to-far-field transformation technique for electrically large antenna arrays. Advances in Computational Mathematics, 41 (5):1015–1034, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9373-0.

Shekhtman:2008:BIP

[She08]  Boris Shekhtman.  Bivari-

ate ideal projectors and their perturbations. Advances in Computational Mathematics, 29(3):207–228, October 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9040-9.

Shen:2009:BDE

[She09] Jianhong (Jackie) Shen. Beamlets are densely embedded in H^{−1}. Advances in Computational Mathematics, 31 (1–3):329–348, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9094-3.

Sherratt:2013:NCB

[She13] Jonathan A. Sherratt. Numerical continuation of boundaries in parameter space between stable and unstable periodic travelling wave (wavetrain) solutions of partial differential equations. Advances in Computational Mathematics, 39(1):175–192, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-012-9273-0.

Shi:2014:QRR

[SHTS14] Lei Shi, Xiaolin Huang, Zheng Tian, and Johan A. K. Suykens. Quantile regression with ℓ_{1}- regularization and Gaussian kernels. Advances in Computational Mathematics, 40(2):517– 551, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
013-9317-0.

San:2015:SPO

[SI15] Omer San and Traian Iliescu.

A stabilized proper orthogonal decomposition reducedorder model for large scale quasigeostrophic ocean circulation. Advances in Computational Mathematics, 41 (5):1289–1319, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9417-0.

Sidi:2017:MPR

[Sid17] Avram Sidi. Minimal polynomial and reduced rank extrapolation methods are related. Advances in Computational Mathematics, 43(1):151–170, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9481-0.

Silbermann:1999:TPD

[Sil99] Bernd Silbermann. Tribute to Professor Dr. rer. nat. habil. Siegfried Prössdorf 02.01.1939— 19.07.1998. Advances in Computational Mathematics, 11(1):i–ii, July 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018903421256.

Sun:2015:NIS

[SJJL15] Zhiyuan Sun, Yuling Jiao, Bangti Jin, and Xiliang Lu. Numerical identification of a sparse Robin coefficient. Advances in Computational Mathematics, 41(1):131–148, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9352-5.

Sun:2005:GFT

[SL05] Jiachang Sun and Huiyuan Li. Generalized Fourier transform on an arbitrary triangular domain. Advances in Computational Mathematics, 22(3):223–248, April 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-7667-8.

Shen:2012:RIP

[SL12] Yi Shen and Song Li. Restricted p-isometry property and its application for nonconvex compressive sensing. Advances in Computational Mathematics, 37(3):441–452, October 2012. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9219-y.

Schauer:2015:RBM

[SL15] Volker Schauer and Christian Linder. The reduced basis method in all-electron calculations with finite elements. Advances in Computational Mathematics, 41(5):1035–1047, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9374-z.

Sommer:2014:OGE

[SLN14] S. Sommer, F. Lauze, and M. Nielsen. Optimization over geodesics for exact principal geodesic analysis. Advances in Computational Mathematics, 40(2):283–313, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9308-1.

Sun:1999:SPD

[SM99] Xingping Sun and V. A.

Menegatto. Strictly positive definite functions on the complex Hilbert sphere. Advances in Computational Mathematics, 11(2–3):105–119, November 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018915723982. Ra-

dial basis functions and their applications.

Salah:2000:CBS

[SM00] I. Ben Salah and P. Maroni. The connection between self-associated two-dimensional vector functionals and third degree forms. Advances in Computational Mathematics, 13(1):51–77, April 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018941924408.

Smoch:2007:SBG

[Smo07] Laurent Smoch. Spectral behaviour of GMRES applied to singular systems. Advances in Computational Mathematics, 27(2):151–166, August 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-007-9029-4.

Sondergaard:2007:GFS

[Søn07] Peter L. Søndergaard. Gabor frames by sampling and periodization. Advances in Computational Mathematics, 27 (4):355–373, November 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-005-9003-y.

Sorokina:2018:BBT

[Sor18] Tatyana Sorokina. Bernstein– Bézier techniques for divergence of polynomial spline vector fields in R^{n}. Advances in Computational Mathematics, 44(1):227–244, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9541-0.

Speleers:2017:HSS

[Spe17]  Hendrik Speleers.  Hierar-

chical spline spaces: quasiinterpolants and local approximation estimates. Advances in Computational Mathematics, 43(2):235–255, April 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9483-y.

Sprott:2002:KGR

[SR02] Kenneth Sprott and Bahram Ravani. Kinematic generation of ruled surfaces. Advances in Computational Mathematics, 17(1–2):115–133, July 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1015211729988. Ad-

vances in geometrical algorithms and representations.

Song:2012:MII

[SRFH12] Guohui Song, John Riddle, Gregory E. Fasshauer, and Fred J. Hickernell. Multivariate interpolation with increasingly flat radial basis functions of finite smoothness. Advances in Computational Mathematics, 36(3):485–501, April 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9192-5.

Sommer:1996:CSW

[SS96] Manfred Sommer and Hans Strauss. A condition of Schoenberg–Whitney type for multivariate spline interpolation. Advances in Computational Mathematics, 5 (1):381–397, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124752.

Schumaker:2004:QST

[SS04] Larry L. Schumaker and Tatyana Sorokina. C^{1} quintic splines on type-4 tetrahedral partitions. Advances in Computational Mathematics, 21 (3–4):421–444, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000032045.
03775.8d.

Sloan:2008:ASU

[SS08] Ian H. Sloan and Alvise Sommariva. Approximation on the sphere using radial basis functions plus polynomials. Advances in Computational Mathematics, 29(2):147–177, August 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9048-1.

Sun:2013:IFW

[SS13] Xudong Sun and Wenchang Sun. Inversion formula for the windowed Fourier transform, II. Advances in Computational Mathematics, 38(1):21–34, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9223-2.

Schumaker:2014:CPS

[SS14] Larry L. Schumaker and Hendrik Speleers. Convexity preserving C^{0} splines. Advances in Computational Mathematics, 40(1):117–135, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9301-8.

Son:2015:MOR

[SS15] Nguyen Thanh Son and Tatjana Stykel. Model order reduction of parameterized circuit equations based on interpolation. Advances in Computational Mathematics, 41(5):1321–1342, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9418-z.

Santin:2016:AEK

[SS16] G. Santin and R. Schaback. Approximation of eigenfunctions in kernel-based spaces. Advances in Computational Mathematics, 42(4):973–993, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-015-9449-5.

Steidl:1998:NFF

[Ste98] Gabriele Steidl. A note on Fast Fourier Transforms for nonequispaced grids. Advances in Computational Mathematics, 9(3–4):337–352, November 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018901926283.

Stonag:1996:CPL

[Sto96] Eduardo D. Stonag. Critical points for least-squares problems involving certain analytic functions, with applications to sigmoidal nets. Advances in Computational Mathematics, 5(1):245–268, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124746.

Stolk:2017:ISD

[Sto17] Christiaan C. Stolk. An improved sweeping domain decomposition preconditioner for the Helmholtz equation. Advances in Computational Mathematics, 43(1):45–76, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9475-y;
http://link.springer.com/
content/pdf/10.1007/s10444-
016-9475-y.pdf.

Strom:1995:BSH

[Str95] Kyrre Strøm. B-splines with homogenized knots. Advances in Computational Mathematics, 3(3):291–308, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02988627.

Struble:2013:CWA

[Str13] Dale W. Struble. The construction of wavelets adapted to compact domains. Advances in Computational Mathematics, 38(3):581–600, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9250-z.

Sun:2012:LIF

[SU12] Qiyu Sun and Michael Unser.

Left-inverses of fractional Laplacian and sparse stochastic processes. Advances in Computational Mathematics, 36(3): 399–441, April 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.
com/article/10.1007/s10444-
011-9183-6;   http://link.
springer.com/content/pdf/
10.1007/s10444-011-9183-6.
pdf.

Sun:2008:FSF

[Sun08] Qiyu Sun. Frames in spaces with finite rate of innovation. Advances in Computational Mathematics, 28(4):301–329, May 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-006-9021-4.

Sun:2010:LRS

[Sun10] Qiyu Sun. Local reconstruction for sampling in shiftinvariant spaces. Advances in Computational Mathematics, 32(3):335–352, April 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9109-0.

Sun:2014:LNF

[Sun14] Qiyu Sun. Localized nonlinear functional equations and two sampling problems in signal processing. Advances in Computational Mathematics, 40(2):415–458, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9314-3.

Shaw:1996:TAF

[SW96] Simon Shaw and J. R. Whiteman. Towards adaptive finite element schemes for partial differential Volterra equation solvers. Advances in Computational Mathematics, 6 (1):309–323, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127710.  John

Crank 80th birthday special issue.

Steinbach:1998:CSE

[SW98] O. Steinbach and W. L. Wendland. The construction of some efficient preconditioners in the boundary element method. Advances in Computational Mathematics, 9(1– 2):191–216, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018937506719. Nu-

merical treatment of boundary integral equations.

Sloan:2004:ESP

[SW04] Ian H. Sloan and Robert S.

Womersley. Extremal systems of points and numerical integration on the sphere. Advances in Computational Mathematics, 21(1–2):107–125, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/

                                        10.1023/B:ACOM.0000016428.
                                        25905.da.

Schaback:2006:LCR

[SW06] R. Schaback and J. Werner. Linearly constrained reconstruction of functions by kernels with applications to machine learning. Advances in Computational Mathematics, 25(1–3):237–258, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7616-1.

Sorokina:2008:MPS

[SW08] T. Sorokina and A. J. Worsey. A multivariate Powell–Sabin interpolant. Advances in Computational Mathematics, 29(1): 71–89, July 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1007/s10444-
                                        007-9041-8.

Sun:2010:RLS

[SW10] Hongwei Sun and Qiang Wu. Regularized least square regression with dependent samples. Advances in Computational Mathematics, 32 (2):175–189, February 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9099-y.

Sauer:1995:MHI

[SX95] Thomas Sauer and Yuan Xu. On multivariate Hermite interpolation. Advances in Computational Mathematics, 4(1– 2):207–259, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF03177515.

Shi:2007:COP

[SX07] Ying Guang Shi and Guoliang Xu. Construction of σ-orthogonal polynomials and Gaussian quadrature formulas. Advances in Computational Mathematics, 27(1):79–94, July 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9033-8.

Shen:2009:PMM

[SX09] Lixin Shen and Yuesheng Xu.

Preface [Mathematical methods for image processing]. Advances in Computational Mathematics, 31(1–3):1–4, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9113-z;
 http://link.springer.com/
content/pdf/10.1007/s10444-
009-9113-z.pdf.

Shu:2006:AMM

[SXYY06] Shi Shu, Jinchao Xu, Ying Yang, and Haiyuan Yu. An algebraic multigrid method for finite element systems on criss-cross grids. Advances in Computational Mathematics, 25(1–3):287–304, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7627-y.

Schneider:2006:ADG

[SXZ06] Reinhold Schneider, Yuesheng Xu, and Aihui Zhou. An analysis of discontinuous Galerkin methods for elliptic problems. Advances in Computational Mathematics, 25(1–3):259–286, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7619-y.

Sun:2009:CLS

[SZ09] Wenchang Sun and Xingwei Zhou. Characterization of local sampling sequences for spline subspaces. Advances in Computational Mathematics, 30(2):153–175, February 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9062-y.

Tang:2017:SID

[Tan17] Sui Tang. System identification in dynamical sampling. Advances in Computational Mathematics, 43(3):555–580, June 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9497-5.

Turner:2016:TDC

[TB16] M. R. Turner and T. J.

Bridges. Time-dependent conformal mapping of doublyconnected regions. Advances in Computational Mathematics, 42(4):947–972, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9448-6;
 http://link.springer.com/
content/pdf/10.1007/s10444-
015-9448-6.pdf.

Twizell:1994:NME

[TBD94] E. H. Twizell, A. Boutayeb, and K. Djidjeli. Numerical methods for eighth-, tenth- and twelfthorder eigenvalue problems arising in thermal instability. Advances in Computational Mathematics, 2(4):407–436, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02521607.

Tian:2016:ACS

[TDG16] Xiaochuan Tian, Qiang Du, and Max Gunzburger. Asymptotically compatible schemes for the approximation of fractional Laplacian and related nonlocal diffusion problems on bounded domains. Advances in Computational Mathematics, 42(6):1363–1380, December 2016. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9466-z.

Temlyakov:1998:BTA

[Tem98] V. N. Temlyakov. The best m-term approximation and greedy algorithms. Advances in Computational Mathematics, 8(4):249–265, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018900431309.

Temlyakov:2000:WGA

[Tem00]  V. N. Temlyakov.  Weak

greedy algorithms. Advances in Computational Mathematics, 12(2–3):213–227, February 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018917218956.

Temlyakov:2001:GAB

[Tem01] V. N. Temlyakov. Greedy algorithms in Banach spaces. Advances in Computational Mathematics, 14(3):277–292, April 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1016657209416.

Temlyakov:2002:CCW

[Tem02] V. N. Temlyakov. A criterion for convergence of weak greedy algorithms. Advances in Computational Mathematics, 17(3):269–280, October 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016061804993.

Temlyakov:2007:GEB

[Tem07] V. N. Temlyakov. Greedy expansions in Banach spaces. Advances in Computational Mathematics, 26(4):431–449, May 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7452-y.

Twizell:1996:SOS

[TGA96] E. H. Twizell, A. B. Gumel, and M. A. Arigu. Second-order, L_{0}-stable methods for the heat equation with time-dependent boundary conditions. Advances in Computational Mathematics, 6(1):333–352, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127712.  John

Crank 80th birthday special issue.

Takhirov:2018:CEM

[TL18] Aziz Takhirov and Alexander Lozovskiy. Computationally efficient modular nonlinear filter stabilization for high Reynolds number flows. Advances in Computational Mathematics, 44(1):295–325, February 2018. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9544-x.

Tommasini:1996:CRL

[Tom96] T. Tommasini. Complexity reduction of least squares problems involving special Vandermonde matrices. Advances in Computational Mathematics, 6(1):77–86, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02127697.

Tornberg:2016:ESS

[Tor16] Anna-Karin Tornberg. The Ewald sums for singly, doubly and triply periodic electrostatic systems. Advances in Computational Mathematics, 42(1):227–248, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-015-9422-3.

Tao:2017:DDM

[TQ17] Zhanjing Tao and Jianxian Qiu. Dimension-by-dimension moment-based central Hermite WENO schemes for directly solving Hamilton–Jacobi equations. Advances in Computational Mathematics, 43 (5):1023–1058, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9515-2.

Tsukanov:2005:MMA

[TS05] I. Tsukanov and V. Shapiro.

Meshfree modeling and analysis of physical fields in heterogeneous media. Advances in Computational Mathematics, 23(1–2):95–124, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1835-3.

Tan:2010:ROB

[TSY10] Lihui Tan, Lixin Shen, and Lihua Yang. Rational orthogonal bases satisfying the Bedrosian identity. Advances in Computational Mathematics, 33(3):285–303, October 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9133-8.

Tran:1998:WBP

[TSZ98] Thanh Tran, Ernst P. Stephan, and Stefan Zaprianov. Waveletbased preconditioners for boundary integral equations. Advances in Computational Mathematics, 9(1–2):233–249, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018993624466. Nu-

merical treatment of boundary integral equations.

Tausch:1998:SKI

[TW98] Johannes Tausch and Jacob White. Second-kind integral formulations of the capacitance problem. Advances in Computational Mathematics, 9(1– 2):217–232, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018973019922. Nu-

merical treatment of boundary integral equations.

Twizell:1996:P

[Twi96] E. H. Twizell. Preface. Advances in Computational Mathematics, 6(1):i–ii, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02127691.

Temlyakov:2011:GAR

[TYY11] V. N. Temlyakov, Mingrui Yang, and Peixin Ye. Greedy approximation with regard to non-greedy bases. Advances in Computational Mathematics, 34(3):319–337, April 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-010-9155-2.

Tasche:2003:REA

[TZ03] Manfred Tasche and Hansmartin Zeuner. Roundoff error analysis of the recursive moving window discrete Fourier transform. Advances in Computational Mathematics, 18(1):65–78, January 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1021254709641.

Tan:2005:FSR

[TZ05] Jieqing Tan and Ping Zhou. On the finite sum representations of the Lauricella functions F_{D}. Advances in Computational Mathematics, 23(4):333–351, November 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1838-0.

Tebou:2007:UBS

[TZ07] Louis T. Tebou and Enrique Zuazua. Uniform boundary stabilization of the finite difference space discretization of the 1-d wave equation. Advances in Computational Mathematics, 26(1–3):337–365, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7629-9.

Teng:2016:BCD

[TZL16] Zhongming Teng, Yunkai Zhou, and Ren-Cang Li. A block Chebyshev–Davidson method for linear response eigenvalue problems. Advances in Computational Mathematics, 42 (5):1103–1128, October 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9455-2.

Urban:1995:DFW

[Urb95] Karsten Urban. On divergencefree wavelets. Advances in Computational Mathematics, 4 (1–2):51–81, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02123473.

Varas:1996:RSM

[Var96] María Leonor S. Varas. Reconstruction of signals from multiscale edges. Advances in Computational Mathematics, 5(1):31–49, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02124734.

Venturino:2000:CLI

[VC00] E. Venturino and G. A. Chandler. Calculation of line integrals in boundary integral methods. Advances in Computational Mathematics, 13(3):293–318, July 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018914513021.

vanderHouwen:1993:PII

[vdH93] P. J. van der Houwen. Preconditioning in implicit initialvalue problem methods on parallel computers. Advances in Computational Mathematics, 1(1):39–60, February 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02070820.

vanderHouwen:1997:PLS

[vdHdS97] P. J. van der Houwen and J. J. B. de Swart. Parallel linear system solvers for Runge–Kutta methods. Advances in Computational Mathematics, 7(1–2):157–181, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018990601750. Par-

allel methods for ODEs.

vanderHouwen:2000:DIR

[vdHS00] P. J. van der Houwen and B. P.

Sommeijer. Diagonally implicit Runge–Kutta methods for 3D shallow water applications. Advances in Computational Mathematics, 12(2–3):229–250, February 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018969203026.

vanderHouwen:1997:WRM

[vdHv97] P. J. van der Houwen and W. A. van der Veen. Waveform relaxation methods for implicit differential equations. Advances in Computational Mathematics, 7(1–2):183–197, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018942718589. Par-

allel methods for ODEs.

vanderMee:2003:SEI

[vdMNS03] Cornelis V. M. van der Mee, M. Z. Nashed, and Sebastiano Seatzu. Sampling expansions and interpolation in unitarily translation invariant reproducing kernel Hilbert spaces. Advances in Computational Mathematics, 19(4):355–372, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1024233232215.

Venturino:1994:LIA

[Ven94]  Ezio Venturino.  A lacu-

nary interpolation algorithm on arbitrary points. Advances in Computational Mathematics, 2(2):223–233, March 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02521109.

Voss:1996:TSA

[VK96] D. A. Voss and A. Q. M. Khaliq. Time-stepping algorithms for semidiscretized linear parabolic PDEs based on rational approximants with distinct real poles. Advances in Computational Mathematics, 6 (1):353–363, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127713.  John

Crank 80th birthday special issue.

Voronin:2017:EAC

[VM17] Sergey Voronin and Per-Gunnar Martinsson. Efficient algorithms for CUR and interpolative matrix decompositions. Advances in Computational Mathematics, 43(3):495–516, June 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9494-8.

Van:2002:ATD

[VW02] Tri Van and Aihua Wood. Analysis of time-domain Maxwell’s equations for 3-D cavities. Advances in Computational Mathematics, 16(2–3):211–228, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014489832230. Mod-

eling and computation in optics and electromagnetics.

Van:2005:FEA

[VW05] Tri Van and Aihua Wood. Finite element analysis of transient electromagnetic scattering problems. Advances in Computational Mathematics, 22(1):21–48, January 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-005-1176-x.

Walz:1995:UAB

[Wal95] Guido Walz. A unified approach to B-spline recursions and knot insertion, with application to new recursion formulas. Advances in Computational Mathematics, 3(1–2):89–100, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02431997.

Wang:2008:SDQ

[Wan08] Yang Wang. Sigma–delta quantization errors and the traveling salesman problem. Advances in Computational Mathematics, 28(2):101–118, February 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-006-9016-1.

Wang:2012:PNM

[Wan12] Y. F. Wang. Preconditioning non-monotone gradient methods for retrieval of seismic reflection signals. Advances in Computational Mathematics, 36(2):353–376, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9207-2.

Warren:1996:BCC

[War96] Joe Warren. Barycentric coordinates for convex polytopes. Advances in Computational Mathematics, 6(1):97–108, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127699.

Ward:2013:EEA

[War13] John Paul Ward. L^{p} error estimates for approximation by Sobolev splines and Wendland functions on R^{d}. Advances in Computational Mathematics, 38(4):873–889, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9263-7.

Watson:1994:SOP

[Wat94] G. A. Watson. The solution of orthogonal Procrustes problems for a family of orthogonally invariant norms. Advances in Computational Mathematics, 2(4):393–405, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02521606.

Watanabe:1996:SML

[Wat96]  Sumio Watanabe.  Solvable

models of layered neural networks based on their differential structure. Advances in Computational Mathematics, 5 (1):205–231, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124744.

Wendland:1995:PPP

[Wen95] Holger Wendland. Piecewise polynomial, positive definite and compactly supported radial functions of minimal degree. Advances in Computational Mathematics, 4(1–2):389–396, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02123482.

Wu:2002:TMP

[WFL02] Wei Wu, Guorui Feng, and Xin Li. Training multilayer perceptrons via minimization of sum of ridge functions. Advances in Computational Mathematics, 17(4):331–347, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016249727555.

Wei:2007:NDR

[WH07] T. Wei and Y. C. Hon. Numerical differentiation by radial basis functions approximation. Advances in Computational Mathematics, 27(3):247–272, October 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-9001-0.

Wang:2018:NMS

[WHS18] Liqun Wang, Songming Hou, and Liwei Shi. A numerical method for solving threedimensional elliptic interface problems with triple junction points. Advances in Computational Mathematics, 44 (1):175–193, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-017-9539-7.

Wierse:1997:NTM

[Wie97] Monika Wierse. A new theoretically motivated higher order upwind scheme on unstructured grids of simplices. Advances in Computational Mathematics, 7(3):303–335, May 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018955121314.

Wieland:2015:IPM

[Wie15] Bernhard Wieland. Implicit partitioning methods for unknown parameter sets: in the context of the reduced basis method. Advances in Computational Mathematics, 41(5):1159–1186, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9404-5.

Wille:1998:ESC

[Wil98] David R. Willé. Experiments in stepsize control for Adams linear multistep methods. Advances in Computational Mathematics, 8(4):335–344, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018960717197.

Williams:1993:LSC

[WK93] Jack Williams and Z. Kalogiratou. Least squares and Chebyshev fitting for parameter estimation in ODEs. Advances in Computational Mathematics, 1(3):357–366, October 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02072016.

Wu:2005:GSF

[WL05] Zong Min Wu and Jian Ping Liu. Generalized Strang–Fix condition for scattered data quasi-interpolation. Advances in Computational Mathematics, 23(1–2):201–214, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1832-6.

Wang:2012:RDN

[WL12] Haibing Wang and Jijun Liu. Recovering the Dirichlet-to-Neumann map in inverse scattering problems using integral equation methods. Advances in Computational Mathematics, 36(2):279–297, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9191-6.

Witkowski:2015:SCN

[WLPV15] T. Witkowski, S. Ling, S. Praetorius, and A. Voigt. Software concepts and numerical algorithms for a scalable adaptive parallel finite element method. Advances in Computational Mathematics, 41(6): 1581–1613, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9405-4.

Wang:2016:MAL

[WLW16] Chuanlong Wang, Chao Li, and Jin Wang. A modified augmented Lagrange multiplier algorithm for Toeplitz matrix completion. Advances in Computational Mathematics, 42(5):1209–1224, October 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9459-y.

Wang:2013:PCS

[WMB13] Chuan-Long Wang, Guo-Yan Meng, and Yan-Hong Bai. Practical convergent splittings and acceleration methods for non-Hermitian positive definite linear systems. Advances in Computational Mathematics, 39(2):257–271, August 2013.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-012-9278-8.

Wang:2013:MPM

[WMY13] Chuan-Long Wang, Guo-Yan Meng, and Xue-Rong Yong. Modified parallel multisplitting iterative methods for non-Hermitian positive definite systems. Advances in Computational Mathematics, 38(4):859–872, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9262-8.

Wenger:2017:CAH

[WOBL17] Theresa Wenger, Sina Ober-Blöbaum, and Sigrid Leyendecker. Construction and analysis of higher order variational integrators for dynamical systems with holonomic constraints. Advances in Computational Mathematics, 43(5):1163–1195, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9520-5.

Womersley:2001:HGC

[WS01] Robert S. Womersley and Ian H. Sloan. How good can polynomial interpolation on the sphere be? Advances in Computational Mathematics, 14(3):195–226, April 2001. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016630227163.

Warren:2007:BCC

[WSHD07] Joe Warren, Scott Schaefer, Anil N. Hirani, and Mathieu Desbrun. Barycentric coordinates for convex sets. Advances in Computational Mathematics, 27(3):319–338, October 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-9008-6.

Wu:2013:SSD

[WSM13] Zongmin Wu, Xingping Sun, and Limin Ma. Sampling scattered data with Bernstein polynomials: stochastic and deterministic error estimates. Advances in Computational Mathematics, 38(1):187–205, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9233-0.

Wu:1995:CSP

[Wu95] Zongmin Wu. Compactly supported positive definite radial functions. Advances in Computational Mathematics, 4(1– 2):283–292, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF03177517.

Wang:2012:SPI

[WW12] Yang Wang and Qiang Wu. Sparse PCA by iterative elimination algorithm. Advances in Computational Mathematics, 36(1):137–151, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9186-3.

Wang:2016:WGF

[WY16] Junping Wang and Xiu Ye. A weak Galerkin finite element method for the Stokes equations. Advances in Computational Mathematics, 42 (1):155–174, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-015-9415-2.

Wang:2017:HPV

[WY17] Lina Wang and Lijun Yi. An h–p version of the continuous Petrov–Galerkin method for Volterra delay-integro-differential equations. Advances in Computational Mathematics, 43(6): 1437–1467, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9531-2.

Wallner:2011:CSA

[WYW11] Johannes Wallner, Esfandiar Nava Yazdani, and Andreas Weinmann. Convergence and smoothness analysis of subdivision rules in Riemannian and symmetric spaces. Advances in Computational Mathematics, 34(2):201–218, February 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9150-7.

Wei:2010:CAC

[WZ10] T. Wei and D. Y. Zhou. Convergence analysis for the Cauchy problem of Laplace’s equation by a regularized method of fundamental solutions. Advances in Computational Mathematics, 33(4):491–510, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9134-7.

Wang:2017:PGT

[WZL17] Tongke Wang, Zhiyue Zhang, and Zhifang Liu. The practical Gauss type rules for Hadamard finite-part integrals using Puiseux expansions. Advances in Computational Mathematics, 43(2):319–350, April 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9487-7.

Wu:2016:FHA

[WZtX16] Gang Wu, Lu Zhang, and Ting ting Xu. A framework of the harmonic Arnoldi method for evaluating φfunctions with applications to exponential integrators. Advances in Computational Mathematics, 42(3):505–541, June 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-015-9433-0.

Xiao:2005:MSA

[XB05] S. P. Xiao and T. Belytschko. Material stability analysis of particle methods. Advances in Computational Mathematics, 23(1–2):171–190, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1817-5.

Xiao:2013:MCC

[XH13] Junmin Xiao and Qiya Hu. Multilevel correction for collocation solutions of Volterra integral equations with proportional delays. Advances in Computational Mathematics, 39 (3–4):611–644, October 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9294-3.

Xiang:2015:EBF

[XHC15] Shuhuang Xiang, Guo He, and Yeol Je Cho. On error bounds of Filon–Clenshaw– Curtis quadrature for highly oscillatory integrals. Advances in Computational Mathematics, 41(3):573–597, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9377-9.

Xiang:2013:CQV

[Xia13]  Dao-Hong Xiang.  Condi-

tional quantiles with varying Gaussians. Advances in Computational Mathematics, 38(4):723–735, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9257-5.

Xie:2008:ENE

[Xie08] Hehu Xie. Extrapolation of the Nédélec element for the Maxwell equations by the mixed finite element method. Advances in Computational Mathematics, 29(2):135–145, August 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9047-2.

Xiong:2015:IES

[XKY15] Xiaogang Xiong, Ryo Kikuuwe, and Motoji Yamamoto. Implicit Euler simulation of onedimensional Burridge–Knopoff model of earthquakes with setvalued friction laws. Advances in Computational Mathematics, 41(6):1475–1493, December 2015. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9398-4.

Xu:2010:SND

[XL10] Huilin Xu and Jijun Liu. Stable numerical differentiation for the second order derivatives. Advances in Computational Mathematics, 33(4):431–447, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9132-9.

Xiao:2010:LNI

[XP10] Quan-Wu Xiao and Zhi-Wei Pan. Learning from nonidentical sampling for classification. Advances in Computational Mathematics, 33(1): 97–112, July 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1007/s10444-
                                        009-9123-x.

Xu:1998:MCF

[Xu98] Yuan Xu. Minimal cubature formulae for a family of radial weight functions. Advances in Computational Mathematics, 8(4):367–380, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018964818105.

Xu:2000:PIS

[Xu00] Yuan Xu. Polynomial interpolation in several variables, cubature formulae, and ideals. Advances in Computational Mathematics, 12(4):363–376, March 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018989707569. Mul-

tivariate polynomial interpolation.

Xu:2004:AMH

[Xu04a] Yuan Xu. Approximation by means of h-harmonic polynomials on the unit sphere. Advances in Computational Mathematics, 21(1–2):37–58, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/B:ACOM.0000016433.
27278.2a.

Xu:2004:PIU

[Xu04b] Yuan Xu. Polynomial interpolation on the unit sphere and on the unit ball. Advances in Computational Mathematics, 20(1–3):247–260, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1025851005416. Ap-

proximation and applications.

Xu:2014:LTE

[Xu14] Da Xu. The long time error analysis in the second order difference type method of an evolutionary integral equation with completely monotonic kernel. Advances in Computational Mathematics, 40(4):881–922, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9331-2.

Xiao:2013:SLA

[XWL13] Yunhai Xiao, Soon-Yi Wu, and Dong-Hui Li. Splitting and linearizing augmented Lagrangian algorithm for subspace recovery from corrupted observations. Advances in Computational Mathematics, 38(4):837–858, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9261-9.

Xu:2012:PSI

[XY12] Yuesheng Xu and Hongqi Yang. Preface [Special issue: Numerical methods for ill-posed problems]. Advances in Computational Mathematics, 36 (2):153–156, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9269-1;
 http://link.springer.com/
content/pdf/10.1007/s10444-
011-9269-1.pdf.

Xie:2015:ASF

[XY15] Hehu Xie and Xiaobo Yin. Acceleration of stabilized finite element discretizations for the Stokes eigenvalue problem. Advances in Computational Mathematics, 41(4):799–812, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9386-8.

Xu:2001:LPF

[XZ01] Jinchao Xu and Aihui Zhou.

Local and parallel finite element algorithms based on twogrid discretizations for nonlinear problems. Advances in Computational Mathematics, 14(4):293–327, May 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1012284322811.

Xu:2003:AWM

[XZ03] Yuesheng Xu and Qingsong Zou. Adaptive wavelet methods for elliptic operator equations with nonlinear terms. Advances in Computational Mathematics, 19(1–3):99–146, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1022816511874. Chal-

lenges in computational mathematics (Pohang, 2001).

Yan:2001:PEE

[Yan01]  Ningning Yan.  A posteri-

ori error estimators of gradient recovery type for elliptic obstacle problems. Advances in Computational Mathematics, 15(1–4):333–361, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1014284306804. A

posteriori error estimation and adaptive computational methods.

Yan:2003:PEE

[Yan03] Ningning Yan. A posteriori error estimators of gradient recovery type for FEM of a model optimal control problem. Advances in Computational Mathematics, 19(1–3):323–336, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1022800401298. Chal-

lenges in computational mathematics (Pohang, 2001).

Yao:2016:CAM

[Yao16]  Xudong Yao.  Convergence

analysis of a minimax method for finding multiple solutions of hemivariational inequality in Hilbert space. Advances in Computational Mathematics, 42(6):1331–1362, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9465-0.

Yang:1997:EAM

[YDF97] Geng Yang, Michel C. Delfour, and Michel Fortin. Error analysis of mixed finite elements for cylindrical shells. Advances in Computational Mathematics, 7(3):261–277, May 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018998903567.

Ye:2000:SFE

[Ye00] Xiu Ye. Stabilized finite element approximations for the Reissner–Mindlin plate. Advances in Computational Mathematics, 13(4):375–386, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016693613626.

Ye:2004:DSE

[Ye04] Xiu Ye. Discontinuous stable elements for the incompressible flow. Advances in Computational Mathematics, 20(4):333–345, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1027363218427.

Guo:2010:CLL

[yGjW10] Ben yu Guo and Tian jun Wang. Composite Laguerre– Legendre spectral method for exterior problems. Advances in Computational Mathematics, 32(4):393–429, May 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9112-5.

Guo:1998:AMA

[yGmW98] Ben yu Guo and Yuan ming Wang. An almost monotone approximation for a nonlinear two-point boundary value problem. Advances in Computational Mathematics, 8(1–2):65–96, January 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018983927675.

Guo:2009:LGC

[yGqW09] Ben yu Guo and Zhong qing Wang. Legendre–Gauss collocation methods for ordinary differential equations. Advances in Computational Mathematics, 30(3):249–280, April 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9067-6.

Guo:2008:IAT

[yGS08] Ben yu Guo and Jie Shen. Irrational approximations and their applications to partial differential equations in exterior domains. Advances in Computational Mathematics, 28(3):237–267, April 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-006-9020-5.

Guo:2003:SPA

[yGSlX03] Ben yu Guo, Jie Shen, and Cheng long Xu. Spectral and pseudospectral approximations using Hermite functions: application to the Dirac equation. Advances in Computational Mathematics, 19(1–3):35–55, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1022892132249. Chal-

lenges in computational mathematics (Pohang, 2001).

Guo:2007:EAS

[yGW07] Ben yu Guo and Li-Lian Wang. Error analysis of spectral method on a triangle. Advances in Computational Mathematics, 26(4):473–496, May 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7471-8.

Ying:2007:CAO

[Yin07] Yiming Ying. Convergence analysis of online algorithms. Advances in Computational Mathematics, 27(3):273–291, October 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-9002-z.

Yang:2012:EAB

[YLBL12] Yidu Yang, Qun Lin, Hai Bi, and Qin Li. Eigenvalue approximations from below using Morley elements. Advances in Computational Mathematics, 36(3):443–450, April 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9185-4.

Yoon:2001:CAA

[Yoo01]  Jungho Yoon.  Computa-

tional aspects of approximation to scattered data by using “shifted” thin-plate splines. Advances in Computational Mathematics, 14(4):329–359, May 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1012205804632.

Ying:2012:LCG

[YWC12] Yiming Ying, Qiang Wu, and Colin Campbell. Learning the coordinate gradients. Advances in Computational Mathematics, 37(3):355–378, October 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9211-6.

Ye:2008:LAG

[YZ08] Gui-Bo Ye and Ding-Xuan Zhou. Learning and approximation by Gaussians on Riemannian manifolds. Advances in Computational Mathematics, 29(3):291–310, October 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9049-0.

Ye:2013:FSI

[YZ13] Weicai Ye and Yongdong Zhang.

A fast solver for integral equations with convolution-type kernel. Advances in Computational Mathematics, 39(1):45–67, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9268-2.

Zheludev:2010:CIS

[ZA10] Valery A. Zheludev and Amir Z.

Averbuch. Computation of interpolatory splines via triadic subdivision. Advances in Computational Mathematics, 32(1):63–72, January 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9087-2.

Zala:1999:WAP

[ZB99] Cedric A. Zala and Ian Barrodale. Warping aerial photographs to orthomaps using thin plate splines. Advances in Computational Mathematics, 11(2–3):211–227, November 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018928026708. Ra-

dial basis functions and their applications.

Zhang:2012:NIG

[ZB12] Qinghui Zhang and Uday Banerjee. Numerical integration in Galerkin meshless methods, applied to elliptic Neumann problem with non-constant coefficients. Advances in Computational Mathematics, 37 (4):453–492, November 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9216-1.

Zhao:2014:RPE

[ZC14] Jikun Zhao and Shaochun Chen. Robust a posteriori error estimates for conforming discretizations of a singularly perturbed reaction–diffusion problem on anisotropic meshes. Advances in Computational Mathematics, 40(4):797–818, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9327-y.

zuCastell:2007:PIU

[zCFX07] Wolfgang zu Castell, Noemí Laín Fernández, and Yuan Xu. Polynomial interpolation on the unit sphere. II. Advances in Computational Mathematics, 26 (1–3):155–171, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-005-7510-5.

Chen:2009:DAM

[zCpST09] Hua zhu Chen, Jin ping Song, and Xue-Cheng Tai. A dual algorithm for minimization of the LLT model. Advances in Computational Mathematics, 31(1–3):115–130, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9097-0.

Zhao:2014:JPC

[ZD14] Lijing Zhao and Weihua Deng.

Jacobian-predictor–corrector approach for fractional differential equations. Advances in Computational Mathematics, 40 (1):137–165, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9302-7.

Zhao:2016:HOF

[ZD16] Lijing Zhao and Weihua Deng.

High order finite difference methods on non-uniform meshes for space fractional operators. Advances in Computational Mathematics, 42(2):425–468, April 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9430-3.

Zhang:2017:NAF

[ZD17] Zhijiang Zhang and Weihua Deng. Numerical approaches to the functional distribution of anomalous diffusion with both traps and flights. Advances in Computational Mathematics, 43(4):699–732, August 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9503-y.

Zhu:2016:ARM

[ZG16] Hongqiang Zhu and Zhen Gao. An h-adaptive RKDG method with troubled-cell indicator for one-dimensional detonation wave simulations. Advances in Computational Mathematics, 42(5):1081–1102, October 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9454-3.

Zhang:2012:NAS

[ZGGW12] Deyue Zhang, Yukun Guo, Chengchun Gong, and Guan Wang. Numerical analysis for the scattering by obstacles in a homogeneous chiral environment. Advances in Computational Mathematics, 36(1):3–20, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9169-9.

Zhang:2001:PEE

[Zha01]  Zhimin Zhang.  A posteri-

ori error estimates on irregular grids based on gradient recovery. Advances in Computational Mathematics, 15(1– 4):363–374, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014221409940. A

posteriori error estimation and adaptive computational methods.

Zhang:2005:PWF

[Zha05] Zhihua Zhang. Periodic wavelet frames. Advances in Computational Mathematics, 22 (2):165–180, February 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-3139-z.

Zhang:2014:SFV

[Zha14] Tie Zhang. Superconvergence of finite volume element method for elliptic problems. Advances in Computational Mathematics, 40(2):399–413, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9313-4.

Zheludev:2006:ISS

[Zhe06] Valery A. Zheludev. Interpolatory subdivision schemes with infinite masks originated from splines. Advances in Computational Mathematics, 25 (4):475–506, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-4149-6.

Zhu:2015:SLC

[Zhu15] Wei Zhu. Simulation of liquid crystal elastomers using Chebyshev spectral method with a new preconditioner. Advances in Computational Mathematics, 41(4):853–879, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9389-5.

Zhi:2016:HWM

[ZhXpZ16] Shi Zhi, Yan hua Xu, and Jun ping Zhao. Haar wavelets method for solving Poisson equations with jump conditions in irregular domain. Advances in Computational Mathematics, 42(4):995–1012, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-015-9450-z.

Zhou:2006:APK

[ZJ06] Ding-Xuan Zhou and Kurt Jetter. Approximation with polynomial kernels and SVM classifiers. Advances in Computational Mathematics, 25(1–3): 323–344, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1007/s10444-
                                        004-7206-2.

Zheng:2016:IFA

[ZL16] Xiaoming Zheng and John Lowengrub. An interface-fitted adaptive mesh method for elliptic problems and its application in free interface problems with surface tension. Advances in Computational Mathematics, 42(5):1225–1257, October 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9460-5.

Zhang:2017:SSS

[ZL17] Jin Zhang and Xiaowei Liu.

Supercloseness of the SD-FEM on Shishkin triangular meshes for problems with exponential layers. Advances in Computational Mathematics, 43(4):759–775, August 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9505-9.

Zozulya:2015:RDI

[Zoz15a]  V. V. Zozulya.  Regular-

ization of divergent integrals: a comparison of the classical and generalized-functions approaches. Advances in Computational Mathematics, 41(4):727–780, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9399-3.

Zozulya:2015:RAR

[Zoz15b] V. V. Zozulya. RETRACTED ARTICLE: Regularization of divergent integrals: An approach using generalized functions. Advances in Computational Mathematics, 41(4):953, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9321-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        013-9321-4.pdf.

Zampieri:2006:ESO

[ZP06] Elena Zampieri and Luca F. Pavarino. An explicit second order spectral element method for acoustic waves. Advances in Computational Mathematics, 25(4):381–401, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7626-z.

Zhang:2003:EMU

[ZPR03] Lei Zhang, Haesun Park, and J. Ben Rosen. Exponential modeling with unknown model order using structured nonlinear total least norm. Advances in Computational Mathematics, 19(1–3):307–322, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1022881907141. Chal-

lenges in computational mathematics (Pohang, 2001).

Zhang:2015:LTS

[ZPY15] Tong Zhang, Damazio Pedro, and JinYun Yuan. A large time stepping viscosity-splitting finite element method for the viscoelastic flow problem. Advances in Computational Mathematics, 41(1):149–190, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9353-4.

Zhu:2013:ARM

[ZQ13] Hongqiang Zhu and Jianxian Qiu. An h-adaptive RKDG method with troubled-cell indicator for two-dimensional hyperbolic conservation laws. Advances in Computational Mathematics, 39(3–4):445–463, October 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-012-9287-7.

Zhang:2001:SFE

[ZSB01a] L. Zhang, T. Strouboulis, and I. Babuška. η%-superconvergence of finite element solutions and error estimators. Advances in Computational Mathematics, 15(1–4):393–404, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014267804986. A

posteriori error estimation and adaptive computational methods.

Zhang:2001:PEF

[ZSB01b] L. Zhang, T. Strouboulis, and I. Babuška. A posteriori estimators for the FEM: analysis of the robustness of the estimators for the Poisson equation. Advances in Computational Mathematics, 15(1–4):375–392, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1014232321825. A

posteriori error estimation and adaptive computational methods.

Zheng:2015:PUP

[ZSHZ15] Haibiao Zheng, Lina Song, Yanren Hou, and Yuhong Zhang. The partition of unity parallel finite element algorithm. Advances in Computational Mathematics, 41(4):937–951, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9392-x.

Zak:2014:NSC

[ZT14] Mohammad Khorsand Zak and Faezeh Toutounian. Nested splitting CG-like iterative method for solving the continuous Sylvester equation and preconditioning. Advances in Computational Mathematics, 40(4):865–880, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9330-3.

Zhang:2015:SFV

[ZT15] Tie Zhang and Lixin Tang.

A stabilized finite volume method for Stokes equations using the lowest order P_{1}– P_{0} element pair. Advances in Computational Mathematics, 41(4):781–798, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9385-9.

Zhang:2008:SCE

[ZV08] Chengjian Zhang and Stefan Vandewalle. Stability criteria for exact and discrete solutions of neutral multidelay-integro-differential equations. Advances in Computational Mathematics, 28(4):383–399, May 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9037-4.

Zeng:1995:NSI

[ZVW95] S. Zeng, C. Vuik, and P. Wesseling. Numerical solution of the incompressible Navier–Stokes equations by Krylov subspace and multigrid methods. Advances in Computational Mathematics, 4(1–2):27–49, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02123472.

Zheng:2012:NRM

[ZW12] G. H. Zheng and T. Wei. A new regularization method for a Cauchy problem of the time fractional diffusion equation. Advances in Computational Mathematics, 36 (2):377–398, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9206-3.

Zeng:2016:DSS

[ZWDD16] Chao Zeng, Meng Wu, Fang Deng, and Jiansong Deng. Dimensions of spline spaces over non-rectangular T-meshes. Advances in Computational Mathematics, 42(6):1259–1286, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1007/s10444-
                                        016-9461-4.

Zwick:1994:PMA

[Zwi94] D. Zwick. A planar minimax algorithm for analysis of coordinate measurements. Advances in Computational Mathematics, 2(4):375–391, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF02521605.

Zou:2012:GBE

[ZXC12] Bin Zou, Zongben Xu, and Xiangyu Chang. Generalization bounds of ERM algorithm with V -geometrically ergodic Markov chains. Advances in Computational Mathematics, 36(1):99–114, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9182-7.

Zhang:2016:OCS

[ZYY16] Hui Zhang, Ming Yan, and Wotao Yin. One condition for solution uniqueness and robustness of both ℓ_{1}- synthesis and ℓ_{1}-analysis minimizations. Advances in Computational Mathematics, 42(6): 1381–1399, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-016-9467-y.

Zhou:2009:HOP

[ZZ09] Xiang-Jun Zhou and Ding-Xuan Zhou. High order Parzen windows and randomized sampling. Advances in Computational Mathematics, 31 (4):349–368, November 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9073-8.